### v2

# Field transformations in functional integral, effective action and functional flow equations

Abstract
Field transformations for the quantum effective action lead to different pictures of a given physical situation, as describing a given evolution of the universe by different geometries. Field transformations for functional flow equations can reveal important physical features, as the appearance of bound states. They also allow for technical simplifications. We make a critical assessment of the virtues and shortcomings of different versions of field transformations. Key issues are demonstrated by the quantum field theory for scalars with a field-dependent coefficient of the kinetic term. Our findings confirm the principle of field relativity for gravity and cosmology based on the quantum effective action.
Introduction
Field transformations are widely used in many areas of physics. In the presence of fluctuations – quantum or thermal – it matters for which object the transformations are performed. For example, in quantum gravity a field transformation of the classical action yields different results as the transformation of the quantum effective action. This difference results in the role of fluctuations. A proper field transformation of the functional integral has to include the transformation of the functional measure. The corresponding Jacobian is often not taken into account, leading to statements that theories related by field transformations are not equivalent on the quantum level. In cosmology and quantum gravity the omission of the Jacobian has led to a long debate on the equivalence of models related by field transformation, see the review Faraoni et al. (1998). In condensed matter physics this issue concerns the difference of making field transformations in the Hamiltonian or the free energy. They are only equivalent if the transformation of the measure factor is taken into account.
In the presence of fluctuations the relevant field equations are the ones obtained by variation of the quantum effective action or free energy with respect to the fields. They are exact identities for a given form of the effective action and contain the relevant physical information, in contrast to the field equations derived by variation of the classical action. Field transformations of the effective action, that we call here “macroscopic field transformations”, lead to identical field equations, simply related by a variable transformations in differential equations. The quantum effective action or free energy contains much more information than only the field equations. The second functional derivative is related to the connected correlation function or inverse propagator, and interaction vertices follow from higher functional derivatives. After a field transformation the higher functional derivatives of the effective action may no longer be directly related to microscopic correlation functions. They still play an important role for macroscopic observations which are often not sensitive to the microscopic details or microscopic correlation functions. A well known example is the universality of critical phenomena where vastly different microscopic theories can lead to the same free energy or effective action. The second functional derivative of the effective action contains the information about the correlation length independently of the microscopic model. We name the functional derivatives of the effective action “macroscopic correlation functions” and discuss how their relation to microscopic correlation functions is affected by field transformations. This paper aims for an overview of the impact of field transformations on different levels. In the view of recent discussions we put particular emphasis on the technical simplifications that can be realized by field transformations in functional flow equations.
Field transformations can help to focus on important physical features and to discard redundant information. For critical phenomena in statistical physics their use has a long tradition Wegner (1974). Various versions of field transformations have been formulated for modern functional flow equations Wetterich (1996); Gies and Wetterich (2002, 2004); Pawlowski (2007); Floerchinger and Wetterich (2009). They are used for making important “composite or collective degrees of freedom” of a model directly accessible or for technical simplifications. Motivated by the elimination of redundant couplings for the counting of relevant parameters at a given fixed point of renormalization group equations Wegner (1974), the proposal of the “essential renormalization group” Baldazzi et al. (2022); Baldazzi and Falls (2021) has argued that couplings which can be removed by field transformations are not observable and need not to be considered for the search of possible scaling solutions. The essential renormalization group entails important technical simplifications for approximate solutions of functional flow equations Knorr (2023a); Knorr et al. (2022a); Laporte et al. (2023); Ihssen and Pawlowski (2023a); Buccio and Percacci (2022); Knorr et al. (2022b); de Brito et al. (2023); Saueressig (2023); Ihssen and Pawlowski (2023b); Kawai and Ohta (2023); Knorr (2023b); Baldazzi et al. (2023). A critical discussion of the merits and shortcomings of this approach seems timely.
Removing couplings by field transformations can be of great help but also entails substantial dangers. This may be demonstrated by two cases. The introduction of renormalized fields removes a field-independent and momentum-independent wave function renormalization $Z$. This quantity is not observable for purely macroscopic observations. The wave function renormalization only plays a role if macroscopic observables are related to microphysical degrees of freedom. For example, it is crucial for understanding Wetterich (1993a); Gräter and Wetterich (1995); v. Gersdorff and Wetterich (2001) the compatibility of the Kosterlitz-Thouless phase transition Kosterlitz and Thouless (1973) with the Mermin-Wagner theorem Mermin and Wagner (1966). If $Z$ depends on some renormalization scale $k$, its logarithmic $k$-derivative, $\eta=-k\partial_{k}\ln Z$, defines an anomalous dimension. This appears in renormalization group or flow equations once $Z$ is removed by a field transformation. The same anomalous dimension often governs the details of the momentum- or field-dependence of the propagator, as encoded by the replacement of $Z$ by a momentum- and field-dependent factor $Z(q,\chi)$ multiplying the kinetic term.
If the dependence of $Z(q,\chi)$ on the momenta $q$ and fields $\chi$ is structureless and the anomalous dimension is small, it seems a promising idea to absorb $Z(q,\chi)$ by a field transformation. This simplification of the propagator facilitates many computations. In particular, for a field-independent $Z(q)$ a linear field transformation moves the information contained in $Z(q)$ to “dressed vertices” by multiplying the vertices with appropriate powers of $Z^{-1/2}(q)$. The dressed vertices are the same for all original choices of fields which are related by a linear momentum-dependent field transformation. In this sense the dressed vertices contain the “physical” information.
Our second example points to the opposite direction. The inverse propagator of electrons in a solid is of the form $P_{F}(q)=\varepsilon(q)-\mu+i(2n+1)\pi T$, where $\mu$ is the chemical potential, $T$ the temperature and the integer $n$ determines the Matsubara frequency. The Fermi surface is determined by $\varepsilon(q_{F})=\mu$. It is located at non-zero Fermi momenta $q_{F}$ even for the simplest case $\varepsilon(q)=q^{2}/(2m)$. In solids $\varepsilon(q)$ can have a rather complex form which fixes the detailed location of the Fermi surface and many other directly observable features. For $T\neq 0$ a regular linear field transformation transforms $P_{F}(q)\to\tilde{P}_{F}(q)=f(q)P_{F}(q)$. We can bring $\tilde{P}_{F}(q)$ to an arbitrarily chosen form without zeros. After the field transformation $\tilde{P}_{F}(q)$ does no longer contain the information about the Fermi surface if we take some standard form as $\tilde{P}_{F}=q^{2}/(2m)+i(2n+1)\pi T$ for a gas of free fermions or bosons at zero density. The corresponding complex function $f(q)$ has removed the chemical potential and all structures in the dispersion relation $\varepsilon(q)$. The details of the dispersion relation involves couplings that parameterize the electron propagator in a solid. For the two-dimensional Hubbard model with next-neighbor and next-to-nearest neighbor hopping parameters $t$ and $t^{\prime}$ one has
$\varepsilon(q)=-2t(\cos q_{x}+\cos q_{y})-4t^{\prime}\cos q_{x}\cos q_{y}\ .$ (1)
One of the couplings that is removed by this field transformation is the ratio $t^{\prime}/t$.
The conclusion that couplings used to parameterize the momentum-dependence of the fermion propagator are unobservable is not justified. The physical information contained in these couplings is moved by the field transformation to the vertices. Only certain products of vertices with powers of $P_{F}^{-1/2}(q)$ are invariant under the momentum-dependent field transformation. If one would use a truncation which neglects the detailed momentum-dependence of vertices, key physical information would even be lost by a field transformation which brings the electron propagator to some standard form.
The danger of discarding important information, which is already visible for linear field transformations, gets greatly enhanced for non-linear field transformations. The key issue is the control to which sector the information about observables is moved by a field transformation. Choosing optimal fields for the solution of a given problem is mainly a question of simplicity. Since truncations of the effective action are usually needed, it is crucial to encode the most relevant physical information in the part that is not discarded by the truncation. If important physical features are visible in the propagator, it seems to be a good idea to keep as much information as possible in the field- and momentum-dependent propagator, extended to low order vertices with perhaps less complete resolution.
Different versions of field transformations for functional flow equations have been proposed in the past. These functional flow equations describe the dependence of the effective average action $\Gamma_{k}$ on a renormalization scale $k$ which typically corresponds to an infrared cutoff. Only fluctuations with momenta greater than $k$ are then included for $\Gamma_{k}$. For $k\to 0$ the effective average action becomes the usual effective action with all quantum or thermal fluctuations included. The present paper tries to compare the advantages and disadvantages of main representatives of field transformations in functional flow equations. Particular emphasis is devoted to the issue of control of the field transformation. We also address the question to which extent the effective actions obtained by different versions of field transformations are equivalent. Furthermore, we assess the relation between macroscopic correlation functions obtained from functional derivatives of the effective action and microscopic correlation functions as defined by expectation values of powers of microscopic fields.
The possibility of field transformations has also important conceptual implications. This is most striking for gravity. A non-linear transformation of the metric field changes the geometry. For example, an expanding universe can be transformed into a shrinking universe Wetterich (2013a); Bars et al. (2013); Wetterich (2014a, b); Li (2014); Hossain et al. (2015); Domènech and Sasaki (2015); Ijjas and Steinhardt (2015); Cai and Piao (2016); Domènech and Sasaki (2016); Kamenshchik et al. (2016); Graef et al. (2017); Amendola et al. (2018); Shimon (2019); Luan and Qiu (2020); Casadio et al. (2021); Koivisto and Zheng (2021); Rubakov and Wetterich (2022); Mukherjee and Sahota (2023); Lombriser (2023). Field relativity Wetterich (2013a) states that macroscopic field transformations of the quantum effective action do not change the observables. Cosmologies with different geometries related by such transformations are strictly equivalent. The present paper confirms this point of view. Field relativity also establishes the equivalence of many models of modified gravity with coupled quintessence Wetterich (2014c).
As a simple example illustrating various aspects of field transformations we focus in this paper on a quantum field theory for a single scalar, with a classical action involving only a term with two derivatives, multiplied by a field-dependent wave function renormalization or “kinetial”. On the one hand a non-linear field transformation can transform away the field-dependence of the kinetial such that the transformed action looks like the one for a free massless scalar field. On the other hand the field-dependence of the kinetial leads to terms in the action involving more than two powers of the field, suggesting interactions. We will relate these various aspects and find that for generic kinetials this model indeed describes an interacting scalar field.
While a general field-dependent kinetial describes a theory with interactions, we find strong indications for an ultraviolet fixed point for which the effective average action is given for $\chi\to 0$ by
$\Gamma_{k}[\chi]=\int_{x}\left(\frac{\kappa_{0}}{2\chi^{2}}\partial^{\mu}\chi\partial_{\mu}\chi+\text{const.}\right)\ .$. ) . (2)
This describes a free massless composite field ${\varphi}\sim\ln\chi$. This ultraviolet fixed point differs from the (trivial) infrared fixed point for which the kinetial is field-independent. The ultraviolet fixed point is characterized by the quantum scale symmetry of a constant multiplicative rescaling $\chi\to\beta\chi$ with fixed metric. In contrast, the infrared fixed point exhibits a different form of quantum scale symmetry for which the metric (or coordinates) are rescaled together with $\chi$. It also is invariant under constant shifts $\chi\to\chi+\delta$. Possible crossover trajectories between the two fixed points are characterized by the presence of interactions. If the ultraviolet fixed point (2) is confirmed this would define a possible ultraviolet completion of an interacting scalar field theory in four dimensions, avoiding the “triviality problem”.
In sect. II we begin with field-transformations for the integration variables of the functional integral defining the quantum field theory. The Jacobian of this transformation modifies the action beyond a simple replacement of the field variables. We demonstrate how non-linear field transformations change the saddle point approximation and perturbation theory. Sect. III turns to field transformations for the quantum effective action for which all effects of quantum fluctuations are included. We distinguish between macroscopic field transformations which only modify the expression of a given effective action in terms of field-variables, and microscopic field transformations which modify in addition the sources used for the definition of the effective action.
In sect. IV we proceed to the field transformations in functional flow equations, focusing first on the macroscopic field transformations Wetterich (1996); Gies and Wetterich (2002). This is contrasted in sect. V with the microscopic field transformations Pawlowski (2007); Baldazzi et al. (2022); Baldazzi and Falls (2021); Ihssen and Pawlowski (2023b) which are the basis of the proposed “essential renormalization group”. In sect. VI we employ a formulation with both fundamental and composite fields in order to clarify further aspects of the microscopic field transformations, in particular the relation between the classical action and the initial values of the flow equations. Sect. VII concentrates on scaling solutions of the flow equations whose existence can make a model ultraviolet complete. Our conclusions are presented in sect. VIII. Two appendices provide a discussion of the model with a field-dependent kinetial without the use of field transformations. This can be used for comparison with treatments involving field transformations. These computations yield indications for the exitence of the ultraviolet fixed point (2).
Field transformations for functional
integrals
Field transformations in a functional integral are conceptually very simple. They are the analogue of variable transformations in ordinary integrals. The Jacobian of the transformation plays an important role and is often difficult to handle in practice. We demonstrate this for a simple model of a scalar field with a field-dependent kinetic term. The classical action involves no potential or mass term, while the coefficient of the term with two derivatives - the kinetial - is allowed to depend on the value of the scalar field. A field transformation can transform this classical action to the one for a free massless scalar field. The Jacobian induces, however, additional terms in the relevant microscopic action. By virtue of the Jacobian the interactions contained in the kinetial are shifted to a potential.
Functional integral
Let us define our model by the “classical” or “microscopic” action
$S[\chi^{\prime}]=\frac{1}{2}\int_{x}K(\chi^{\prime})\partial^{\mu}\chi^{\prime}\partial_{\mu}\chi^{\prime}\ ,$ (3)
with $\partial^{\mu}\chi^{\prime}=\eta^{\mu\nu}\partial_{\nu}\chi^{\prime}$, $\eta^{\mu\nu}$ the metric of flat Euclidean or Minkowski space, $\int_{x}$ the integral over $d$-dimensional space $\mathbb{R}^{d}$ with Cartesian coordinates $x^{\mu}$, and $\partial_{\mu}\chi^{\prime}=\partial\chi^{\prime}/\partial x^{\mu}$. For Euclidean signature the action defines a probability density $p[\chi^{\prime}]$ for field configurations $\chi^{\prime}(x)$,
$p[\chi^{\prime}]=Z^{-1}\exp\big{(}-S[\chi^{\prime}]\big{)}\ ,$ (4)
with partition function given by the functional integral
$Z=\int\mathcal{D}\chi^{\prime}\,\exp\big{(}-S[\chi^{\prime}]\big{)}\ .$ (5)
The functional measure is formally defined by
$\int\mathcal{D}\chi^{\prime}=\prod_{x}\int_{-\infty}^{\infty}\text{d}\chi^{\prime}(x)\ .$ (6)
This formal definition may be considered as the continuum limit of some lattice regularization. A field $\chi^{\prime}$ with linear measure (6) and action expressed as functional of $\chi^{\prime}$ will later be called a “canonical field”.
One is typically interested in correlation functions as
$G(x,y)=\langle\chi^{\prime}(x)\chi^{\prime}(y)\rangle-\langle\chi^{\prime}(x)\rangle\langle\chi^{\prime}(y)\rangle\ ,$ (7)
or similar higher order correlation functions. As usual, the expectation value for an observable is given by
$\langle A\rangle=Z^{-1}\int\mathcal{D}\chi^{\prime}A[\chi^{\prime}]\exp\big{(}-S[\chi^{\prime}]\big{)}\ .$ (8)
(We will later generalize this expression to the presence of sources.) If the “kinetial” $K(\chi^{\prime})$ is not constant, the model involves interactions, as manifest in the classical field equation obtained by variation of the action
$K\partial^{\mu}\partial_{\mu}\chi^{\prime}=-\frac{1}{2}\frac{\partial K}{\partial\chi^{\prime}}\partial^{\mu}\chi^{\prime}\partial_{\mu}\chi^{\prime}\ .$ (9)
We will see that these interactions can lead to complex phenomena as spontaneous symmetry breaking.
Field transformations
By a non-linear field transformation
${\varphi}^{\prime}(x)={\varphi}^{\prime}\big{(}\chi^{\prime}(x)\big{)}\ ,$ (10)
defined by
$\frac{\partial{\varphi}^{\prime}}{\partial\chi^{\prime}}=K^{1/2}(\chi^{\prime})\ ,$ (11)
one can bring the classical action (3) to the form of a free scalar field theory
$S=\frac{1}{2}\int_{x}\partial^{\mu}{\varphi}^{\prime}\partial_{\mu}{\varphi}^{\prime}\ .$ (12)
This is, however, not the relevant action for the scalar field ${\varphi}^{\prime}$. The non-linear field transformation involves a Jacobian in the functional integral
$Z=$ $\,\int\mathcal{D}\chi^{\prime}\,e^{-S[\chi^{\prime}]}=\int\mathcal{D}{\varphi}^{\prime}\,J[{\varphi}^{\prime}]e^{-S\big{[}\chi^{\prime}({\varphi}^{\prime})\big{]}}$
$=$ $\,\int\mathcal{D}{\varphi}^{\prime}\,e^{-\big{(}S[{\varphi}^{\prime}]-\ln J[{\varphi}^{\prime}]\big{)}}=\int\mathcal{D}{\varphi}^{\prime}\,e^{-\bar{S}[{\varphi}^{\prime}]}\ ,$ (13)
with
$J[{\varphi}^{\prime}]=\prod_{x}\left|\frac{\partial\chi^{\prime}(x)}{\partial{\varphi}^{\prime}(x)}\right|=\exp\left(-\varepsilon^{-d}\int_{x}\ln\left|\frac{\partial{\varphi}^{\prime}}{\partial\chi^{\prime}}\right|\right)\ .$ (14)
Here $\varepsilon$ is a typical length scale (lattice distance) related to the regularization, defined by $\sum_{x}=\varepsilon^{-d}\int_{x}$.
As a consequence, the action for ${\varphi}^{\prime}$ involves a potential term reflecting the interactions
$\bar{S}[{\varphi}^{\prime}]=\int_{x}\left\{\frac{1}{2}\partial^{\mu}{\varphi}^{\prime}\partial_{\mu}{\varphi}^{\prime}+V({\varphi}^{\prime})\right\}\ ,$ (15)
where
$V({\varphi}^{\prime})=\frac{\varepsilon^{-d}}{2}\ln K[{\varphi}^{\prime}]\ .$ (16)
We suppose that $K[\chi^{\prime}]$ is positive for all $\chi^{\prime}$, such that taking the positive root in eq. (11) results in ${\varphi}^{\prime}$ monotonically increasing with $\chi^{\prime}$.
Inversely, many scalar theories with a potential and action (15) can be cast by a non-linear field transformation into the form (3), with kinetial
$K=\exp\big{(}2\varepsilon^{d}V\big{)}\ .$ (17)
After the field transformation ${\varphi}^{\prime}\to\chi^{\prime}({\varphi}^{\prime})$ the potential is no longer present, $V(\chi^{\prime})=0$. It has been absorbed by the Jacobian. As an example we may take
$V=\frac{\varepsilon^{-d}}{2\lambda}\mathrm{tanh}^{2}\left\{\frac{\lambda\varepsilon^{d/2}}{2}\big{(}{\varphi}^{\prime 2}-{\varphi}_{0}^{2}\big{)}\right\}\ ,$ (18)
where ${\varphi}_{0}^{2}$ is a positive or negative constant. For small $|{\varphi}^{\prime 2}-{\varphi}_{0}^{2}|\ll\varepsilon^{-d/2}/\lambda$ we can expand
$V=\frac{\lambda}{8}\big{(}{\varphi}^{\prime 2}-{\varphi}_{0}^{2}\big{)}^{2}+\ldots$ (19)
For ${\varphi}_{0}^{2}>0$ and small $\lambda>0$ this is a typical model with spontaneous symmetry breaking in the perturbative range. The kinetial (17) reads
$K=\exp\left\{\frac{1}{\lambda}\mathrm{tanh}^{2}\left[\frac{\lambda\varepsilon^{d/2}}{2}\big{(}{\varphi}^{\prime 2}-{\varphi}_{0}^{2}\big{)}\right]\right\}\ .$ (20)
If the potential $V({\varphi}^{\prime})$ in a model with canonical kinetic term leads to spontaneous symmetry breaking, the same holds after the field transformation in the model without potential, but with a non-trivial kinetial. This is a simple demonstration that a non-trivial kinetial can lead to complex phenomena and should not be “discussed away”.
For ${\varphi}^{\prime 2}$ near ${\varphi}_{0}^{2}$ one has $K\approx 1$ such that $\chi^{\prime}$ and ${\varphi}^{\prime}$ almost coincide. On the other hand, for ${\varphi}^{\prime 2}\to\infty$ one infers $K\approx\exp(1/\lambda)$ and therefore
$\chi^{\prime}\approx e^{-\frac{1}{2\lambda}}{\varphi}^{\prime}+\bar{\chi}\ ,$ (21)
where the constant $\bar{\chi}$ can be found by numerical integration of eq. (11). We choose $\chi^{\prime}({\varphi}^{\prime}=0)=0$ such that $\chi^{\prime}(-{\varphi}^{\prime})=-\chi^{\prime}({\varphi}^{\prime})$ and $K_{0}=K(\chi^{\prime}=0)\approx 1+\lambda\varepsilon^{d}{\varphi}_{0}^{4}/4$. For ${\varphi}_{0}^{2}>0$ and $\chi^{\prime}_{0}=\chi^{\prime}({\varphi}_{0})$ one has $K(\chi^{\prime}_{0})=1$. In this case $K(\chi^{\prime})$ interpolates smoothly between $K(\chi^{\prime}=0)=K_{0}$, $K(\chi^{\prime}_{0})=1$ and $K\big{(}|\chi^{\prime}|\to\infty\big{)}\approx\exp(1/\lambda)$. A minimum of $K(\chi^{\prime})$ at $\chi^{\prime}_{0}\neq 0$ results in a minimum of $V({\varphi}^{\prime})$ at ${\varphi}_{0}={\varphi}^{\prime}(\chi^{\prime}_{0})$. Nothing prevents us to make this choice of $K(\chi^{\prime})$. Our simple example demonstrates that large classes of scalar theories, including ones with non-trivial features as spontaneous symmetry breaking, can be described by the classical action (3) for a suitable kinetial $K(\chi^{\prime})$.
The partition function does not depend on the choice of the integration variable $\chi^{\prime}$ or ${\varphi}^{\prime}$, provided one accounts properly for the Jacobian,
$Z=\int\mathcal{D}\chi^{\prime}\,\exp\big{(}-S[\chi^{\prime}]\big{)}=\int\mathcal{D}{\varphi}^{\prime}\,\exp\big{(}-\bar{S}[{\varphi}^{\prime}]\big{)}\ .$ (22)
Also the probability distribution (for Euclidean signature) is the same
$Z^{-1}\mathcal{D}\chi^{\prime}\,\exp\big{(}-S[\chi^{\prime}]\big{)}=Z^{-1}\mathcal{D}{\varphi}^{\prime}\,\exp\big{(}-\bar{S}[{\varphi}^{\prime}]\big{)}\ .$ (23)
One concludes that all correlation functions are identical for both formulations, provided they are expressed properly in terms of $\chi^{\prime}$ or ${\varphi}^{\prime}$. For example, the connected two-point function (7) involves non-linear expressions in ${\varphi}^{\prime}$ once $\chi^{\prime}({\varphi}^{\prime})$ is inserted. Simple correlation functions for one choice of fields may look rather complicated for another choice of fields. Nevertheless, a possible expectation value $\langle\chi^{\prime}(x)\rangle=\chi^{\prime}_{0}$ translates directly to an expectation value $\langle{\varphi}^{\prime}(x)\rangle={\varphi}^{\prime}(\chi^{\prime}_{0})$.
Saddle point approximation
In principle, it does not matter which fields are chosen for the functional integral. This does no longer hold for given approximations. We demonstrate this for the saddle point approximation for our model. In the tree approximation one evaluates $\langle{\varphi}^{\prime}\rangle$ by a solution of the field equation derived from the classical action $\bar{S}[{\varphi}^{\prime}]$. For our example with ${\varphi}_{0}^{2}>0$ this results in $\langle{\varphi}^{\prime}(x)\rangle=\pm{\varphi}_{0}$, as given by the minimum of the potential $V({\varphi}^{\prime})$. In contrast, the tree approximation for $\langle\chi^{\prime}\rangle$, defined by a solution of the field equation (9) derived by variation of $S[\chi^{\prime}]$, admits arbitrary constant $\langle\chi^{\prime}(x)\rangle=\text{const.}$. The tree approximation in the two pictures looks rather different. This is not surprising, since the tree approximation does not account for the role of the Jacobian appearing in the treatment of fluctuations. Performing a loop expansion in both schemes the difference in the loop contributions has to match the difference in the tree approximations such that the result of $\langle{\varphi}^{\prime}(x)\rangle$ is identical in both schemes.
Let us first consider a saddle point approximation in the $\chi$-scheme with action (3). In one-loop order the effective action reads (see below)
$\Gamma[\chi]=S[\chi]+\frac{1}{2}\mathrm{tr}\big{\{}\ln S^{(2)}[\chi]\big{\}}\ ,$ (24)
with $S^{(2)}$ the second functional derivative of $S$,
$S^{(2)}(x,y)=-\Big{[}$ $\,K\partial^{\mu}\partial_{\mu}+\partial_{\chi}K\partial^{\mu}\chi\partial_{\mu}+\frac{1}{2}\partial^{2}_{\chi}K\partial^{\mu}\chi\partial_{\mu}\chi$
$\,+\partial_{\chi}K\partial^{\mu}\partial_{\mu}\chi\Big{]}(x)\delta(x-y)\ .$ (25)
Here all derivatives are taken with respect to $x$ and $\partial_{\chi}K=\partial K/\partial\chi$. The argument $\chi(x)$ of the effective action is the “macroscopic” or “background” field which can be identified with the expectation value $\langle\chi^{\prime}(x)\rangle$ in the presence of suitable sources.
The potential term $U(\chi)$ in the effective action obtains by evaluating eq. (24) for a constant background field,
$\Gamma[\chi]=\int_{x}U(\chi)=\frac{1}{2}\mathrm{tr}\big{\{}\ln K(\chi)+\ln\big{(}-\partial^{\mu}\partial_{\mu}\big{)}\big{\}}\ .$ (26)
For the first term we evaluate the trace in position space,
$\mathrm{tr}\ln K=\sum_{x}\ln K=\varepsilon^{-d}\int_{x}\ln K\ ,$ (27)
and we extract
$U(\chi)=\frac{\varepsilon^{-d}}{2}\ln K(\chi)\ .$ (28)
This is precisely the potential (16) which is present in the tree approximation in the ${\varphi}$-scheme. The trace for the second term may be evaluated in momentum space where $-\partial^{\mu}\partial_{\mu}$ translates to the squared momentum $q^{2}=q^{\mu}q_{\mu}$. The second term yields a field-independent constant which plays no role.
For a general $x$-dependent field $\chi(x)$ we write
$\mathrm{tr}\big{\{}\ln S^{(2)}\big{\}}=\mathrm{tr}\ln K+\mathrm{tr}\ln\Delta\ ,$ (29)
with
$\Delta=K^{-1}S^{(2)}=-\big{(}\partial^{\mu}\partial_{\mu}+B^{\mu}\partial_{\mu}+C\big{)}\ ,$ (30)
where
$B^{\mu}=\partial_{\chi}\ln K\partial^{\mu}\chi\ ,$ (31)
and
$C=\frac{1}{2K}\partial_{\chi}^{2}K\partial^{\mu}\chi\partial_{\mu}\chi+\partial_{\chi}\ln K\partial^{\mu}\partial_{\mu}\chi\ .$ (32)
The first term in eq. (29) yields the potential (28), while the second term contributes derivative terms to the effective action.
This correction to the kinetic terms needs a regularization, as can be seen by evaluating traces of functions of $\Delta$ in the heat kernel expansion. Expanding in the number of derivatives one finds
$\mathrm{tr}f(\Delta)=\frac{1}{16\pi^{2}}\int_{x}\int_{0}^{\infty}\text{d}zf(z)\big{(}(z+H(\chi)\partial^{\mu}\chi\partial_{\mu}\chi+\ldots\big{)}\ .$ (33)
Here $H(\chi)$ is defined by
$\int_{x}H(\chi)\partial^{\mu}\chi\partial_{\mu}\chi=\int_{x}\big{(}C-\frac{1}{4}B^{\mu}B_{\mu}\big{)}\ ,$ (34)
which yields
$H(\chi)=\frac{3}{4}\big{(}\partial_{\chi}\ln K\big{)}^{2}-\frac{1}{2K}\partial_{\chi}^{2}K\ .$ (35)
The dots in eq. (33) stand for terms with more than two derivatives. For $f(z)=\ln(z)$ the $z$-integral does not converge. Even without specifying a regularization it is clear, however, that the term $\sim H(\chi)$ results in a renormalization of the kinetial, $K\to K+cH$. This does not affect the minimum of the potential which determines the expectation value $\langle\chi^{\prime}(x)\rangle=\chi_{0}$. We conclude that the expectation value of the scalar field is the same for the tree approximation in the ${\varphi}$-scheme and the one-loop approximation in the $\chi$-scheme.
Going further by including the one-loop correction in the ${\varphi}$-scheme one finds a renormalization of the effective scalar potential. (The one-loop correction needs again a regularization.) This modifies the expectation value $\langle{\varphi}^{\prime}(x)\rangle$. In turn, this change has to be matched by higher loop terms in the $\chi$-scheme. We conclude that a change of variables in the functional integral is allowed, in principle. In practice, it may be rather cumbersome. This holds even if the Jacobian is computable as for our case.
We may consider the particular case
$K=\frac{\kappa_{0}}{\chi^{\prime 2}}\ .$ (36)
In this case $H(\chi)$ vanishes. On the other hand, one obtains in leading order of the saddle point expansion a potential which diverges for $\chi\to 0$
$U=-\varepsilon_{d}\big{(}\ln(\chi)-\frac{1}{2}\ln\kappa_{0}\big{)}\ .$ (37)
We conclude that the possible ultraviolet fixed point (2) is unlikely to arise from a microscopic action (3) with linear measure (6). It could rather be connected to a measure $\sim\text{d}\chi^{\prime}/\chi^{\prime}$ which respects the multiplicative scale symmetry and is linear in ${\varphi}^{\prime}$. In this case one deals with a free massless scalar field ${\varphi}=\sqrt{\kappa_{0}}\ln(\chi/\sqrt{\kappa_{0}})$. For the possible ultraviolet fixed point this should describe the behavior for $\chi\to 0$, ${\varphi}\to-\infty$. A possible crossover to the trivial fixed point for $\chi$ needs an instability towards $\chi\to\infty$, ${\varphi}\to\infty$. A possibility is a measure which is linear in $\chi^{\prime}$ for large $|\chi^{\prime}|$, and $\sim\text{d}\chi^{\prime}/\chi^{\prime}$ for small $|\chi^{\prime}|$.
Field transformations for the effective action
The issue of a change of field-variables is rather different on the level of the (quantum) effective action $\Gamma$. The computation of $\Gamma$ performs already the functional integral. Once the effective action is computed or its form assumed, no further integration over fields is needed and the Jacobian for integration variables plays no longer a role. The quantum field equation and the macroscopic correlation functions are obtained by taking functional derivatives of the effective action. On this level one can perform arbitrary transformations of field variables. Physical observables are often expressed in terms of the macroscopic correlation functions, which in turn are given by functional derivatives of $\Gamma$. The expectation values of observables remain the same if the field transformations are inserted in the functional derivatives of the effective action. Field equations are differential equations. They can be solved for an arbitrary choice of field-variables. In short, the physics remains invariant under an arbitrary change of macroscopic field-variables on which a given effective action depends. We call the transformations of the variables in the effective action “macroscopic field transformations”. What is affected by the macroscopic field transformations is the relation between the macroscopic correlation functions and the microscopic correlation functions defined by expectation values of powers of microscopic fields. We discuss this issue in detail.
Effective action as functional integral
In place of the most commonly used definition of the effective action by the Legendre transform of the logarithm of the partition function in the presence of sources, we employ here an equivalent definition by an implicit functional integral Berges et al. (2002) (background field integral),
$\Gamma[\chi]=$
$-\ln\int\mathcal{D}\chi^{\prime}\,\exp\left\{-S[\chi^{\prime}]+\int_{x}\frac{\partial\Gamma}{\partial\chi(x)}\big{(}\chi^{\prime}(x)-\chi(x)\big{)}\right\}\ .$ (38)
This is an integro-differential equation due to the presence of the sources
$j(x)=\frac{\partial\Gamma}{\partial\chi(x)}\ .$ (39)
The exact quantum field equation in the absence of sources is given by
$\frac{\partial\Gamma}{\partial\chi(x)}=0\ .$ (40)
In the form (III) it is apparent that a change of the integration variable $\chi^{\prime}(x)$ does not change the integral $\Gamma[\chi]$, provided that the Jacobian is properly included.
Taking a functional derivative of eq. (III) with respect to $\chi(x)$ yields the identity
$\int_{y}\Gamma^{(2)}(x,y)\big{(}\langle\chi^{\prime}(y)\rangle-\chi(y)\big{)}=0\ .$ (41)
Here we define the expectation value of an observable in the presence of sources by
$\langle A\rangle=\frac{\int\mathcal{D}\chi^{\prime}\,A[\chi^{\prime}]\exp\big{\{}-S[\chi^{\prime}]+\int_{x}j(x)\big{(}\chi^{\prime}(x)-\chi(x)\big{)}\big{\}}}{\int\mathcal{D}\chi^{\prime}\,\exp\big{\{}-S[\chi^{\prime}]+\int_{x}j(x)\big{(}\chi^{\prime}(x)-\chi(x)\big{)}\big{\}}}\ ,$ (42)
and $\Gamma^{(2)}$ is the second functional derivative of $\Gamma$,
$\Gamma^{(2)}(x,y)=\frac{\partial^{2}\Gamma}{\partial\chi(x)\partial\chi(y)}\ .$ (43)
If $\Gamma^{(2)}$ is invertible, one concludes that the macroscopic field $\chi(x)$ equals the expectation value of the microscopic field $\chi^{\prime}(x)$ in the presence of a suitable source
$\chi(x)=\langle\chi^{\prime}(x)\rangle\ .$ (44)
Taking a derivative of eq. (44) with respect to $\chi(z)$ expresses the connected two-point function $G(y,z)$ as the inverse of $\Gamma^{(2)}$,
$\int_{y}\Gamma^{(2)}(x,y)G(y,z)=\delta(x-z)\ ,$ (45)
with $G$ given by the expectation value (7) in the presence of sources. Taking further functional derivatives of eq. (45) the effective action is found to be the generating functional for the one-particle-irreducible Green functions for the field $\chi^{\prime}(x)$.
In terms of the fluctuations around the macroscopic field,
$\hat{\chi}(x)=\chi^{\prime}(x)-\chi(x)\ ,$ (46)
we can write the effective action in the form
$\Gamma[\chi]=-\ln\int\mathcal{D}\hat{\chi}\,\exp\left\{-S[\chi+\hat{\chi}]+\int_{x}\frac{\partial\Gamma}{\partial\chi(x)}\hat{\chi}(x)\right\}\ .$ (47)
This step assumes that the measure $\mathcal{D}\hat{\chi}$ is defined such that $\mathcal{D}\hat{\chi}=\mathcal{D}\chi^{\prime}$. In this form the saddle point approximation (24) follows as an expansion of the argument of the exponential on the r.h.s. of eq. (47) in powers of $\hat{\chi}$. One inserts in lowest order $\partial\Gamma/\partial\chi(x)=\partial S/\partial\chi(x)$, and performs the Gaussian integral for the term quadratic in $\hat{\chi}$. As before, the loop integrals need a regularization.
A change of the integration variable from $\chi^{\prime}$ to ${\varphi}^{\prime}$ does not change the functional $\Gamma[\chi]$ defined by eq. (III). Of course, for the functional integral the Jacobian has to be taken into account. It becomes apparent why in the preceding section the saddle point expansions for the $\chi$-scheme and the ${\varphi}$-scheme differ. In the ${\varphi}$-scheme the step from eq. (III) to eq. (47) requires to take out the Jacobian from the measure such that $\int\mathcal{D}{\varphi}^{\prime}=\prod_{x}\int\text{d}{\varphi}(x)$ obeys $\mathcal{D}\hat{\varphi}=\mathcal{D}{\varphi}^{\prime}$. Once the Jacobian is included in the classical action, $\bar{S}[{\varphi}^{\prime}]=S[{\varphi}^{\prime}]-\ln J[{\varphi}^{\prime}]$, and we express the source term $\int_{x}\partial\Gamma/\partial\chi)(\chi^{\prime}-\chi)$ in terms of ${\varphi}^{\prime}$, one obtains the same $\Gamma$ for either choice of integration variable. The saddle point expansions in the $\chi$- and ${\varphi}$-schemes start, however, from a different tree approximation $S[\chi^{\prime}]\neq\bar{S}[{\varphi}^{\prime}]$, and also have different loop contributions since $S^{(2)}$ differs from $\bar{S}^{(2)}$.
Transformations of microscopic and macroscopic fields
As stated above, once the effective action is computed one is free to make any macroscopic variable change from $\chi$ to ${\varphi}[\chi]$. The derivative $\partial\Gamma/\partial{\varphi}(x)$ is easily expressed in terms of $\partial\Gamma/\partial\chi(x)$, and similarly for higher functional derivatives as $\partial^{2}\Gamma/\partial{\varphi}(x)\partial{\varphi}(y)$. The relation between functional derivatives and correlation functions for microscopic fields needs some care, however. Let us consider non-linear field transformations of the macroscopic and microscopic fields
${\varphi}(x)=N_{m}\big{(}\chi(x)\big{)}\chi(x)\ ,\quad{\varphi}^{\prime}(x)=N\big{(}\chi^{\prime}(x)\big{)}\chi^{\prime}(x)\ ,$ (48)
with
$\frac{\partial{\varphi}}{\partial\chi}=\tilde{N}_{m}(\chi)\ ,\quad\frac{\partial{\varphi}^{\prime}}{\partial\chi^{\prime}}=\tilde{N}(\chi^{\prime})\ ,\quad\tilde{N}_{m}=N_{m}+\frac{\partial N_{m}}{\partial\ln\chi}\ .$ (49)
Even though we may use the same function $N$ in both expressions in eq. (48), we stress that we deal with different transformations. For a better distinction we use the index $m$ for the macroscopic field transformations for which $N_{m}$ depends on the expectation value $\chi=\langle\chi^{\prime}\rangle$. In contrast, for the microscopic field transformation $N$ is a function of the microscopic field $\chi^{\prime}$.
We may compare the inverse of the second functional derivative of $\Gamma$ with respect to ${\varphi}$ and the connected two point function for ${\varphi}^{\prime}$. With
$\Gamma_{({\varphi})}^{(2)}(x,y)=\frac{\partial^{2}\Gamma}{\partial{\varphi}(x)\partial{\varphi}(y)}$ (50)
$=\tilde{N}_{m}^{-1}(x)\frac{\partial^{2}\Gamma}{\partial\chi(x)\partial\chi(y)}\tilde{N}_{m}^{-1}(y)+\frac{\partial^{2}\chi(y)}{\partial{\varphi}(x)\partial{\varphi}(y)}\frac{\partial\Gamma}{\partial\chi(y)}\ ,$
one has for $\partial\Gamma/\partial\chi(y)=0$
$\big{(}\Gamma_{({\varphi})}^{(2)}\big{)}^{-1}(x,y)$ $=\tilde{N}_{m}(x)G(x,y)\tilde{N}_{m}(y)$ (51)
$=\tilde{N}_{m}\big{(}\langle\chi^{\prime}(x)\rangle\big{)}\tilde{N}_{m}\big{(}\langle\chi^{\prime}(y)\rangle\big{)}$
×$\quad\times\big{(}\langle\chi^{\prime}(x)\chi^{\prime}(y)\rangle-\langle\chi^{\prime}(x)\rangle\langle\chi^{\prime}(y)\rangle\big{)}\ .$
On the other side, the connected two-point function for ${\varphi}^{\prime}$ reads
$G_{({\varphi})}(x,y)=$ $\,\langle{\varphi}^{\prime}(x){\varphi}^{\prime}(y)\rangle-\langle{\varphi}^{\prime}(x)\rangle\langle{\varphi}^{\prime}(y)\rangle$
$=$ $\,\langle N\big{(}\chi^{\prime}(x)\big{)}\chi^{\prime}(x)N\big{(}\chi^{\prime}(y)\big{)}\chi^{\prime}(y)\rangle$
$-\langle N\big{(}\chi^{\prime}(x)\big{)}\chi^{\prime}(x)\rangle\langle N\big{(}\chi^{\prime}(y)\big{)}\chi^{\prime}(y)\rangle\ .$ (52)
This expression involves higher correlation functions for $\chi^{\prime}$. The expressions (51) and (III) differ even if one uses the same functions for $N_{m}$ and $N$. Their explicit relation is not obvious. A further difference arises for $\partial\Gamma/\partial\chi(y)\neq 0$.
In summary, we have defined the effective action $\Gamma_{\chi}[\chi]$ by the implicit relation (III). Here the index $\chi$ of $\Gamma$ indicates the form of the source term linear in $\chi^{\prime}-\chi$. The correlation functions for the microscopic field $\chi^{\prime}$ are simply related to functional derivatives of $\Gamma_{\chi}[\chi]$ with respect to $\chi$. By a macroscopic field transformation $\chi=\chi[{\varphi}]$ we obtain
$\Gamma_{\chi}[{\varphi}]=\Gamma_{\chi}\big{[}\chi[{\varphi}]\big{]}\ .$ (53)
The functional derivatives of $\Gamma_{\chi}[{\varphi}]$ with respect to ${\varphi}$ do not generate the correlation functions for the microscopic field ${\varphi}^{\prime}$ which obtains from $\chi^{\prime}$ by an analogous field transformation. They rather multiply the correlation functions for $\chi^{\prime}$ with suitable functions of the expectation value $\chi$. In this sense $\chi$ is singled out as the “canonical field” for the effective action $\Gamma_{\chi}$. Fields ${\varphi}[\chi]$ obtained by a non-linear field transformation are no longer canonical as far as the relation of functional derivatives of $\Gamma_{\chi}[{\varphi}]$ to correlation functions of microscopic fields are concerned. They may be called “composite fields”.
Macroscopic correlation functions
We may define “macroscopic correlation functions” for ${\varphi}$ by suitable functional derivatives of $\Gamma_{\chi}$ with respect to ${\varphi}(x)$. For example, the macroscopic connected two-point function for ${\varphi}$ is given by the inverse of $\Gamma_{({\varphi})}^{(2)}$ in eq. (50). In contrast, the “microscopic correlation functions” are given by suitable expectation values of powers of the microscopic field ${\varphi}^{\prime}(x)$. For the canonical field $\chi^{\prime}$ the macroscopic and microscopic correlation functions coincide. For the non-canonical field ${\varphi}^{\prime}$ we have seen that in general the macroscopic and microscopic correlation functions differ.
For many problems the knowledge of the microscopic correlation function plays actually no role. Only the macroscopic correlation functions matter, and therefore only suitable macroscopic variable transformations in the effective action $\Gamma_{\chi}$ are involved. As an example, we may consider the field equation in the presence of a source,
$\frac{\partial\Gamma_{\chi}}{\partial\chi(x)}=j(x)\ .$ (54)
This translates directly to the field equation for ${\varphi}(x)$
$\frac{\partial\Gamma_{\chi}}{\partial{\varphi}(x)}=\tilde{N}_{m}^{-1}\big{(}{\varphi}(x)\big{)}j(x)=j_{\varphi}(x)\ .$ (55)
Let us denote the solution of the field equation in the absence of sources by ${\varphi}_{0}(x)=N_{m}\big{(}\chi_{0}(x)\big{)}\chi_{0}(x)$, where $\partial\Gamma/\partial\chi(x)(\chi_{0})=0$. For small $j_{\varphi}(x)$ one can solve the linearized field equation, with solution denoted by $\bar{\varphi}(x)={\varphi}_{0}(x)+\delta{\varphi}(x)$,
$\int_{y}\frac{\partial^{2}\Gamma}{\partial{\varphi}(x)\partial{\varphi}(y)}$ $\delta{\varphi}(y)=j_{\varphi}(x)\ ,$
$\delta{\varphi}(x)=$ $\,\int_{y}\big{(}\Gamma_{\varphi}^{(2)}\big{)}^{-1}(x,y)j_{\varphi}(y)\ .$ (56)
The linearized solution only involves the macroscopic correlation function $(\Gamma_{\varphi}^{(2)})^{-1}$. It is the macroscopic rather than the microscopic correlation function which describes the response of the macroscopic system, as characterized by the value of ${\varphi}$, to a small (macroscopic) source.
The role of the macroscopic correlation functions is particularly apparent for critical systems. Many, sometimes vastly different, microscopic models can lead to the same universal critical behavior which is encoded in the form of the free energy of effective action $\Gamma$. Important quantities as the correlation length can be extracted directly from $\Gamma$ for an arbitrary choice of fields, reflecting properties of the second functional derivative $\Gamma^{(2)}$. The microscopic correlation functions play no role here – they can be actually very different for the different microscopic models. Only the macroscopic fields play a role for the universal critical behavior. In terms of the microscopic fields of a given microscopic model they are often composite fields.
In summary, on the level of the effective action a detailed understanding of field transformations of microscopic variables is only needed for the relation between macroscopic and microscopic correlation functions. As long as only the macroscopic level is concerned we can simply consider non-linear transformations of the macroscopic fields as variable transformations for a given functional and its derivatives. It is the analogue to variable transformations for arguments of a function.
Different effective actions
The choice of the effective action is not unique. Different effective actions can be defined by different source terms. For example, we may specify a different effective action $\Gamma_{\varphi}[{\varphi}]$ by replacing in eq. (III) the source term by $\int_{x}\partial\Gamma/\partial{\varphi}(x)\big{(}{\varphi}^{\prime}(x)-{\varphi}(x)\big{)}$. This corresponds to a canonical source term for the “composite field” ${\varphi}^{\prime}$,
$\Gamma_{\varphi}[{\varphi}]=$
$-\ln\int\mathcal{D}\chi^{\prime}\,\exp\left\{-S[\chi^{\prime}]+\int_{x}\frac{\partial\Gamma_{\varphi}}{\partial{\varphi}(x)}\big{(}{\varphi}^{\prime}(x)-{\varphi}(x)\big{)}\right\}\ .$ (57)
The macroscopic correlation functions for ${\varphi}$ derived from $\Gamma_{\varphi}[{\varphi}]$ equal the microscopic correlation functions for ${\varphi}^{\prime}$. The difference to the previous version of the effective action $\Gamma_{\chi}[{\varphi}]$ arises only from the different source term. (Both $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$ are independent of the choice of the integration variable.) As far as the relation between macroscopic and microscopic correlation functions is concerned the effective action admits a “canonical field choice”. The canonical field is $\chi(x)$ for $\Gamma_{\chi}$ and ${\varphi}(x)$ for $\Gamma_{\varphi}$. Only for the canonical field choice the macroscopic correlation functions equal the microscopic ones.
Effective action for fundamental and composite fields
Further light on this issue can be shed by introducing sources for both $\chi^{\prime}$ and ${\varphi}^{\prime}$ and defining an effective action depending both on the macroscopic fields $\chi$ and ${\varphi}$, treated as independent variables,
$\Gamma[\chi,{\varphi}]=$
$-\ln\int\mathcal{D}\chi^{\prime}\,\exp\bigg{\{}-S[\chi^{\prime}]+\int_{x}\bigg{[}\frac{\partial\Gamma}{\partial\chi(x)}\big{(}\chi^{\prime}(x)-\chi(x)\big{)}$
$\quad+\frac{\partial\Gamma}{\partial{\varphi}(x)}\big{(}N\big{(}\chi^{\prime}(x)\big{)}\chi^{\prime}(x)-{\varphi}(x)\big{)}\bigg{]}\bigg{\}}\ .$ (58)
This results in
$\chi(x)=\langle\chi^{\prime}(x)\rangle\ ,\quad{\varphi}(x)=\langle{\varphi}^{\prime}(x)\rangle=\langle N\big{(}\chi^{\prime}(x)\big{)}\chi^{\prime}(x)\rangle\ ,$ (59)
where expectation values are taken in the presence of the two sources
$\tilde{j}_{\chi}(x)=\frac{\partial\Gamma}{\partial\chi(x)}\ ,\quad\tilde{j}_{\varphi}(x)=\frac{\partial\Gamma}{\partial{\varphi}(x)}\ .$ (60)
Let us denote by ${\varphi}_{0}[\chi]$ the solution of the partial field equation
$\frac{\partial\Gamma[\chi,{\varphi}]}{\partial{\varphi}(x)}\big{(}{\varphi}_{0}[\chi]\big{)}=0\ .$ (61)
This defines a macroscopic field transformation by
${\varphi}_{0}[\chi]=N_{m}(\chi)\chi\ .$ (62)
In general, the function $N_{m}$ differs from $N$ as used in the microscopic field transformation in eqs. (III) (59).
Inserting this $\chi$-dependent solution into eq. (III) we recover
$\Gamma_{\chi}[\chi]=\Gamma\big{[}\chi,{\varphi}_{0}[\chi]\big{]}\ .$ (63)
Defining similarly $\chi_{0}[{\varphi}]$ by the solution of the partial field equation
$\frac{\partial\Gamma[\chi,{\varphi}]}{\partial\chi(x)}\big{(}\chi_{0}[{\varphi}]\big{)}=0\ ,$ (64)
yields $\Gamma_{\varphi}[{\varphi}]$. In general, $\Gamma_{\varphi}[{\varphi}]$ and $\Gamma_{\chi}[{\varphi}]$ differ,
$\Gamma_{\chi}[{\varphi}]$ $=\Gamma\left[\frac{{\varphi}}{N_{m}({\varphi})},{\varphi}_{0}\bigg{[}\frac{{\varphi}}{N_{m}({\varphi})}\bigg{]}\right]\ ,$
$\Gamma_{\varphi}[{\varphi}]$ $=\Gamma\big{[}\chi_{0}[{\varphi}],{\varphi}\big{]}\ .$ (65)
Nevertheless, the two functionals $\Gamma_{\chi}$ and $\Gamma_{\varphi}$ coincide on the solution $(\chi_{0},{\varphi}_{0})$ of the full field equation
$\frac{\partial\Gamma}{\partial\chi}[\chi_{0},{\varphi}_{0}]=0\ ,\quad\frac{\partial\Gamma}{\partial{\varphi}}[\chi_{0},{\varphi}_{0}]=0\ ,$ (66)
where
$\Gamma_{\chi}[{\varphi}_{0}]=\Gamma_{\varphi}[{\varphi}_{0}]=\Gamma[\chi_{0},{\varphi}_{0}]\ .$ (67)
Besides the conceptual clarification an effective action depending both on a field $\chi$ and a “composite” of this field ${\varphi}[\chi]$ is often useful for practical purposes. An example are strong interactions in elementary particle physics where meson fields can be introduced as bilinears in the quark fields.
Frame invariance
For a given effective action $\Gamma_{\chi}$ or $\Gamma_{\varphi}$ the fields $\chi^{\prime}(x)$ or ${\varphi}^{\prime}(x)$ can be considered as different coordinates for field-space, and similarly for $\chi(x)$ or ${\varphi}(x)$. Observables are independent of the coordinate choice or choice of “frame”. This frame independence can be formulated on different levels. Microscopic frame invariance states that observables are independent of the choice of microscopic fields $\chi^{\prime}(x)$ or ${\varphi}^{\prime}(x)$. A given choice of effective action as $\Gamma_{\chi}[\chi]$ or $\Gamma_{\varphi}[{\varphi}]$ is per se invariant under microscopic frame transformations since those only concern the change of the integration variables in an integral. This supposes, of course, that the Jacobian is properly taken into account. All correlation functions and observables that can be computed from the effective action take the same values in all microscopic frames. For this type of microscopic frame transformations the sources are kept fixed, i.e. $j(x)=\partial\Gamma_{\chi}/\partial\chi(x)$ in the functional integral (III) is not changed. As a consequence the source term changes its form if $\chi^{\prime}$ is replaced by $\chi^{\prime}({\varphi}^{\prime})$. A different form of microscopic frame invariance, for which the sources are transformed simultaneously, will be discussed below.
Macroscopic frame transformations concern field transformations of the macroscopic fields $\chi(x)$ or ${\varphi}(x)$. Observables that can be expressed by macroscopic correlation functions can equally be computed from $\Gamma_{\chi}[\chi]$ or $\Gamma_{\chi}[{\varphi}]=\Gamma_{\chi}\big{[}\chi[{\varphi}]\big{]}$. They take the same values provided we replace properly in all functional derivatives
$\frac{\partial}{\partial\chi(x)}=\int_{y}\frac{\partial{\varphi}(y)}{\partial\chi(x)}\frac{\partial}{\partial{\varphi}(y)}\ .$ (68)
This simple statement is macroscopic frame invariance. For invertible field transformations it implies directly that the quantum field equations in the absence of sources can be equally computed from $\Gamma_{\chi}[\chi]$ or $\Gamma_{\chi}[{\varphi}]$,
$\frac{\partial\Gamma_{\chi}[\chi]}{\partial\chi(x)}=0\iff\frac{\partial\Gamma_{\chi}[{\varphi}]}{\partial{\varphi}(x)}=0\ .$ (69)
Macroscopic frame invariance extends to field equations in the presence of sources if the sources are transformed properly according to eq. (55).
Macroscopic frame invariance is relevant for cosmology. In any theory of quantum gravity the relevant field equations are the quantum field equations obtained by variation of the quantum effective action. The relevant fields are the macroscopic metric together with possible additional macroscopic fields as scalar fields. Different “metric frames” correspond to different choices of the macroscopic metric field. They are related by macroscopic frame transformations. An example are the well known Weyl scalings H.Weyl (1918); Dicke (1962) or conformal transformations which multiply the metric by a function of a scalar field. Macroscopic frame invariance states that the quantum field equations derived from variation of the effective action are equivalent for all metric frames. This equivalence of cosmological equations on the quantum level has been advocated for a long time Wetterich (1988).
In contrast, in a quantum context a similar equivalence does not hold for the classical field equations. Two classical actions $S[\chi^{\prime}]$ and $S[{\varphi}^{\prime}]=S\big{[}\chi^{\prime}[{\varphi}^{\prime}]\big{]}$ lead, of course, to the same classical field equations. As a starting point for a quantum theory the two actions are not equivalent, however, unless supplemented by an appropriate Jacobian in the functional measure. This Jacobian leads effectively to different classical actions $S[\chi^{\prime}]$ and $\bar{S}[{\varphi}^{\prime}]$. Even if the Jacobian is included (which is usually not done) the loop expansion in any finite order will lead to different results, as we have seen in section II. In the past, these facts have led to some confusion about the equivalence of different metric frames.
Beyond the quantum field equations, the macroscopic frame invariance also extends to the characterization of fluctuations. The “primordial cosmic fluctuations” in inflationary cosmology correspond to the macroscopic correlation function, with power spectrum directly proportional to the macroscopic connected two-point function in momentum space. It is rather obvious that the metric fluctuations in the observable range of scales correspond to fluctuations of some type of macroscopic metric. The details of microscopic correlation functions for degrees of freedom defined at length scales much shorter than the Planck length play no role - the macroscopic metric could even be a composite of different microscopic fields, as scalar or vector fields. Details of the relation between microscopic and macroscopic correlations could only play a role if they affect the form of the effective action.
The primordial cosmic fluctuation spectrum can indeed be computed by inverting the second functional derivative of the effective action Wetterich (2015a, 2017). The power spectrum of primordial fluctuations can be shown directly to be a frame invariant quantity with respect to Weyl scalings Wetterich (2016). In the context of quantum gravity and cosmology the macroscopic frame invariance has been named “field relativity” Wetterich (2013a, 2015b). Since geometry depends on the choice of the metric, macroscopic frame invariance implies the relativity of geometry. Geometry remains invariant only under the particular metric transformations that correspond to a change of the coordinates of space-time (diffeomorphisms). Under more general metric transformations, as Weyl scalings, the geometry changes.
(In)equivalent source terms
A more subtle issue concerns field transformations accompanied by changes of source terms in the definition of the effective action. They are also often named frame transformations. For this version of microscopic frame transformations the effective action changes due to the change of the source term. While a macroscopic frame transformation changes $\Gamma_{\chi}[\chi]\to\Gamma_{\chi}[{\varphi}]$, this version of a microscopic frame transformation amounts to the map $\Gamma_{\chi}[\chi]\to\Gamma_{\varphi}[{\varphi}]$. It is this version of microscopic frame transformations that will be generalized to the microscopic field transformations of flow equations Pawlowski (2007) which are the basis of the “essential renormalization group” Baldazzi et al. (2022); Baldazzi and Falls (2021).
This type of microscopic frame transformation maps the functional integral (III) for $\Gamma_{\chi}[\chi]$ to the functional integral (III) for $\Gamma_{\varphi}[{\varphi}]$ or
$\Gamma_{\varphi}[{\varphi}]=$ $\,-\ln\int\mathcal{D}{\varphi}^{\prime}\,\exp\bigg{\{}-\bar{S}[{\varphi}^{\prime}]$
$+\int_{x}\frac{\partial\Gamma_{\varphi}}{\partial{\varphi}(x)}\big{(}{\varphi}^{\prime}(x)-{\varphi}(x)\big{)}\bigg{\}}\ .$ (70)
Both microscopic and macroscopic fields are transformed. A change of the integration variable has no effect on the effective action. The change of the macroscopic field is a change of the variables in terms of which the effective action is expressed. The core of this type of microscopic frame transformation is therefore the non-trivial transformation of the source term.
A possible view considers the source terms only as an auxiliary construction and evaluates correlation functions at vanishing sources $j(x)=0$, or equivalently for solutions of the field equations obtained by variation of the effective action. The final result for microscopic correlation functions is then evaluated by a functional integral in the absence of a source term according to eq. (42) with $j(x)=0$. These correlation functions are, of course, independent of the precise definition of the source term, provided that the final result is obtained for a vanishing source term.
On the other hand, we have seen explicitly that the effective actions $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$, which are distinguished by different source terms, are different functionals. They describe different relations between microscopic and macroscopic correlation functions. In principle, the microscopic correlation functions for a given microscopic model defined by the classical action $S[\chi^{\prime}]$ and the functional measure $\mathcal{D}\chi^{\prime}$ can be formally inferred both from $\Gamma_{\chi}[{\varphi}]$ and from $\Gamma_{\varphi}[{\varphi}]$. In practice, the way how they are derived can be rather different, however. The difference of the functionals $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$ has to be compensated by different relations how a given observable or microscopic correlation function in the absence of sources is related to functional derivatives of $\Gamma_{\chi}[{\varphi}]$ or $\Gamma_{\varphi}[{\varphi}]$. While this relation exists formally, it is typically much too complicated to be of practical use. Whenever one wants to exploit the behavior of the effective action away from the solution of the quantum field equations one has to consider $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$ as different functionals. Only the field equations in the absence of sources, and the values of the effective action for solutions of the field equations are identical for $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$.
These remarks are relevant for the macroscopic correlation functions evaluated on solutions of the field equations. The second functional derivatives of $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$ differ, as we have demonstrated explicitly. The question arises which effective action should be used for the definition of macroscopic correlation functions. In principle, this is determined by the question which macroscopic correlation functions are relevant for a given observation.
For many questions the difference between $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$ plays no role for real observations in practice. This is related to the fact that observations typically concern length scales much larger than the microscopic length scales on which the theory is defined. Universality induced by the renormalization flow washes out differences in the microscopic definition. In the momentum range relevant for observations $\Gamma_{\chi}[{\varphi}]$ and $\Gamma_{\varphi}[{\varphi}]$ are typically the same, or distinguished only by somewhat different values of renormalizable couplings. One can then use $\Gamma_{\chi}[{\varphi}]$ or $\Gamma_{\varphi}[{\varphi}]$ equivalently for the computation of macroscopic correlation functions.
The main conclusions of this section state that macroscopic field transformations (frame transformations) for the arguments of the quantum effective action do not affect the physics. Macroscopic correlation functions obtain by functional derivatives of the effective action and transform covariantly. They are relevant for many situations of interest. Only if one wants to establish a relation between the macroscopic correlation functions and the microscopic correlation functions a canonical choice of fields (canonical frame) is singled out.
Different versions of the quantum effective action correspond to different choices of source terms. They are related by a type of microscopic frame transformation which also transforms the source term. These different versions are strictly equivalent as far as solutions of the quantum field equations in the absence of sources are concerned. They often remain (approximately) equivalent for the macroscopic correlation functions as well.
Field transformations in flow equation
A conceptually, and often also practically, convenient way to compute the effective action employs a functional flow equation. In this approach the fluctuation effects are included stepwise by variation of a suitable cutoff. The change of the effective action for an infinitesimal step typically involves a momentum integration which covers effectively only a finite momentum range. As a consequence, the loop integrals are finite and no further regularization is needed. The cutoff involves a characteristic momentum scale $k$. The change of the effective action by a variation of $k$ obeys an exact functional flow equation Wetterich (1993b). Conceptually, one starts at very high $k$ (or infinite $k$) with the microscopic effective action. This should be related to the “classical action” which defines the functional integral. Following the flow equations to $k\to 0$ yields the quantum effective action which includes all fluctuation effects. In practice, a truncation of the form of the effective action is needed in order to obtain approximate solutions of the functional flow equation.
Field transformations for the flow equation can be implemented in several ways. “Macroscopic field transformations” Wetterich (1996) amount to variable transformations for a given flow equation. They correspond to a simple change of variables for a differential equation. In contrast, “microscopic field transformations” Pawlowski (2007) change the cutoff and the source term used for the definition of the flow equation. A microscopic field transformation therefore corresponds to a map onto a different flow equation which is not related to the original flow equation by a simple change of variables. Microscopic variable transformations also affect the relation between the classical action and the microscopic effective action. This brings back indirectly the issue of the Jacobian in the functional integral. Despite their different conceptual status the macroscopic and microscopic variable transformations have in practice often very similar effects. They are equivalent for linear field transformations. We focus in this section on the macroscopic field transformations and discuss the microscopic field transformations in sects. VVI.
We demonstrate many of the issues encountered for field transformations by a particular example. This concerns the flow of the field-dependent coefficient of the derivative term (kinetial) in a simple scalar field theory. For comparison, we present in two appendices A and B the flow equation for the kinetial without performing a field transformation. In appendix A we omit the effect of the scalar potential on the flow of the kinetial. The contributions of the potential are included in appendix B. These computations can serve as a benchmark for comparison with the results obtained by field transformations in the flow equation. We will see that the flow equations are not changed by $k$-independent field transformations. The computations of the appendices AB therefore cover this case as well. We find strong indications for the existence of the ultraviolet fixed points (2) from these computations.
Effective average action and flow equation
We first present an intuitive definition of the effective average action $\Gamma_{k}$ in complete analogy to eq. (47), adding only a cutoff term for the fluctuations
[…]
$-\frac{1}{2}\int_{x,y}$ $\hat{\chi}(x)\mathcal{R}_{k}(x,y)\hat{\chi}(y)\bigg{\}}\ .$ (71)
The term $\sim\mathcal{R}_{k}$ serves as a cutoff, typically an infrared cutoff which suppresses fluctuations with $q^{2}<k^{2}$. It introduces the renormalization scale $k$. We assume $\mathcal{R}_{k=0}=0$, such that for $k=0$ the effective average action becomes the quantum effective action $\Gamma_{\chi}[\chi]$ defined by eq. (47).
While intuitive, the relation (IV) is not the basic definition of the effective average action. The basic definition of $\Gamma_{k}$ is similar to eq. (III),
$\Gamma_{k}[\chi]=$ $\,\Gamma_{k}^{\prime}[\chi]-\Delta_{k}S[\chi]\ ,$
$\Gamma_{k}^{\prime}[\chi]=$ $\,-\ln\int\mathcal{D}\chi^{\prime}\,\exp\bigg{\{}-S[\chi^{\prime}]-\Delta_{k}S[\chi^{\prime}]$
$\quad\quad+\int_{x}\frac{\partial\Gamma_{k}^{\prime}}{\partial\chi(x)}\big{(}\chi^{\prime}(x)-\chi(x)\big{)}\bigg{\}}\ ,$ (72)
where the cutoff term $\Delta_{k}S$ is quadratic in $\chi^{\prime}$
$\Delta_{k}S[\chi^{\prime}]=\frac{1}{2}\int_{x,y}\chi^{\prime}(x)\mathcal{R}_{k}(x,y)\chi^{\prime}(y)\ .$ (73)
Here $\Gamma_{k}^{\prime}$ is equivalent to the Legendre transform of the Schwinger functional for a microscopic action including the cutoff term. Subtracting this cutoff term is important for making contact to the microscopic action and to implement the simple one-loop form of the exact flow equation.
The definitions (IV) and (IV) are identical only for $\mathcal{D}\hat{\chi}=\mathcal{D}\chi^{\prime}$. In this case they are related by a simple shift of integration variables in the exponent, employing with $\mathcal{R}_{k}(x,y)=\mathcal{R}_{k}(y,x)$ the relation
$\frac{\partial\Gamma_{k}^{\prime}}{\partial\chi(x)}=\frac{\partial\Gamma_{k}}{\partial\chi(x)}+\int_{y}\mathcal{R}(x,y)\chi(y)\ .$ (74)
For non-linear functional measures $\mathcal{D}\hat{\chi}$ differs from $\mathcal{D}\chi$ and the relation (IV) is no longer valid.
The relation (IV) is a central tool for establishing that the microscopic effective action $\Gamma_{k\to\infty}$ equals the classical action. This follows directly if $\mathcal{R}_{k}$ diverges for $k\to\infty$. The linearity of the functional measure does no longer hold after a non-linear microscopic field transformation. As a consequence, the microscopic effective action and the classical action are no longer identical. We will see in sect. VI that their relation brings back the Jacobian even if we do not change the functional measure. It is the linearity of the functional measure, together with the linearity of the source term, that singles out a canonical choice of fields for the flow equation.
The definition (IV) continues to hold even if $\mathcal{D}\hat{\chi}=\mathcal{D}\chi^{\prime}$ is not realized. For $k=0$ one recovers $\Gamma_{k}[\chi]=\Gamma_{k}^{\prime}[\chi]=\Gamma_{\chi}[\chi]$, as defined by eq. (III). Observing that $\Gamma_{k}^{\prime}$ is the effective action for a modified classical action $S+\Delta_{k}S$ we obtain the relations
$\langle\chi^{\prime}(x)\rangle=\chi(x)\ ,\quad\langle\hat{\chi}(x)\rangle=0\ ,$ (75)
and
$G(x,y)=$ $\,\langle\hat{\chi}(x)\hat{\chi}(y)\rangle=\big{(}\Gamma_{k}^{\prime(2)}\big{)}^{-1}(x,y)$
$=$ $\,\big{(}\Gamma_{k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}(x,y)\ .$ (76)
Here the expectation values are evaluated for the classical action $S+\Delta_{k}S$ in the presence of the source $\partial\Gamma^{\prime}/\partial\chi(x)$. For canonical fields one can use equivalently the weight factor given by the exponential in eq. (IV).
An exact functional flow equation is obtained by taking a $k$-derivative of eq. (IV),
$\partial_{k}\Gamma_{k}=\frac{1}{2}\int_{x,y}\partial_{k}\mathcal{R}_{k}(x,y)\langle\hat{\chi}(y)\hat{\chi}(x)\rangle-\partial_{k}\int_{x}\frac{\partial\Gamma_{k}}{\partial\chi(x)}\langle\hat{\chi}(x)\rangle\ .$ (77)
Inserting the relations (75), (IV) one obtains
$\partial_{k}\Gamma_{k}=\frac{1}{2}\mathrm{tr}\left\{\partial_{k}\mathcal{R}_{k}\big{(}\Gamma_{k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}\right\}\ .$ (78)
Expressing the trace in momentum space and choosing diagonal $\mathcal{R}_{k}$,
$\mathcal{R}_{k}(q,q^{\prime})=$ $\,R_{k}(q^{2})\delta(q,q^{\prime})\ ,$
$\delta(q,q^{\prime})=$ $\,(2\pi)^{d}\prod_{\mu}\delta(q_{\mu}-q^{\prime}_{\mu})\ ,$ (79)
yields
$\partial_{k}\Gamma_{k}=\frac{1}{2}\int_{q}\partial_{k}R_{k}(q^{2})\big{(}\Gamma_{k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}(q,q)\ ,$ (80)
where
$\int_{q}=\frac{1}{(2\pi)^{d}}\prod_{\mu}\int_{-\infty}^{\infty}\text{d}q_{\mu}\ .$ (81)
The single momentum integral reflects the one-loop form of the exact flow equation. The effects of higher loops arise from the non-trivial field- and momentum dependence of $\Gamma_{k}^{(2)}$.
For the example of diagonal $\Gamma_{k}^{(2)}$,
$\Gamma_{k}^{(2)}(q,q^{\prime})=\big{(}q^{2}+m^{2}\big{)}\delta(q,q^{\prime})\ ,$ (82)
the flow equation becomes
$\partial_{k}\Gamma_{k}=\frac{1}{2}\int_{x}\int_{q}\partial_{k}R_{k}(q^{2})\big{(}q^{2}+m^{2}+R_{k}(q^{2})\big{)}^{-1}\ ,$ (83)
where we employ with total volume of space $\Omega$
$\delta(q,q)=\Omega=\int_{x}\ .$ (84)
The diagonal form (82) typically obtains for $x$-independent macroscopic fields $\chi(x)=\chi$, with $m^{2}$ depending on $\chi$. Constant $\chi$ yields the part of $\Gamma_{k}$ which corresponds to the effective potential, $\Gamma_{k}=\int_{x}U_{k}(\chi)+\ldots$. The momentum integral (83) is ultraviolet finite if $R_{k}$ decays sufficiently fast for $q^{2}\to\infty$, and infrared finite for a suitable infrared cutoff $R_{k}(q^{2})$.
The finiteness of the flow equation generalizes to arbitrary $x$-dependent macroscopic fields $\chi(x)$. The r.h.s. of the flow equation (78) is well defined and needs no regularization. In this sense the flow equation constitutes its own “ERGE-regularization scheme”. The regularization conditions in the loop expansion are replaced by the initial value of $\Gamma_{k}$ for very large $k$, e.g. $k\to\infty$ or $k=\Lambda$, with $\Lambda$ some intrinsic UV-cutoff scale as for lattice models $\Lambda\sim\varepsilon^{-1}$.
Macroscopic field transformations
The flow equation is a functional differential equation. As such, we can solve it for an arbitrary choice of field variables. Often it is useful to employ $k$-dependent field variables
${\varphi}={\varphi}_{k}[\chi]\ .$ (85)
This corresponds to a macroscopic $k$-dependent field transformation in the effective average action. Evaluating the flow equation at fixed ${\varphi}$ instead of fixed $\chi$ induces a modification of the flow equation Wetterich (1996),
$\partial_{t}\Gamma_{k}|_{\varphi}=\partial_{t}\Gamma_{k}|_{\chi}-\int_{x}\frac{\partial\Gamma_{k}}{\partial{\varphi}(x)}\partial_{t}{\varphi}(x)|_{\chi}\ .$ (86)
Here we use the flow variable $t=\ln(k/k_{0})$, $\partial_{t}=k\partial_{k}$, and
$\partial_{t}{\varphi}(x)|_{\chi}=k\frac{\partial{\varphi}_{k}[\chi]}{\partial k}(x)=\tilde{\gamma}_{k}[{\varphi}]\ .$ (87)
The last identity expresses $\partial_{t}{\varphi}|_{\chi}$ as a functional of ${\varphi}$. The original flow equation (78) obtains for canonical fields. For composite fields depending on $k$ an additional term appears in eq. (86).
A simple example concerns the dimensionless “scale invariant scalar fields” in four dimensions,
${\varphi}(x)=\tilde{\chi}(x)=\frac{\chi(x)}{k}\ ,\quad\partial_{t}{\varphi}(x)|_{\chi}=-{\varphi}(x)\ ,$ (88)
such that a “scaling form” of the flow equation reads
$\bigg{(}\partial_{t}-\int_{x}\tilde{\chi}(x)$ $\frac{\partial}{\partial\tilde{\chi}(x)}\bigg{)}\Gamma_{k}$
$=\frac{1}{2}\mathrm{tr}\left\{\big{(}2+\partial_{t}\big{)}\tilde{r}_{k}\big{(}\tilde{\Gamma}_{k}^{(2)}+\tilde{r}_{k}\big{)}^{-1}\right\}\ ,$ (89)
where $\Gamma_{k}=\Gamma_{k}[\tilde{\chi}]$ and
$\Gamma_{k}^{(2)}=k^{2}\tilde{\Gamma}_{k}^{(2)}\ ,\quad\mathcal{R}_{k}=k^{2}\tilde{r}_{k}\ .$ (90)
In close analogy we can define renormalized scalar fields by factoring out a $k$-dependent wave function renormalization.
A second example involves momentum dependent rescalings of the Fourier modes $\chi(q)$ in momentum space
${\varphi}(q)=N_{m}\left(\frac{q^{2}}{k^{2}}\right)\chi(q)=N_{m}(y)\chi(q)\ ,\quad y=\frac{q^{2}}{k^{2}}\ .$ (91)
With
$\partial_{t}{\varphi}(q)|_{\chi}=-2\frac{\partial\ln N_{m}}{\partial\ln y}{\varphi}(q)\ ,$ (92)
the flow equation for ${\varphi}$ reads
$\left(\partial_{t}+2\int_{q}\frac{\partial\ln N_{m}}{\partial\ln y}(q){\varphi}(q)\frac{\partial}{\partial{\varphi}(q)}\right)\Gamma_{k}[{\varphi}]$
$=\frac{1}{2}\int_{q}\left(\partial_{t}-4\frac{\partial\ln N_{m}}{\partial\ln y}\right)\tilde{R}_{k}(q)\big{(}\tilde{\Gamma}_{k}^{(2)}+\tilde{\mathcal{R}}_{k}\big{)}^{-1}(q,q)\ ,$ (93)
where now
$\tilde{\Gamma}_{k}^{(2)}(q,q^{\prime})=$ $\,\frac{\partial^{2}\Gamma_{k}}{\partial{\varphi}(q)\partial{\varphi}(q^{\prime})}=N_{m}^{-1}(q)N_{m}^{-1}(q^{\prime})\Gamma_{k}^{(2)}(q,q^{\prime})\ ,$
~$\tilde{\mathcal{R}}_{k}(q,q^{\prime})=$ $\,N_{m}^{-1}(q)N_{m}^{-1}(q^{\prime})R_{k}(q^{2})\delta(q,q^{\prime})\ .$ (94)
Suitable forms of $N_{m}(q)$ may be used in order to absorb some of the momentum dependence of $\Gamma_{k}^{(2)}(q,q^{\prime})$ into a simple $\tilde{\Gamma}_{k}^{(2)}(q,q^{\prime})$. We emphasize, however, that vertices corresponding to derivatives with respect to ${\varphi}(q)$ receive now additional factors of $N_{m}(q)$. The momentum dependence of the propagator is shifted to the vertices.
For linear macroscopic field transformations we can adapt the cutoff function such that it becomes simple in terms of the composite fields. For our example ~$\tilde{\mathcal{R}}_{k}$ can be chosen to be a simple function of $q^{2}$. A linear field transformation does not change the quadratic form of the cutoff term in the functional integral and does not induce a field-dependent Jacobian. We will see in the next section that eq. (IV) obtains equivalently from a microscopic field transformation.
Non-linear field transformations
For a non-linear macroscopic field transformation the modified flow equation shows new more complex features. This is due to the fact that $\partial_{t}\Gamma_{k}|_{\chi}$ has to be expressed in terms of ${\varphi}$. Similarly, the quantity $\tilde{\gamma}_{k}[{\varphi}]$ may not always remain simple. In particular, the flow equation involves now a connection term. For
$\frac{\partial}{\partial\chi(x)}=\int_{u}\tilde{N}_{m}(x,u)\frac{\partial}{\partial{\varphi}(u)}\ ,\quad\tilde{N}_{m}(x,u)=\frac{\partial{\varphi}(u)}{\partial\chi(x)}\ ,$ (95)
the second derivative of $\Gamma$ with respect to $\chi$ results in a type of covariant derivative with respect to ${\varphi}$,
$\frac{\partial^{2}\Gamma_{k}}{\partial\chi(x)\partial\chi(y)}=\int_{u,v}\tilde{N}_{m}(x,u)\tilde{N}_{m}(y,v)$
×$\quad\quad\times\left[\frac{\partial^{2}\Gamma_{k}}{\partial{\varphi}(u)\partial{\varphi}(v)}+\int_{z}C(u,v,z)\frac{\partial\Gamma_{k}}{\partial{\varphi}(z)}\right]\ ,$ (96)
with “connection”
$C(u,v,z)=\int_{s,w}\frac{\partial\chi(s)}{\partial{\varphi}(u)}\frac{\partial\chi(w)}{\partial{\varphi}(v)}\frac{\partial^{2}{\varphi}(z)}{\partial\chi(s)\partial\chi(w)}\ .$ (97)
This connection results in a correction term,
$\Delta(u,v)=\int_{z}C(u,v,z)\frac{\partial\Gamma_{k}}{\partial{\varphi}(z)}\ ,$ (98)
which appears in the flow equation (86) since $\partial_{t}\Gamma|_{\chi}$ involves $\partial^{2}\Gamma/\big{(}\partial\chi(x)\partial\chi(y)\big{)}$.
For a local field transformation, $\partial{\varphi}(x)/\partial\chi(y)=\tilde{N}_{m}(y)\delta(x-y)$, the connection reads
$C(u,v,z)=$ $\,-\frac{\partial\tilde{N}_{m}^{-1}}{\partial\chi}(u)\delta(u-v)\delta(u-z)\ ,$
$\Delta(u,v)=$ $\,-\frac{\partial\tilde{N}_{m}^{-1}}{\partial\chi}(u)\frac{\partial\Gamma_{k}}{\partial{\varphi}(u)}\delta(u-v)\ .$ (99)
With
~$\tilde{\mathcal{R}}_{k}(u,v)=\int_{s,w}\tilde{N}_{m}^{-1}(u,s)\tilde{N}_{m}^{-1}(v,w)\mathcal{R}_{k}(s,w)$ (100)
one obtains the flow equation
$\partial_{t}\Gamma_{\chi,k}[{\varphi}]=\frac{1}{2}\mathrm{tr}$$\left\{\tilde{N}_{m}^{-1}\partial_{t}\mathcal{R}_{k}\big{(}\tilde{N}_{m}^{-1}\big{)}^{T}\big{(}\Gamma_{({\varphi})}^{(2)}+\Delta+\tilde{\mathcal{R}}_{k}\big{)}^{-1}\right\}$
$-\int_{x}\frac{\partial\Gamma_{k}}{\partial{\varphi}(x)}\partial_{t}{\varphi}(x)|_{\chi}\ .$ (101)
The second functional derivative with respect to ${\varphi}$, $\Gamma_{({\varphi})}^{(2)}(x,y)=\partial^{2}\Gamma_{k}/\partial{\varphi}(x)\partial{\varphi}(y)$, is effectively regulated by ~$\tilde{\mathcal{R}}_{k}$ which becomes field-dependent. Even if one finds a field transformation which brings $\Gamma_{({\varphi})}^{(2)}+\Delta$ to a simple form, there remains an additional complication from the field dependence of ~$\tilde{\mathcal{R}}_{k}$. The cutoff acts no longer homogeneously for all field values.
The field transformation discussed in this section only involves the macroscopic fields. The effective average action and the flow equation is not changed, they are only expressed in terms of different variables. The same holds for the field equations that follow in the absence of sources, $\partial\Gamma_{k}/\partial\chi(x)=0$ or $\partial\Gamma_{k}/\partial{\varphi}(x)=0$. These differential equations are simply expressed in different variables. For a non-linear field transformation there is no direct relation between the “composite macroscopic field” ${\varphi}$ and the expectation value of some microscopic composite field. For ${\varphi}(x)=f\big{(}\chi(x)\big{)}$ the relation ${\varphi}(x)=\big{\langle}f\big{(}\chi^{\prime}(x)\big{)}\big{\rangle}$ does not hold, except for linear functions $f$. In this sense the canonical field $\chi(x)$ is singled out by the relation $\chi(x)=\langle\chi^{\prime}(x)\rangle$.
Elimination of couplings
Field transformations of the macroscopic fields can be used to bring either the flow equation or the field equation into a form easier to solve approximately or simpler to understand. An example for the transformation of the flow equation is the elimination of certain couplings Gies and Wetterich (2002, 2004); Braun (2009); Baier et al. (2005); Salmhofer (2007) or simplification of coupling functions Friederich et al. (2011); Denz et al. (2020). A given flowing coupling is typically defined by some functional derivative of $\Gamma_{k}$. Assume a setting where this coupling vanishes for the “microscopic effective action” $\Gamma_{k=\Lambda}$ with very large (or infinite) $\Lambda$. The flow of this coupling obtains by a suitable functional derivative of the flow equation (IV). In general, the coupling will be generated by the flow. One can now choose a field transformation defined by the property that the flow of this particular coupling vanishes. This poses a constraint on the choice of ${\varphi}_{k}[\chi]$. This procedure is useful provided that for this choice of ${\varphi}_{k}[\chi]$ no other important unwanted couplings are generated.
As an example we consider a truncation of the effective average action similar to eq. (3)
$\Gamma_{\chi,k}[\chi]=\frac{1}{2}\int_{x}K(\chi)\partial^{\mu}\chi\partial_{\mu}\chi\ ,$ (102)
where $K(\chi)$ is a $k$-dependent function of $\chi$. With the non-linear field transformation ${\varphi}[\chi]$ defined by
$\frac{\partial{\varphi}}{\partial\chi}=\tilde{N}_{m}=K^{1/2}\ ,\quad\Gamma_{\chi,k}[{\varphi}]=\frac{1}{2}\int_{x}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}\ ,$ (103)
the effective average action is the one for a free massless scalar field. As long as $K_{k}(\chi)$ is positive for all $\chi$ this field transformation exists for every $k$. Naively, one could conclude that fluctuations of a free field do not generate interactions or a scalar potential and that for every $k$ we deal with a free theory for composite fields.
In order to see that this is not correct in general, we investigate the flow equation with a suitable macroscopic variable transformation which keeps the kinetic term (103) fixed for all $k$. This will also fix the $k$-dependence of $K_{k}(\chi)$. For the modified flow equation (IV) one employs
$\Gamma_{({\varphi})}^{(2)}=$ $\,-\partial^{\mu}\partial_{\mu}=-\partial^{2}\ ,\quad\frac{\partial\Gamma_{k}}{\partial{\varphi}(x)}=-\partial^{\mu}\partial_{\mu}{\varphi}(x)\ ,$
$\Delta=$ $\,-\frac{\partial K^{-1/2}}{\partial\chi}\partial^{\mu}\partial_{\mu}{\varphi}\ .$ (104)
With
${\varphi}=N_{m}\chi\ ,\quad\partial_{t}{\varphi}=\partial_{t}N_{m}\chi=\partial_{t}\ln N_{m}{\varphi}\ ,$ (105)
the last term in eq. (IV) becomes
$-\int_{x}\frac{\partial\Gamma}{\partial{\varphi}(x)}\partial_{t}{\varphi}(x)=$ $\,-\int_{x}A({\varphi})\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}\ ,$
$A({\varphi})=$ $\,\frac{\partial}{\partial{\varphi}}\big{(}\partial_{t}N_{m}\chi\big{)}\ ,$ (106)
where $\partial_{t}N_{m}$ is evaluated for fixed $\chi$.
Extracting from a derivative expansion of the first term on the r.h.s. of eq. (IV) the coefficient $B({\varphi})$ of the term with two derivatives yields
$\partial_{t}\Gamma[{\varphi}]=\int_{x}\big{[}B({\varphi})-A({\varphi})\big{]}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}+\dots$ (107)
One can now determine $\partial_{t}(N_{m}\chi)$ such that $A({\varphi})=B({\varphi})$,
$\frac{\partial}{\partial{\varphi}}\partial_{t}(N_{m}\chi)=$ $\,B({\varphi})\ ,$
$\partial_{t}N_{m}(k,\chi)=$ $\,\frac{1}{\chi}\int_{\chi_{0}}^{\chi}\text{d}\chi^{\prime}\left(N_{m}(\chi^{\prime})+\frac{\partial N_{m}(\chi^{\prime})}{\partial\ln\chi^{\prime}}\right)B(\chi^{\prime})\ .$ (108)
In this case the canonical kinetic term (103) for ${\varphi}$ remains preserved. In other words, possible couplings $\sim{\varphi}^{2}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}$ etc. are eliminated by this $k$-dependent macroscopic field transformation.
Nevertheless, the flow equation with initial condition (102) for $k=\Lambda$ has not a free massless scalar field as a solution. The dots in eq. (107) indicate terms in the effective average action with zero or more than two derivatives. These terms are generated by the flow even if not present in the microscopic effective action. For example, the flow equation for a non-trivial potential obtains by evaluating eq. (IV) for homogeneous ${\varphi}$. If not present at $\Lambda$, this flow generates a potential for $k<\Lambda$. For a non-linear field transformation the r.h.s. of eq. (IV) does not vanish due to the ${\varphi}$-dependence of $\tilde{N}_{m}$.
If one has found $\partial_{t}N_{m}(k,\chi)$ as a solution of eq. (IV) one can infer for the flow of the kinetial $K(\chi)$ the relation
$\partial_{t}K^{1/2}(k,\chi)=\left(1+\frac{\partial}{\partial\ln\chi}\right)\partial_{t}N_{m}(k,\chi)\ .$ (109)
This can be compared with the flow of the kinetial computed without a field transformation in appendix A. Without a truncation the two computations should yield identical results for $K(k,\chi)$. The effective actions $\Gamma_{\chi,k}[\chi]$ and $\Gamma_{\chi,k}[{\varphi}]$ are related by a simple variable transformation. This $k$-dependent macroscopic field transformation also maps the flow equation for $\partial_{t}\Gamma|_{\chi}$ and $\partial_{t}\Gamma|_{\varphi}$ into each other. If a kinetial $K(k,\chi)$ obeys the flow equation at fixed $\chi$, the field transformation is constructed such that the kinetic term for ${\varphi}$ takes for all $k$ the simple form (103). The kinetial in the ${\varphi}$-scheme does not change. For the truncation (102) this is precisely the condition (IV). For the present example the direct method of appendix A seems actually to be much simpler than the use of a field transformation.
It is sometimes argued that couplings contained in the kinetial $K(\chi)$ are “inessential” since they can be removed by a field transformation which leads to a canonical kinetic term. Field transformations only move, however, the information contained in the couplings from one place to another. Only in the truncation (102) the precise form of $K(\chi)$ does not matter since this action describes a free massless composite scalar field ${\varphi}$ for arbitrary monotonic $K(\chi)$. As soon as one computes the flow of other parts in the effective action, in particular the flow of the potential, the information contained in $K(\chi)$ becomes important. In appendix B we demonstrate explicitly the importance of the potential for the flow of the kinetial $K$. One concludes that the omission of the dots in eq. (107) can be strongly misleading. We are aware that omitting the potential is a rather drastic truncation. Our example is thought to underline, however, a general issue for field transformations in truncated flow equations. One has to ensure that the field transformation does not move important information into parts that are not covered by the truncation.
Flow equations with cutoff and sources for composite fields
In this section we investigate a different form of field-transformations in flow equations. In this version the variable transformation is performed on the level of the precise definition of the effective average action by choosing a different source term and a different form of the infrared cutoff, now adapted to composite fields. Since the field transformation leads to a different functional integral we call it a microscopic field transformation, in distinction to the macroscopic field transformation which only changes the arguments of the effective average action, as discussed in the preceding section. Microscopic field transformations are closely related to the possibility of defining different effective actions, as $\Gamma_{\chi}$ or $\Gamma_{\varphi}$ in sect. III. They extend this concept to non-zero $k$. We will see in sect. VI that for the microscopic field transformation the Jacobian reappears in the initial value for the flow equations at very high $k$. This happens independently of a change of variables for the functional measure or not. The appearance of the Jacobian underlines the close connection of the microscopic field transformations to the microscopic field transformations in the functional integral in sect. II. The microscopic field transformations discussed in this section have been derived in ref. Pawlowski (2007). They correspond to the ones employed for the “essential renormalization group” Baldazzi et al. (2022); Baldazzi and Falls (2021)
Effective average action for composite fields
As an alternative to the effective average action $\Gamma_{\chi,k}$ defined in the preceding section, we define a different version of the effective average action $\Gamma_{{\varphi},k}$ by taking the cutoff piece quadratic in the fields ${\varphi}^{\prime}_{k}[\chi^{\prime}]$ and using a source given by $\partial\Gamma^{\prime}_{{\varphi},k}/\partial{\varphi}(x)$,
$\Gamma^{\prime}_{{\varphi},k}[{\varphi}]=-\ln\int\mathcal{D}\chi^{\prime}\exp\bigg{\{}-S[\chi^{\prime}]$ (110)
$-\frac{1}{2}\int_{x,y}{\varphi}^{\prime}_{k}(x)\mathcal{R}_{k}(x,y){\varphi}^{\prime}_{k}(y)+\int_{x}\frac{\partial\Gamma_{{\varphi},k}^{\prime}}{\partial{\varphi}(x)}\big{(}{\varphi}^{\prime}_{k}(x)-{\varphi}(x)\big{)}\bigg{\}}\ .$
Here ${\varphi}^{\prime}_{k}(x)$ is a functional of $\chi^{\prime}$, ${\varphi}^{\prime}_{k}[\chi^{\prime}](x)$ which can depend on $k$. For the effective action we subtract again the cutoff piece
$\Gamma_{{\varphi},k}[{\varphi}]=\Gamma^{\prime}_{{\varphi},k}[{\varphi}]-\frac{1}{2}\int_{x,y}{\varphi}(x)\mathcal{R}_{k}(x,y){\varphi}(y)\ .$ (111)
For $k=0$ and $\mathcal{R}_{k=0}=0$ this effective average action becomes the quantum effective action $\Gamma_{\varphi}[{\varphi}]$, as defined by eq. (III). As discussed in section III, the effective actions $\Gamma_{\varphi}[{\varphi}]$ and $\Gamma_{\chi}[{\varphi}]$ are not identical, due to the different form of the source term. Thus $\Gamma_{\chi,k}$ and $\Gamma_{{\varphi},k}$ differ for all $k$, including $k=0$. An exception are the values for solutions of the field equations which coincide for $k=0$.
We emphasize that the definition (110), (111) keeps the original measure $\mathcal{D}\chi^{\prime}$ and the classical action $S[\chi^{\prime}]$ unchanged. There is no need to switch to $k$-dependent integration variables ${\varphi}^{\prime}_{k}(x)$, which would introduce a $k$-dependent Jacobian and make the classical action $k$-dependent once expressed as a functional of ${\varphi}^{\prime}_{k}$. We could, however, use different integration variables ${\varphi}^{\prime}$. In this form $\Gamma_{{\varphi},k}$ takes a form analogous to eq. (IV) provided we include the Jacobian by replacing the microscopic action $S[\chi^{\prime}]$ by $\bar{S}[{\varphi}^{\prime}]$. This gives a hint why the Jacobian will be back for the initial conditions of $\Gamma_{{\varphi},k}$ even if we do not change the integration variable. With the integration variable ${\varphi}^{\prime}$ the flow equation evaluated at fixed ${\varphi}$ will receive an additional contribution from the $k$-dependence of the Jacobian or $\bar{S}[{\varphi}^{\prime}]$. On the other hand, keeping the original integration variables $\chi^{\prime}$ the flow equation will receive additional contributions from the $k$-dependence of the variable transformation. The two views are equivalent.
Taking the functional derivative of $\Gamma_{{\varphi},k}$ with respect to ${\varphi}(x)$ establishes similar to eqs. (41), (44) the identity
$\langle{\varphi}^{\prime}(x)\rangle={\varphi}(x)\ ,$ (112)
where the expectation value is taken in the presence of the source $\partial\Gamma^{\prime}_{{\varphi},k}/\partial{\varphi}(x)$ and the cutoff term $\sim\mathcal{R}_{k}$. Similarly, the derivative of eq. (112) with respect to ${\varphi}(y)$ yields an identity corresponding to eq. (45)
$\int_{y}\frac{\partial^{2}\Gamma_{{\varphi},k}^{\prime}}{\partial{\varphi}(x)\partial{\varphi}(y)}\big{(}\langle{\varphi}^{\prime}_{k}(y){\varphi}^{\prime}_{k}(z)\rangle-{\varphi}(y){\varphi}(z)\big{)}=\delta(x-z)\ .$ (113)
Flow equation for “composite” cutoff and sources
The flow equation receives an additional contribution from the $k$-dependence of ${\varphi}^{\prime}_{k}[\chi]$,
$\partial_{k}\Gamma_{{\varphi},k}[{\varphi}]=$ $\,\frac{1}{2}\mathrm{tr}\Big{\{}\partial_{k}\mathcal{R}_{k}\big{(}\Gamma_{{\varphi},k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}\Big{\}}$
$-\int_{x}\frac{\partial\Gamma_{{\varphi},k}}{\partial{\varphi}(x)}\langle\partial_{k}{\varphi}^{\prime}_{k}(x)\rangle$ (114)
$+\int_{x,y}\mathcal{R}_{k}(x,y)\big{\langle}\big{(}{\varphi}^{\prime}_{k}(y)-{\varphi}(y)\big{)}\partial_{k}{\varphi}^{\prime}_{k}(x)\big{\rangle}\ .$
The $k$-derivative of ${\varphi}^{\prime}_{k}$ is taken at fixed $\chi^{\prime}$. The additional terms in the flow equation involve $\partial_{k}{\varphi}^{\prime}_{k}$ and therefore vanish for a $k$-independent field transformation. Expressed in terms of $\chi^{\prime}$ the additional terms involve higher correlation functions.
For the last term in eq. (V) we can employ a general identity for operators $\mathcal{A}[{\varphi}^{\prime}]$,
$\big{\langle}\big{(}{\varphi}^{\prime}(x)-{\varphi}(x)\big{)}\mathcal{A}[{\varphi}^{\prime}]\big{\rangle}=\int_{y}G(x,y)\frac{\partial}{\partial{\varphi}(y)}\langle\mathcal{A}[{\varphi}^{\prime}]\rangle\ ,$ (115)
with $G(x,y)=\big{(}\Gamma_{{\varphi},k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}(x,y)$ given by eq. (IV). This is found directly by using the definition of the expectation value in the presence of sources or the associated value of ${\varphi}$, and the infrared cutoff term
$\mathcal{A}[{\varphi}]=\langle\mathcal{A}[{\varphi}^{\prime}]\rangle=\frac{\int\mathcal{D}\chi^{\prime}\mathcal{A}[{\varphi}^{\prime}]\exp\big{\{}-F[\chi^{\prime}]\big{\}}}{\int\mathcal{D}\chi^{\prime}\exp\big{\{}-F[\chi^{\prime}]\big{\}}}\ ,$ (116)
with
$F[\chi^{\prime}]=S[\chi^{\prime}]+\Delta_{k}S[\chi^{\prime}]-\int_{y}\frac{\partial\Gamma^{\prime}_{{\varphi},k}}{\partial{\varphi}(y)}\big{(}{\varphi}^{\prime}(y)-{\varphi}(y)\big{)}\ .$ (117)
Here $\Delta_{k}S$ involves the cutoff term (second term in the curled bracket in eq. (110)), which depends on $\chi^{\prime}$ through ${\varphi}^{\prime}(\chi^{\prime})$. Transporting the ${\varphi}$-derivative inside the integral,
$\frac{\partial}{\partial{\varphi}(x)}\mathcal{A}[{\varphi}]=\int_{y}\frac{\partial^{2}\tilde{\Gamma}}{\partial{\varphi}(x)\partial{\varphi}(y)}\big{\langle}\big{(}{\varphi}^{\prime}(y)-{\varphi}(y)\big{)}\mathcal{A}[{\varphi}^{\prime}]\rangle\ ,$ (118)
and multiplication with $G$ yields eq. (115).
Inserting for $\mathcal{A}[{\varphi}^{\prime}]$ the logarithmic $k$-derivative of ${\varphi}^{\prime}_{k}(x)$, with expectation value
$\gamma(x)=\langle\partial_{t}{\varphi}^{\prime}_{k}(x)\rangle\ ,$ (119)
one arrives at the flow equation
$\partial_{t}\Gamma_{{\varphi},k}=$ $\,\frac{1}{2}\int_{x,y}\bigg{[}\partial_{t}\mathcal{R}_{k}(x,y)+2\int_{z}\frac{\partial\gamma(z)}{\partial{\varphi}(x)}\mathcal{R}_{k}(z,y)\bigg{]}$
×$\times\big{(}\Gamma_{{\varphi},k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}(y,x)-\int_{x}\frac{\partial\Gamma_{{\varphi},k}}{\partial{\varphi}(x)}\gamma(x)\ .$ (120)
This is the form of the flow equation discussed in refs. Pawlowski (2007); Baldazzi et al. (2022); Baldazzi and Falls (2021); Ihssen and Pawlowski (2023b), in analogy to the discussion of frame independence by Wegner Wegner (1974), see also Salmhofer (2007). The quantity $\gamma(x)$ is often called the “RG-kernel”.
This flow equation shows similarities with the flow equation (IV) for $\Gamma_{\chi,k}[{\varphi}]$. For linear field transformations the macroscopic and microscopic field transformations lead to identical results. For the macroscopic field transformations the connection term vanishes. For the microscopic field transformations the term involving $\partial\gamma/\partial{\varphi}$ does not depend on ${\varphi}$. The flow equation (V) coincides with the flow equation (86) provided that one chooses a suitable adapted cutoff $\mathcal{R}_{k}$. This may be verified by reproducing eq. (IV) for the transformation (91). For eq. (87) one has $\tilde{\gamma}_{k}[{\varphi}]=\gamma_{k}[{\varphi}]$.
For non-linear field transformations one finds some important differences. As may be expected, the main difference concerns the infrared cutoff, which is now a bilinear form in the field ${\varphi}^{\prime}$ instead of $\chi^{\prime}$. As a consequence, there is no “connection term” $\Delta$ as in eq. (IV). Also the factors $\tilde{N}_{m}^{-1}$ reflecting the “inhomogeneity” of the cutoff translated to the composite field are absent. The “correction terms” involving $\gamma$ in eq. (V) have a somewhat different form as compared to eq. (IV).
We observe that the relation between the source and the macroscopic composite field ${\varphi}$ depends on $k$. For a fixed source the macroscopic field depends on $k$. Different prescriptions on what is kept fixed during the flow lead to different versions of flow equations even if the same cutoff function for the composite field ${\varphi}^{\prime}$ is used. An early version of a flow equation with cutoff for “composite” fields ${\varphi}^{\prime}$ can be found in ref. Gies and Wetterich (2002). All the different versions of field transformations and corresponding definitions of $\Gamma_{k}[{\varphi}]$ have been used in practice for a simplification of flow equations, for example by absorbing unwanted terms in the effective action by a suitable choice of $k$-dependent “composite fields”. This can be achieved by the “correction terms” induced by $k$-dependent fields canceling terms induced by the trace term in the flow equation. Flowing bosonization is a good example of this concept Gies and Wetterich (2004); Braun (2009). Also the ideas of “essential renormalization flow” Baldazzi et al. (2022); Baldazzi and Falls (2021) or “optimized flow” Ihssen and Pawlowski (2023b) rely on the elimination of couplings in the effective average action by a suitable choice of field variables.
RG-kernel
The version with infrared cutoff and source for composite fields offers for non-linear transformations an important advantage as compared to eq. (IV): The cutoff can be adapted to the propagator for composite fields. The inhomogeneity of the cutoff due to the field-dependent factors $\tilde{N}_{m}^{-1}$ can be avoided. There is, however, a price to be paid for this setting. First, the RG-kernel is, in general, not easily computable for a given field transformation. Inversely, for a given RG-kernel the underlying microscopic field transformation is often not known. Second, for non-linear microscopic field transformations the direct contact of the effective action for large $k$ to the classical action is lost. We address the second issue in the next section and discuss here some consequences of the non-linearity of the field transformation.
For ${\varphi}^{\prime}_{k}$ depending non-linearly on $\chi^{\prime}$ the $k$-derivative typically involves higher correlation functions of ${\varphi}^{\prime}_{k}$. As an example, we take
${\varphi}^{\prime}_{k}(x)=\chi^{\prime}(x)+g(k)\chi^{\prime}(x)^{2}\ ,$ (121)
with
$\partial_{k}{\varphi}^{\prime}_{k}(x)=\partial_{k}g(k)\chi^{\prime}(x)^{2}=\partial_{k}\ln g(k)\big{(}{\varphi}^{\prime}_{k}-\chi^{\prime}({\varphi}^{\prime}_{k})\big{)}\ ,$ (122)
where
$\chi^{\prime}({\varphi}^{\prime}_{k})=\frac{1}{2g}\big{(}\sqrt{1+4g{\varphi}^{\prime}_{k}}-1\big{)}\ .$ (123)
Expanding around some arbitrary ${\varphi}^{\prime}_{k0}$, arbitrarily high powers of ${\varphi}^{\prime}_{k}$ appear in $\partial_{k}{\varphi}^{\prime}_{k}(x)$. This results for the flow equation (V), (V) in the presence of correlation functions for ${\varphi}_{k}^{\prime}$ of arbitrary high order. Those could be computed, in principle, by expressions involving high functional derivatives. This destroys, however, the simple closed form of the flow equation. At this point the $k$-dependence of ${\varphi}^{\prime}_{k}$ reflects only the choice of composite field ${\varphi}^{\prime}_{k}[\chi^{\prime}]$. In general, this choice needs not to be related to the flow of the effective action.
One may go the opposite way and specify the RG-kernel $\gamma$ such that the flow equations simplify. In this case the microscopic transformation ${\varphi}_{k}^{\prime}(\chi^{\prime})$ is often no longer known and the existence of an invertible transformation of the microscopic variables is not guaranteed. One may try to construct the microscopic field transformation ${\varphi}^{\prime}_{k}[\chi^{\prime}]$ from a given $\gamma_{k}[{\varphi}]$. For this purpose we integrate the equation,
$k\partial_{k}{\varphi}(k;\chi)=\gamma_{k}\big{[}{\varphi}(k;\chi)\big{]}\ ,$ (124)
with boundary condition
${\varphi}(\Lambda;\chi)=\chi\ .$ (125)
One may take large $\Lambda$ associated to a scale in microphysics or $\Lambda\to\infty$. The solution of this differential equation ${\varphi}(k;\chi)$ describes how the expectation value $\langle{\varphi}^{\prime}_{k}(\chi^{\prime})\rangle={\varphi}(k;\chi)$ evolves with $k$ if for $k=\Lambda$ one has $\langle{\varphi}^{\prime}_{\Lambda}(\chi^{\prime})\rangle=\chi$.
Eq. (124) and its solution concerns the expectation value and needs translation to ${\varphi}^{\prime}_{k}(\chi^{\prime})$. A microscopic field transformation ${\varphi}^{\prime}_{k}(\chi^{\prime})$ with ${\varphi}^{\prime}_{\Lambda}(\chi^{\prime})=\chi^{\prime}$ obeys $\langle{\varphi}^{\prime}_{\Lambda}(\chi^{\prime})\rangle=\langle\chi^{\prime}\rangle$ and we can choose $\chi=\langle\chi^{\prime}\rangle$. One may try to use for all $k$ the ansatz
${\varphi}^{\prime}_{k}(\chi^{\prime})={\varphi}(k;\chi^{\prime})\ ,$ (126)
with ${\varphi}(k;\chi)$ the solution of eqs. (124), (125). Evaluating the $k$-derivative of the expectation value of eq. (126) yields
$\langle\partial_{t}{\varphi}^{\prime}_{k}(\chi^{\prime})\rangle=\big{\langle}\gamma_{k}\big{[}{\varphi}^{\prime}_{k}(\chi^{\prime})\big{]}\big{\rangle}\ .$ (127)
On the other hand, the definition of the RG-kernel reads
$\langle\partial_{t}{\varphi}^{\prime}_{k}(\chi^{\prime})\rangle=\gamma_{k}\big{[}\langle{\varphi}^{\prime}_{k}(\chi^{\prime})\rangle\big{]}\ .$ (128)
Eqs. (127) and (128) are compatible only for linear RG-kernels. If $\gamma_{k}$ depends non-linearly on ${\varphi}$ the ansatz (126) leads to a contradiction since $\langle({\varphi}^{\prime}_{k}(\chi^{\prime}))^{n}\rangle\neq\langle{\varphi}^{\prime}_{k}(\chi^{\prime})\rangle^{n}$. For non-linear RG-kernels this type of reconstruction does not work. We see no simple way to find the microscopic field transformation ${\varphi}^{\prime}_{k}(\chi^{\prime})$ for a given non-linear $\gamma_{k}[{\varphi}]$.
If one is able to find the solution ${\varphi}(k;\chi)$ for a linear RG-kernel, one can then judge if ${\varphi}^{\prime}_{k}(\chi^{\prime})$ remains invertible for all $k$. In contrast, for a non-linear RG-kernel it seems to be very hard in practice to verify a given $\gamma_{k}[{\varphi}]$ corresponds to a regular microscopic field transformation ${\varphi}^{\prime}_{k}(\chi^{\prime})$.
Linear RG-kernel
For linear RG-kernels $\partial_{t}{\varphi}^{\prime}_{k}$ is linear in ${\varphi}^{\prime}_{k}$. This can be realized for a particular family of non-linear field transformations ${\varphi}^{\prime}_{k}(\chi^{\prime})$.
Let us consider local variable transformations of the form
${\varphi}^{\prime}=N(\chi^{\prime})\chi^{\prime}\ ,$ (129)
with RG-kernel
$\gamma(x)=\big{\langle}\big{(}\partial_{t}N(\chi^{\prime})\chi^{\prime}\big{)}(x)\big{\rangle}\ .$ (130)
For general functions $N(\chi^{\prime})$ there is no simple expression for $\gamma(x)$. For the special case
$\partial_{t}N(\chi^{\prime})=a(k)N(\chi^{\prime})+\frac{b(k)}{\chi^{\prime}}\ ,\quad\partial_{t}{\varphi}^{\prime}=a{\varphi}^{\prime}+b\ ,$ (131)
the expression $\partial_{t}{\varphi}^{\prime}=\partial_{t}N(\chi^{\prime})\chi^{\prime}$ is linear in ${\varphi}^{\prime}$ and therefore implies a linear RG-kernel
$\gamma(x)=a(k){\varphi}(x)+b(k)\ .$ (132)
This particular choice of the field ${\varphi}^{\prime}(\chi^{\prime})$ yields the simple flow equation
$\partial_{t}\Gamma_{{\varphi},k}=$ $\,\frac{1}{2}\mathrm{tr}\Big{\{}\big{(}\partial_{t}+2a(k)\big{)}\mathcal{R}_{k}\big{(}\Gamma_{{\varphi},k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}\Big{\}}$
$-\int_{x}\frac{\partial\Gamma_{{\varphi},k}}{\partial{\varphi}(x)}\big{(}a(k){\varphi}(x)+b(k)\big{)}\ .$ (133)
In the appendix C we employ microscopic field transformations with a linear RG-kernel in order to discuss under which circumstances the flowing effective action (102) yields a model of free composite scalar fields. We conclude that using linear RG-kernels (132) the correction terms for non-zero $a$ or $b$ in the flow equation (V) can be employed to achieve certain simplifications. For this restricted set of field transformations one can keep control of the microscopic transformations.
This is no longer the case for more general non-linear RG-kernels $\gamma({\varphi})$ to which we turn next. The possibility to assume a suitable non-linear RG-kernel without explicit knowledge of the microscopic field transformation is further discussed in sect. VII.
Power and dangers of non-linear microscopic field transformations
Microscopic field transformations with a non-linear RG-kernel offer very rich possibilities to change the functional form of the effective average action and associated flow equations. This may be regarded as a powerful tool for a simplification of the flow equations. Terms in the flow equations which are responsible for technical difficulties can be removed by a suitable choice of the RG-kernel. Some proponents of the essential renormalization group suggest that all couplings that can be removed by a suitable choice of a non-linear RG-kernel can be regarded as redundant and play no role for observations. For example, it has been argued that in quantum gravity the cosmological constant is redundant in this sense.
This view is dangerous in two aspects. First, it is often not guaranteed that a given choice of a non-linear RG-kernel $\gamma[{\varphi}]$ can indeed be realized by a regular microscopic field transformation. We have seen that in general the relation between the RG-kernel and the field transformation is rather complex. In many cases it is not accessible in practice. The requirement of regularity of the microscopic field transformation imposes constraints on the allowed form of $\gamma[{\varphi}]$. Omitting these constraints which are unknown in practice presents then the danger that the chosen form of $\gamma[{\varphi}]$ is actually not consistent with a microscopic definition of the effective average action.
The second danger is related to the necessity of truncations. In practice, this is often the most important shortcoming. Assuming that a regular field transformation producing a given RG-kernel exists, the flow equation (V) is an exact identity for a well defined effective action which generates correlation functions for composite fields. A solution of this flow equation requires in most cases a truncation of the general form of the effective average action. The danger of non-linear microscopic field transformations consists in the problem that important physical information can be moved to sectors which are no longer resolved for a given truncation.
As an example we may again consider the momentum dependence of the propagator for electrons in a solid. It contains central information on the Fermi surface and many central quantities for observations. The field dependence of this propagator encodes key aspects as the possibility of spontaneous symmetry breaking. By a suitable choice of the RG-kernel one may transform this momentum- and field-dependence to a simple standard form which no longer reflects the details of the Fermi surface. As discussed in the introduction, this does not mean that the Fermi surface or the effects of spontaneous symmetry breaking on the fermion propagator are not observable. The field transformation only shifts the information about the physical Fermi surface to higher order correlation functions. If a truncation omits the detailed momentum- and field-dependence of vertices, the use of the field transformation can remove key physical information by the truncation.
In the context of the simple scalar field theory we will demonstrate this issue by the role of the effective scalar potential. The use of suitable RG-kernels $\gamma[{\varphi}]$ can both produce and remove the effective potential in the effective average action. Removing the potential by a microscopic field transformation shifts the relevant physical information to sectors that are more difficult to access. A too simple truncation of these sectors will discard important physical information.
Effective scalar potential
The linear RG-kernel (132) is special in the sense that no scalar potential for ${\varphi}$ is generated. For non-linear kernels the derivative $\partial\gamma/\partial{\varphi}$ will depend on ${\varphi}$. Then the flow generates a potential even if one starts at high $k$ without a potential. We demonstrate this for a setting with constant kinetial $K=Z$.
The flow equation for the effective potential is found by evaluating the flow of $\Gamma_{k}$ for a homogeneous scalar field ${\varphi}(x)={\varphi}$, $\Gamma_{k}=\int_{x}U_{k}({\varphi})$. We assume that in this case also $\gamma$ is homogeneous, $\gamma=\gamma({\varphi})$. The flow equation for $U_{k}$ obtains from eq. (V) as
$\partial_{t}U_{k}=$ $\,\frac{1}{2}\int_{q}\left[\partial_{t}\mathcal{R}_{k}(q^{2})+2\frac{\partial\gamma}{\partial{\varphi}}\mathcal{R}_{k}(q^{2})\right]\big{[}Zq^{2}+\mathcal{R}_{k}(q^{2})\big{]}^{-1}$
$+\Delta\partial_{t}U\ .$ (134)
Here $\Delta\partial_{t}U$ indicates terms proportional to derivatives of $U$ with respect to ${\varphi}$ or contributions from fluctuations of other fields. As long as these derivatives are small we can employ on the r.h.s. of the flow equation $\Gamma_{k}=(Z/2)\int_{x}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}$. The ${\varphi}$-dependent part of the potential therefore flows as
$\partial_{t}\left(\frac{\partial U}{\partial{\varphi}}\right)=\frac{\partial^{2}\gamma}{\partial{\varphi}^{2}}\int_{q}\frac{\mathcal{R}_{k}(q^{2})}{Zq^{2}+\mathcal{R}_{k}(q^{2})}=Ck^{4}\frac{\partial^{2}\gamma}{\partial{\varphi}^{2}}\ .$ (135)
If we start at very high $k=\Lambda$ with a flat potential, $\partial U/\partial{\varphi}=0$, a non-zero slope will develop for $k<\Lambda$. For positive $\partial^{2}\gamma/\partial{\varphi}^{2}>0$ this slope is negative and leads to a tendency towards spontaneous symmetry breaking. For negative $\partial^{2}\gamma/\partial{\varphi}^{2}<0$ the potential develops typically a minimum at ${\varphi}=0$ with a positive mass term $m_{\varphi}^{2}\sim k^{4}\partial^{2}\gamma/\partial{\varphi}^{2}$. (We assume here a discrete symmetry ${\varphi}\to-{\varphi}$ for which $\gamma(-{\varphi})=-\gamma({\varphi})$, $\gamma({\varphi})=a{\varphi}+d{\varphi}^{3}+\ldots$.)
One could also be tempted to use the non-linear field transformations in order to remove the effective scalar potential. Imagine that the fluctuations of additional fields, say fermions with a Yukawa coupling to the scalar, induce on the r.h.s. of eq. (V) terms of the type
$\Delta\partial_{t}U=A_{1}k^{2}{\varphi}^{2}+A_{2}{\varphi}^{4}\ .$ (136)
One could then take an RG-kernel of the form
$\gamma=-\frac{A_{1}{\varphi}^{3}}{3Ck^{2}}-\frac{A_{2}{\varphi}^{5}}{5Ck^{4}}\ ,$ (137)
such that the combined flow of $\partial_{t}(\partial U/\partial{\varphi})$ vanishes. If a microscopic field transformation leading to the RG-kernel (137) exists, the potential for ${\varphi}$ remains flat for all $k$. In the language of the “essential renormalization group” it would seem that one can remove the potential as “irrelevant couplings”.
Conservation of physical information
The possibility to remove the effective potential by a non-linear field transformation may seem surprising at first sight. Fermion fluctuations typically induce terms of the form
$A_{1}=\frac{a_{1}k^{2}}{k^{2}+m^{2}}\ ,\quad A_{2}=-\frac{a_{2}k^{4}}{(k^{2}+m^{2})^{2}}\ ,$ (138)
with $m$ the mass of the fermions depending on ${\varphi}$ via a Yukawa coupling. Without removing the potential by the field transformation (137) the fermion fluctuations would induce for $k\to 0$ a potential indicating spontaneous symmetry breaking,
$U_{k\to 0}=-\frac{\mu^{2}}{2}{\varphi}^{2}+\frac{\lambda}{8}{\varphi}^{4}+\dots\ .$ (139)
(We assume a vanishing potential at some microscopic scale $\Lambda$.) The question arises where to see this spontaneous symmetry breaking after the potential has been eliminated by a field transformation.
Field transformations can only move physical features to other sectors. On the one side the last term in the flow equation (V) generates a more complex scalar kinetic term according to
$\int_{x}\partial^{\mu}\partial_{\mu}{\varphi}\gamma({\varphi})=-\int_{x}\frac{\partial\gamma}{\partial{\varphi}}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}\ .$ (140)
On the other side, this term also induces more complex interactions between the scalar and fermions $\psi$ (with $h$ a Yukawa coupling) according to
$\int_{x}-\gamma(x)\frac{\partial}{\partial{\varphi}(x)}\int_{y}h{\varphi}(y)\overline{\psi}(y)\psi(y)=-\int_{x}h\gamma({\varphi})\overline{\psi}\psi\ .$ (141)
The physical information about spontaneous symmetry breaking induced by fermion fluctuation is only lost if a truncation discards these modifications of the kinetial and Yukawa coupling. Only if a truncation keeps the field-dependence of the kinetial and scalar-fermion coupling one can recover the spontaneous symmetry breaking even for a field choice that eliminates the effective potential.
It may be rather difficult to extract the features of spontaneous symmetry breaking from the flow of these higher correlation functions. Given the simplicity of the picture with a scalar potential it seems preferable to keep as much of the information as possible in the form of the scalar potential. Similar considerations apply to the physical information about field-dependent structures in the propagator or two-point function, as a gap induced by spontaneous symmetry breaking. While non-linear field transformations may allow one to move the information contained in the two-point function to other sectors, it will be, in general, more cumbersome to extract the relevant information from those higher correlation functions.
At this point it may be useful to highlight an important difference between macroscopic and microscopic field transformations. The macroscopic field transformations do not change the effective average action. It only expresses a given functional in terms of different field-variables. In this way one can explicitly follow the action of the transformation once a given approximation to $\Gamma_{k}$ is given. For example, a transformation ${\varphi}(x)=f^{-1}_{k}\big{(}\chi(x)\big{)}$ transforms $U(\chi)\to U({\varphi})=U\big{(}f_{k}({\varphi})\big{)}$. A regular transformation $f_{k}$ can transform an arbitrary convex potential $U(\chi)$ to $U({\varphi})=m^{2}{\varphi}^{2}+U_{0}$ for a suitable range of fields. This holds for arbitrarily small $m^{2}$. A potential with spontaneous symmetry breaking of the discrete symmetry $\chi\to-\chi$ cannot be transformed to this form, however. Monotonically decreasing potentials can be brought to the form $U=M^{4}\exp(-{\varphi}/M)$ for a suitable range of fields. Transforming the potential to a standard form shifts information to the kinetial. Standard forms of the potential have proven to be useful for inflationary cosmology Wetterich (2013b); Kallosh et al. (2014); Wetterich (2015c); Galante et al. (2015) or dynamical dark energy Hebecker and Wetterich (2001); Wetterich (2003).
Microscopic field transformations change the flow equation and the effective average action beyond the simple use of different variables. Their action is much less under control and it is often not obvious to which sector the relevant information is shifted. For example, one may doubt if the non-linear RG-kernel which removes the scalar potential completely can be realized by a suitable regular field transformation.
Flow equations for fundamental and composite fields
Two issues remain to be settled if one uses the IR-cutoff for composite fields. The first concerns the relation between the microscopic action $S[\chi^{\prime}]$ and the microscopic effective action $\Gamma_{{\varphi},k}$ for very large $k=\Lambda$, or for $k\to\infty$. The second involves the relation between $\Gamma_{{\varphi},k}[{\varphi}]$ and expectation values and correlations for the “fundamental field” $\chi^{\prime}$. In particular one would like to know the relation between $\langle\chi^{\prime}\rangle$ and ${\varphi}$. Knowing $\langle\chi^{\prime}\rangle=F[{\varphi}]$ one could employ a macroscopic variable transformation $\chi=F[{\varphi}]$ in order to compute $\Gamma_{{\varphi},k}[\chi]$. This relation,
$\langle\chi^{\prime}\rangle=F[{\varphi}]\ ,$ (142)
may differ from the microscopic relation
${\varphi}^{\prime}=$ $\,H[\chi^{\prime}]=N(\chi^{\prime})\chi^{\prime}\ ,$
$\chi^{\prime}=$ $\,H^{-1}[{\varphi}^{\prime}]=N^{-1}[{\varphi}^{\prime}]{\varphi}^{\prime}\ .$ (143)
One can address these issues within the two-field formalism introduced in sect. III.
Effective average action for fundamental and
composite fields
The two-field effective action in eq. (III) can be extended by adding an infrared cutoff terms $\mathcal{R}_{k,\chi}$ for the fundamental fields, and $\mathcal{R}_{k,{\varphi}}$ for the composite fields
$\Gamma_{k}^{\prime}[\chi,{\varphi}]=-\ln\int\mathcal{D}\chi^{\prime}\,\exp\bigg{\{}-S[\chi^{\prime}]$
$-\frac{1}{2}\int_{x,y}\Big{[}\chi^{\prime}(x)\mathcal{R}_{k,\chi}(x,y)\chi^{\prime}(y)$
$\quad\quad\quad\quad+H\big{(}\chi^{\prime}\big{)}(x)\mathcal{R}_{k,{\varphi}}(x,y)H\big{(}\chi^{\prime}\big{)}(y)\Big{]}$
$+\int_{x}\bigg{[}\frac{\partial\Gamma_{k}^{\prime}}{\partial\chi(x)}\big{(}\chi^{\prime}(x)-\chi(x)\big{)}$
$\quad\quad+\frac{\partial\Gamma_{k}^{\prime}}{\partial{\varphi}(x)}\big{(}H\big{(}\chi^{\prime}\big{)}(x)-{\varphi}(x)\big{)}\bigg{]}\bigg{\}}\ .$ (144)
For $\mathcal{R}_{k,\chi}=0$ and $\partial\Gamma_{k}^{\prime}/\partial\chi(x)=0$ one recovers the effective action $\Gamma_{{\varphi},k}[{\varphi}]$ defined by eq. (110). We may define $\Gamma_{{\varphi},k}[\chi,{\varphi}]$ by setting in eq. (VI) the infrared cutoff for the microscopic fields to zero, $R_{k,\chi}=0$. Solving the field equation for $\chi[{\varphi}]$,
$\frac{\partial\Gamma_{{\varphi},k}^{\prime}[\chi,{\varphi}]}{\partial\chi(x)}\big{(}\chi_{0}[{\varphi}]\big{)}=0\ ,$ (145)
one recovers
$\Gamma_{{\varphi},k}^{\prime}[{\varphi}]=\Gamma_{{\varphi},k}^{\prime}\big{[}\chi_{0}[{\varphi}],{\varphi}\big{]}\ .$ (146)
On the other hand, additional information contained in $\Gamma_{{\varphi},k}^{\prime}[\chi,{\varphi}]$ for arbitrary $\chi$ permits us to compute correlation functions for $\chi^{\prime}$ in the presence of sources for $\chi^{\prime}$ and ${\varphi}^{\prime}$, and in the presence of the IR-cutoff $\mathcal{R}_{k,{\varphi}}$. These are the relations
$\chi(x)=\langle\chi^{\prime}(x)\rangle\ ,$ (147)
or
$\int_{y}\langle\chi^{\prime}(x)\chi^{\prime}(y)\rangle_{c}\frac{\partial^{2}\Gamma_{{\varphi},k}^{\prime}[\chi,{\varphi}]}{\partial\chi(y)\partial\chi(z)}=\delta(x-z)\ .$ (148)
In particular, the solution of the field equation for $\chi$ in eq. (145) defines $\langle\chi^{\prime}\rangle=F[{\varphi}]$ by
$\langle\chi^{\prime}\rangle=\chi_{0}[{\varphi}]=F[{\varphi}]\ ,$ (149)
since for this solution the source term for the microscopic field vanishes.
We emphasize that the relation $\langle\chi^{\prime}\rangle=F[{\varphi}]$ cannot be obtained from the effective average action $\Gamma_{{\varphi},k}$. It needs the additional information provided by the two-field formalism.
Initial value problem
The two field formalism sheds light on the initial value problem for $\Gamma_{{\varphi},k}$. We will see that for the initial value of the flow at large $k$ or $k\to\infty$ the issue of the Jacobian in the microscopic field transformation for the functional integral comes back in an indirect way. For this finding it does not matter if we change the microscopic integration variables or not. The flow equation for $\Gamma_{{\varphi},k}[\chi,{\varphi}]$ is the same as for $\Gamma_{{\varphi},k}[{\varphi}]$, e.g. eq. (V), except for the additional argument $\chi$ of the effective action. For its solution we need the initial value of $\Gamma_{{\varphi},k}[\chi,{\varphi}]$ for a large value $k=\Lambda$ or $k\to\infty$. The initial value of $\Gamma_{{\varphi},\Lambda}[{\varphi}]$ is then obtained from eq. (146).
By subtraction of the macroscopic cutoff term the functional expression for $\Gamma_{{\varphi},k}[\chi,{\varphi}]$ reads
$\Gamma_{{\varphi},k}[\chi,{\varphi}]=\Gamma_{{\varphi},k}^{\prime}[\chi,{\varphi}]-\frac{1}{2}\int_{x,y}{\varphi}(x)\mathcal{R}_{k,{\varphi}}(x,y){\varphi}(y)$
$=$ $\,-\ln\int\mathcal{D}\chi^{\prime}\,\exp\bigg{\{}-S[\chi^{\prime}]$
$-\frac{1}{2}\int_{x,y}\big{[}H\big{(}\chi^{\prime}\big{)}(x)-{\varphi}(x)\big{]}\mathcal{R}_{k,{\varphi}}(x,y)\big{[}H\big{(}\chi^{\prime}\big{)}(y)-{\varphi}(y)\big{]}$
$+\int_{x}\bigg{[}\frac{\partial\Gamma_{{\varphi},k}}{\partial{\varphi}(x)}\big{[}H\big{(}\chi^{\prime}\big{)}(x)-{\varphi}(x)\big{]}$
$\quad\quad+\frac{\partial\Gamma_{{\varphi},k}}{\partial\chi(x)}\big{[}\chi^{\prime}(x)-\chi(x)\big{]}\bigg{]}\bigg{\}}\ .$ (150)
For $k\to\infty$ the exponent in the functional integral is dominated by the term ${}\sim\mathcal{R}_{k,{\varphi}}\sim k^{2}$. We want to evaluate $\Gamma_{{\varphi},\Lambda}[\chi,{\varphi}]$ in the limit $k=\Lambda\to\infty$.
For $k\to\infty$ we employ $H(\chi^{\prime})=F^{-1}(\chi^{\prime})$ and define $\hat{\chi}$ by
$\chi^{\prime}=F({\varphi})+\hat{\chi}\ .$ (151)
One infers
$H(\chi^{\prime})=F^{-1}\big{(}F({\varphi})+\hat{\chi}\big{)}={\varphi}+A({\varphi})\hat{\chi}+\ldots$ (152)
where
$A({\varphi})=\frac{\partial H}{\partial\chi^{\prime}}\big{(}F({\varphi})\big{)}\ .$ (153)
Keeping only the leading terms and making a shift in the integration variable yields
$\Gamma_{{\varphi},k}[\chi,{\varphi}]=S\big{[}F({\varphi})\big{]}-\int_{x}\frac{\partial\Gamma_{{\varphi},k}}{\partial\chi(x)}\big{(}F({\varphi})-\chi\big{)}(x)+\Delta\Gamma_{k}[\chi,{\varphi}]\ ,$ (154)
with
[…]
$\quad\quad-\frac{1}{2}\int_{x,y}\big{(}A({\varphi})\hat{\chi}\big{)}(x)\mathcal{R}_{k,{\varphi}}(x,y)\big{(}A({\varphi})\hat{\chi}\big{)}(y)$
$+\int_{x}\bigg{[}\frac{\partial\Gamma_{{\varphi},k}}{\partial{\varphi}(x)}A\big{(}{\varphi}\big{)}(x)+\frac{\partial\Gamma_{{\varphi},k}}{\partial\chi(x)}-\frac{\partial S[\chi]}{\partial\chi(x)}\big{(}F({\varphi})\big{)}\bigg{]}\hat{\chi}(x)\bigg{\}}\ .$ (155)
We focus on solutions of the partial field equation (145) for which one infers
$\Gamma_{{\varphi},k}[{\varphi}]=S\big{[}F({\varphi})\big{]}+\Delta\Gamma_{k}\big{[}\chi_{0}({\varphi}),{\varphi}\big{]}\ .$ (156)
In the limit $k\to\infty$ we can take $\mathcal{R}_{k}(x,y)=k^{2}\delta(x-y)$. Contributions from the term linear in $\hat{\chi}(x)$ in the exponent in eq. (VI) are suppressed by $k^{-2}$ and we end with the Gaussian functional integral
$\Delta\Gamma_{k}[\chi,{\varphi}]=-\ln\int\mathcal{D}\hat{\chi}\,\exp\bigg{\{}-\frac{1}{2}\int_{x}k^{2}A^{2}\big{(}{\varphi}\big{)}(x)\hat{\chi}^{2}(x)\bigg{\}}\ .$ (157)
The field-dependent part of this term reads
$\Delta$ $\Gamma_{k}[\chi,{\varphi}]=-\ln\det A^{-1}({\varphi})$
$=$ $\,-\mathrm{tr}\ln A^{-1}({\varphi})=\mathrm{tr}\ln A({\varphi})=\int_{x}\varepsilon^{-4}\ln\big{[}A\big{(}{\varphi}\big{)}(x)\big{]}\ .$ (158)
It depends on $\chi$, but not on ${\varphi}$. With
$A({\varphi})=\frac{\partial{\varphi}^{\prime}}{\partial\chi^{\prime}}\bigg{|}_{F({\varphi})}\ ,$ (159)
we realize that a term corresponding to the Jacobian in eqs. (II)- (16) arises now on the macroscopic level. The initial value of $\Gamma_{{\varphi},k}[{\varphi}]$ for $k\to\infty$ involves the Jacobian of the non-linear microscopic field transformation.
This is consistent with the fact that one could have done a transformation of integration variables on the microscopic level. With a functional integral over ${\varphi}^{\prime}$ and a classical action $\bar{S}[{\varphi}^{\prime}]$, the limit $k\to\infty$ of the effective average action results directly in $\Gamma_{k\to\infty}[{\varphi}]=\bar{S}[{\varphi}]$, which contains the contribution of the Jacobian in eq. (II). Unfortunately, one recovers the dependence on the regularization encoded in $\varepsilon^{-4}$. This may have the consequence that the effective action $\Gamma_{{\varphi},k}$ may be difficult to handle in practice for the extraction of macroscopic information for a given microscopic model. In particular, starting with $S[\chi^{\prime}]$ involving only a non-linear kinetic term (3) it seems unavoidable that the microscopic effective action $\Gamma_{{\varphi},k\to\infty}$ involves a scalar potential and does not correspond to a model of a free massless scalar field. In summary, the initial values of the effective average action $\Gamma_{{\varphi},k}[{\varphi}]$ are precisely the ones corresponding to a variable transformation in the functional integral. As long as the variable transformation ${\varphi}^{\prime}[\chi^{\prime}]$ is independent of $k$ one retains the standard flow equation. The correction terms in eq. (V) arise if one decides to make the microscopic field transformation $k$-dependent.
Fundamental and composite fields
We conclude that the effective action and the functional flow equations often single out in practice a preferred choice of “canonical or fundamental fields”. These are the fields for which the measure is linear in the fields, $\int\mathcal{D}\chi^{\prime}=\int\mathcal{D}\hat{\chi}$, for $\hat{\chi}=\chi^{\prime}-\chi$, and for which the classical action $S[\chi^{\prime}]$ is supposed to be known. These are typically the fields in terms of which a given model is formulated. In principle, microscopic field transformations to “composite fields” which depend non-linearly on $\chi^{\prime}$ are possible. They require, however, control of a Jacobian, which is typically not given in practice. This holds in an obvious way for the variable transformations in the functional integral. The problem reappears if we define the effective action and flow equations with sources and infrared cutoff for the composite field. It now concerns the initial value of the flow for large $k$. This inherits the Jacobian even if we keep the fundamental field $\chi^{\prime}$ as the integration variable. The Jacobian can be avoided if for $k=\Lambda$ the composite field coincides with the fundamental field.
While the two-field formalism finds its limitations if restricted to a vanishing infrared cutoff for $\chi^{\prime}$, it can provide important information and simplifications if we also introduce a non-vanishing cutoff $\mathcal{R}_{k,\chi}$ for the fundamental field $\chi^{\prime}$. Having at our disposal two cutoffs $\mathcal{R}_{k,\chi}$ and $\mathcal{R}_{k,{\varphi}}$, we can arrange the setting such that for $k\to\infty$ only $\mathcal{R}_{k,\chi}$ matters, implying the simple initial value of the microscopic effective action
$\Gamma_{k\to\infty}[\chi,{\varphi}]=S[\chi]\ .$ (160)
The infrared cutoff $\mathcal{R}_{k,{\varphi}}$ may only become relevant at lower $k$ where composite fields as ${\varphi}$ play an important role. Using suitable identities one can reformulate the two-field effective action as a functional integral over two fields $\chi^{\prime}$ and ${\varphi}^{\prime}$loerchinger and Wetterich (2009). Both fields are then formally on equal footing. This type of approach has found wide applications, for example for the description of mesons in a model of strong interactions involving quarks and gluons as fundamental fields Gies and Wetterich (2004); Ellwanger and Wetterich (1994); Jungnickel and Wetterich (1996); Berges et al. (1999); Jungnickel and Wetterich (1997); Meggiolaro and Wetterich (2001); Brouzakis and Tetradis (2004); Braun and Gies (2006); Braun (2009); Braun et al. (2011); Herbst et al. (2013); Haas et al. (2013); Mitter and Schaefer (2014); Herbst et al. (2014); Fu et al. (2020), see also Fukushima et al. (2011); Skokov et al. (2012); Eser and Blaizot (2021); Helmboldt et al. (2015); Mitter et al. (2015); Eser et al. (2015); Sedrakian et al. (2018); Resch et al. (2019); Braun et al. (2018a, b); Alkofer et al. (2019); Jakovác and Patkós (2019); Braun et al. (2020); Fejős and Patkós (2021); Braun and Schallmo (2022).
Scaling solutions
The functional flow equations may admit scaling solutions for which dimensionless functions of suitable renormalized dimensionless fields do no longer depend on $k$. In this way a given model can be extrapolated to arbitrarily large $k$, thereby constituting an ultraviolet completion. The limit $k\to\infty$ corresponds to an ultraviolet fixed point. The existence of the ultraviolet fixed point induces renormalizability of the theory. If the interactions are not small at the fixed point one encounters non-perturbative renormalizability or asymptotic safety Weinberg (1980).
As long as one is interested only in the existence of possible scaling solutions or fixed points and associated universality classes, no knowledge of the microscopic effective action $\Gamma_{k}$ for $k\to\infty$ is needed. The corresponding form of $\Gamma_{k\to\infty}$ is inferred a posteriori from the scaling solution. The freedom of non-linear field transformations may be helpful to find scaling solutions. In particular, these transformations can be very useful in order to identify the renormalized fields in terms of which the scaling solution becomes independent of $k$. For the microscopic field transformations in the flow equation a non-linear RG-kernel $\gamma[{\varphi}]$ can reduce the complexity of the flow equation and facilitate the detection of possible scaling solutions. Having found a scaling solution establishes a universality class for the corresponding model. The existence of the universality class is often sufficient for many purposes. No detailed microscopic realization of this universality class may be needed.
For our simple scalar model the use of suitable RG-kernels for $\Gamma_{{\varphi},k}$ permits the statement that there exists the universality class of a free massless scalar field, corresponding to the choice of the microscopic effective action
$\Gamma_{{\varphi},k\to\infty}=\frac{1}{2}\int_{x}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}\ .$ (161)
This statement is rather trivial. Since the relation between $\chi^{\prime}$ and ${\varphi}$ is not known this does not help much to answer the question what is the physics of the model given by eqs. (3), (6). Eq. (161) is related to eq. (2) by a macroscopic field transformation and therefore compatible with the existence of such an ultraviolet fixed point. For the interesting question if a crossover trajectory links this fixed point to the trivial fixed point for $k\to 0$ an elimination of the kinetial by a suitable RG-kernel does not seem to be helpful.
For more complex settings the possible scaling solutions and universality classes may not be known. An example is quantum gravity. In this case the flow equation (V) for $\Gamma_{{\varphi},k}$ could be helpful in order to identify candidates for scaling solutions. The use of an arbitrary functional $\gamma[{\varphi}]$ in order to eliminate unwanted couplings (sometimes called “inessential couplings”) is accompanied, however, by two issues. The first concerns the fact that for arbitrary $\gamma[{\varphi}]$ the relation between fundamental and composite fields is not known explicitly. In particular, there is no control if this relation is invertible or not. As a consequence, it is not guaranteed that a scaling solution found for unrestricted $\gamma[{\varphi}]$ is indeed a scaling solution for the model. Promoting such a candidate to a full scaling solution would require to find the microscopic field transformation that leads to the form of $\gamma[{\varphi}]$ used, and to establish its regularity.
One may argue that it is sufficient to find a solution of the coupled system of flow equations (119), (V), without the need of the existence of an associated microscopic field transformation. Eq. (119) is then taken purely as a change in the space of macroscopic fields ${\varphi}$, without the realization that ${\varphi}$ is the expectation value of some microscopic composite field. The problem with this view is that one could replace in this case the flow equation (V) by many other forms of modified flow equations. The specific form of eq. (V) is rooted in the microscopic formulation of the functional integral (V). If no such functional integral exists, the status of $\Gamma_{{\varphi},k}$ and its relation to correlation functions becomes unclear. It is no longer guaranteed that some microscopic realization exists for the ultraviolet completion associated to the scaling solution.
In short, many of the scaling solutions found for the “essential renormalization group” may be fake, either because no microscopic realization exists, or due to a shift of important physical information into sectors that are not resolved sufficiently by the truncation. This seems to happen for quantum gravity if higher order invariants based on polynomials of the curvature tensor are absorbed by the RG-kernel of microscopic field transformations. In this approach a large number of scaling solutions appears formally Knorr (2023b); Baldazzi et al. (2023). It is not a simple issue to decide which ones are physical or fake.
As a second issue the scaling solutions found by microscopic field transformations with arbitrary non-linear RG-kernels may not be complete. For our case of a scalar theory one could define $\gamma[{\varphi}]$ such that in first order of a derivative expansion the effective action is restricted effectively to a potential and a kinetic term with constant coefficient. The (approximate) scaling solution for this model in four dimensions is a “trivial theory” for which the quartic scalar coupling flows slowly towards zero as $k\to 0$. Due to a Landau pole this theory is not “ultraviolet complete” in the sense that it cannot be extended to $k\to\infty$. The only full scaling solution in this setting is the free massless scalar field. This statement does not answer the question if additional scaling solutions for the models of the type of (3) could exist. If scaling solutions with a non-trivial field dependence of the kinetial of the type discussed in appendices AB exist, the transformation to a canonical kinetic term may have moved the relevant information to the part that has been discarded by the truncation.
An example of missing scaling solutions for the essential renormalization group is asymptotic freedom in quantum gravity. In the presence of invariants quadratic in the curvature tensor quantum gravity is renormalizable Stelle (1977) and asymptotically free in perturbation theory Fradkin and Tseytlin (1982); Avramidy and Barvinsky (1985). By use of functional renormalization it has been possible to follow the flow of the relevant couplings into and inside the non-perturbative domain Sen et al. (2022). Scaling solutions have been found which link the asymptotically free ultraviolet fixed point to an infrared fixed point Sen et al. (2023). For the essential renormalization group the couplings multiplying the terms quadratic in the curvature tensor are eliminated by a field transformation. It seems hard to find asymptotic freedom and the corresponding scaling solution in this approach.
Conclusions
Transformations of the field variables for functional flow equations are a powerful and versatile tool, both for technical simplifications and the understanding of key physical features. General non-linear field transformations open, however, a very wide area. Pushed to extremes, especially for cases where the regularity of the field transformation is not under control, they entail the danger that important physical information becomes hidden. In case of approximations to the most general form of the effective action (truncations) key physical information may even get lost.
The present paper addresses a systematic discussion of field transformations, starting from variables in the functional integral and proceeding to transformations of field variables in the effective action and for flow equations for associated scale-dependent effective average actions. In order to illustrate the conceptual discussion with a practical case we have focused often on a simple scalar model with a field-dependent kinetic term.
For the field transformations in functional flow equations we distinguished between two different concepts – the macroscopic and the microscopic field transformations. The macroscopic field transformations are transformations of the variables in terms of which a given effective average action is expressed. They do not change the effective average action and the associated field equations and macroscopic correlation functions. These quantities are only expressed in terms of different “field coordinates”. As familiar from differential equations, a suitable choice of variables can be very helpful to find (approximate) solutions. In quantum field theory the transformation to renormalized fields can absorb part of the microscopic information which does not affect the macroscopic observations. The use of dimensionless renormalized fields reveals the possible existence of scaling solutions. Selecting suitable composite fields can be of great help to make the relevant macroscopic degrees of freedom easily visible.
The advantage of the use of macroscopic field transformations is the explicit control of the transformation at every stage. On the other hand, the fact that the effective average action is not changed may be viewed as a disadvantage. Macroscopic field transformations do not adapt the cutoff to suitable non-linear composite fields. For non-linear field transformations the appearance of connection terms in second functional derivatives may sometimes constitute a further technical disadvantage.
Microscopic field transformations change the effective average action beyond a simple change of variables. This change results from the use of different cutoff functions and different sources employed in the defining functional integral for the effective average action. For non-linear field transformations a non-trivial difference between the field-transformed and the original effective action remains in the limit $k\to 0$ for which all fluctuations are integrated out. The difference in the source terms remains even if the cutoff is removed.
For linear microscopic field transformations the different source terms simply result in a linear macroscopic field transformation. For linear transformations the macroscopic and microscopic field transformations are equivalent. This does no longer hold for non-linear transformations. Macroscopic correlation functions are defined by functional derivatives of the effective action. In the non-linear case there is no simple map between the macroscopic correlation functions with or without the microscopic field transformation. Microscopic field transformations typically loose the direct contact to the microscopic correlation functions of the original model. This contact is maintained for the macroscopic field transformations as long as the transformation to canonical fields is known explicitly.
Microscopic field transformations also affect the initial conditions for the flow at $k\to\infty$ or very large $k$. Relating this “microscopic effective average action”, e.g. $\Gamma_{k}$ for very large $k$, to the “classical action” $S$ used to define the functional integral for the model involves a Jacobian. This Jacobian is the same as the one appearing for a transformation of the field variables in the functional integral. The Jacobian remains field-independent only for linear field transformations. It can therefore be omitted only if at some “initial” or “microscopic” scale $\Lambda$ the composite field equals the fundamental field modulo a linear field transformation.
A simple possible view on the microscopic field transformations associates them to transformations of the integration variables in the functional integral. As long as variable transformations do not depend on $k$ the flow equations are not affected. The correction terms involving the RG-kernel $\gamma[{\varphi}]$ arise from the $k$-dependence of the field transformations. The additional terms can be associated to an additional $k$-dependence resulting from a $k$-dependent Jacobian, combined with a $k$-dependent macroscopic field transformation. The difference of the modified flow equations for non-linear macroscopic and microscopic field transformations results from the Jacobian which depends both on fields and on $k$. This explains why for linear field transformations the macroscopic and microscopic transformations are identical for suitably chosen cutoff functions.
The advantage of the use of microscopic field transformations is the great flexibility if arbitrary non-linear RG-kernels are used. In particular, the cutoff can be adapted to composite fields reflecting the relevant degrees of freedom. Important technical improvements can be achieved by employing the correction terms in order to remove technically annoying terms from the flow equations. The disadvantage of the microscopic field transformations with non-linear RG-kernel is the loss of control of the transformation. It is generally not known which microscopic field transformation produces a given RG-kernel. This relation can be rather complicated. Regular microscopic field transformations impose constraints on the allowed RG-kernels. These constraints are not known in practice, even not approximately, such that solutions of the flow equations obtained for arbitrary $\gamma[{\varphi}]$ are not guaranteed to be valid.
The lack of direct access to the microscopic field transformation makes it hard to follow to which sector of the effective average action a given physical information is moved by the use of arbitrary non-linear $\gamma[{\varphi}]$. This can be rather dangerous for truncated solutions since important information can be moved to the part which is truncated away.
Finally, the initial values for the flow equation encoded in the microscopic effective average action are connected to the classical action defining the functional integral of the model by a relation involving a Jacobian that may be difficult to handle. This disadvantage does not affect the search for scaling solutions and associated universality classes.
For the search for scaling solutions the removal of technical complications by the use of suitable RG-kernels can enlarge the accessible space of truncations. This may contribute to the robustness of scaling solutions found. On the other hand, some of the scaling solutions can be discarded by the truncation and may no longer be visible if arbitrary RG-kernels are employed. The set of scaling solutions found in this way may not be complete. The set of scaling solutions found for the flow equations of the “essential renormalization group” may also contain fake solutions which do not correspond to scaling solutions of the model. Obstructions may be overlooked if the regularity constraints on $\gamma[{\varphi}]$ are discarded. While the set of all scaling solutions for the essential renormalization group may be neither all valid nor complete, there remains a good chance that relevant scaling solutions are found.
The view that couplings which can be removed by the use of suitable RG-kernels are not observable is not justified. Field transformations only move physical information between different sectors. The expression of a given observable in terms of fields and functional derivatives of the effective action is affected by a field transformation.
Both the macroscopic and the microscopic field transformations have their advantages and disadvantages, justifying their use adapted to a given purpose. A setting that combines the advantages and removes disadvantages is the two-field formalism using both fundamental and composite fields. We have not discussed here this formalism in detail and refer to ref. Floerchinger and Wetterich (2009) (or ref. Pawlowski (2007) for similar considerations). The prize to pay for the two-field formalism is the higher complexity due to additional couplings for a description in terms of two fields.
The present paper partly answers the question if the functional flow equation distinguishes between fundamental and composite fields. A first observation states that the correction terms arise only from the $k$-dependence of field transformations. Fields that are related by a $k$-independent field transformation induce identical flow equations. On this level the functional flow equation does not single out a particular choice of fields. We may call fields “canonical” if the classical action is expressed in terms of the canonical fields $\chi^{\prime}(x)$ and the functional measure is linear in $\chi^{\prime}(x)$, $\mathcal{D}\big{[}\chi^{\prime}(x)+a(x)\big{]}=\mathcal{D}\chi^{\prime}(x)$. Fields related to $\chi^{\prime}(x)$ by a non-linear field transformation are no longer canonical. For many purposes one may associate the canonical fields with fundamental fields and call fields obtained by a non-linear field transformation composite. This naming is appropriate if the fundamental fields are the fields in terms of which a model is formulated. We note, however, that there is no fundamental meaning to this concept since for a different choice of the effective action the composite fields may become canonical. On a conceptual level all choices of fields are equivalent.
On the conceptual level there is no difference between canonical and non-canonical fields. Switching to the classical action $\bar{S}[{\varphi}^{\prime}]$ and defining the effective action as $\Gamma_{\varphi}[{\varphi}]$ the non-canonical field becomes canonical. In practice, a given canonical field is often associated with the fundamental field. The non-canonical field is then considered as a composite field. A $k$-dependent field transformation is no longer defined purely on the microscopic level. Fields related to the canonical fundamental field by a $k$-dependent field transformation can always be called “composite fields” or “dressed fields”. In this case the appearance of correction terms in the flow equation indicates “compositeness” of the fields. These dressed fields include the renormalized fields. We infer from this short discussion that the notion of composite fields depends on the context.
We have discussed scalar field theories with a field-dependent kinetial $K$. This serves as a concrete example for the role of non-linear field transformations on different levels and for different versions. In particular we propose that for small values of the scalar field $\chi$ there exist scaling solutions where $K\sim\chi^{-2}$. They are connected to an ultraviolet fixed point with the quantum scale symmetry of a multiplicative rescaling of $\chi$. If this fixed point is confirmed, and if a flow trajectory relates it to the trivial fixed point for $\chi\to\infty$ or $k\to 0$, this setting can lead to an ultraviolet complete scalar quantum field theory with interactions in four dimensions, overcoming the triviality problem.
The most direct evidence for this fixed point comes from the direct computations in app. AB with a field-dependent cutoff respecting the scale symmetry. One may ask if non-linear field transformations can yield additional evidence for the existence of this fixed point. A simple part of the answer is the statement that there exist versions of the effective action and the flow equations for which this fixed point can be firmly established. This is rather trivial, since for these versions one deals with a free theory for composite scalar fields. The interesting question asks if a free theory for composite fields in the ultraviolet can be connected by a flow trajectory to a free theory for fundamental fields in the infrared. For this issue the field transformations have so far not brought substantial additional insight. Nevertheless, the present paper sets the stage for an investigation of this question.
Despite our sometimes critical assessment of the dangers of non-linear field transformations we consider such field transformations as a very rich and powerful tool for the understanding of the physics of a given model. Care is needed, however, in order to keep limitations in mind and to refrain from too general statements.
Acknowledgment
The author would like to thank J. Pawlowski for fruitful discussions.
Appendix A Flow equation for field dependent
kinetial
In this appendix we discuss the flow equation for a field-dependent coefficient of the kinetic term or “kinetial” $K$ in the “leading kinetial approximation”. For this truncation only the term with two derivatives is kept in the effective average action, while the potential and higher derivative terms are omitted. This truncation could be valid in regions of field space for which $K$ is very large and depends on $\chi$. Renormalized couplings for non-derivative interactions scale with inverse powers of $K$ and could therefore only have a small effect. A field-independent $K$ can be absorbed as a wave function renormalization, but interactions due to the $\chi$-dependence of the kinetial do matter. One may check the validity of the leading kinetial approximation by computing the induced potential and investigate its role on the flow of the kinetial $K$. This will be done in the appendix (B).
In the present appendix no field transformations are performed. This direct computation of the flow of the kinetial can serve as a point of comparison for computations using field transformations. We briefly discuss possible scaling solutions in the leading kinetial approximation. In this truncation we find evidence for the ultraviolet fixed point (2). We treat here $\chi$ as the canonical scalar field for which the flow equation (78) holds without corrections. As long as field transformations remain independent of $k$ the same flow equation holds for composite fields as ${\varphi}\sim\ln\chi$. For $k$-independent flow equations the microscopic Jacobian does not influence the flow equation. It matters only for the initial conditions for the solutions for $k\to\infty$.
Flow equation
The flow equation for the truncation (102) can be inferred from the computation in ref. Wetterich (1993c); Seide and Wetterich (1999); v. Gersdorff and Wetterich (2001) by omitting contributions from the potential. We focus on four dimensions, $d=4$, where one finds
$\partial_{t}K(\chi)=$ $\,-\frac{k^{2}}{32\pi^{2}K_{0}^{2}}\left(\frac{\partial K}{\partial\chi}\right)^{2}$
×$\times\left[m_{4}^{8}\left(0,\eta_{K},\frac{K}{K_{0}}\right)-\frac{9}{2}l_{2}^{6}\left(0,\eta_{K},\frac{K}{K_{0}}\right)\right]$
$-\frac{k^{2}}{16\pi^{2}K_{0}}\frac{\partial^{2}K}{\partial\chi^{2}}l_{1}^{4}\left(0,\eta_{K},\frac{K}{K_{0}}\right)\ .$ (162)
Here $K_{0}$ is a type of reference kinetial that will be defined later. It enters the definition of the cutoff function
$\mathcal{R}_{k}(q^{2})=K_{0}k^{2}r(y)\ ,\quad y=\frac{q^{2}}{k^{2}}=\frac{x}{k^{2}}\ ,$ (163)
with “cutoff anomalous dimension” $\eta_{K}$ defined by
$\eta_{K}=-\partial_{t}\ln K_{0}\ .$ (164)
The momentum integrals $l_{1}^{4}$, $l_{2}^{6}$ and $m_{4}^{8}$ are defined as
$l_{1}^{4}\left(0,\eta_{K},\frac{K}{K_{0}}\right)=$ $\,\frac{K_{0}}{2k^{2}}\int_{0}^{\infty}\text{d}x\,x\frac{\partial_{t}\mathcal{R}_{k}(x)}{P^{2}(x)}\ ,$
$l_{2}^{6}\left(0,\eta_{K},\frac{K}{K_{0}}\right)=$ $\,\frac{K_{0}^{2}}{k^{2}}\int_{0}^{\infty}\text{d}x\,x^{2}\frac{\partial_{t}\mathcal{R}_{k}(x)}{P^{3}(x)}\ ,$
$m_{4}^{8}\left(0,\eta_{K},\frac{K}{K_{0}}\right)=$ $\,\frac{2K_{0}^{2}}{k^{2}}\int_{0}^{\infty}\text{d}x\,x^{4}\bigg{[}\frac{\partial_{t}\mathcal{R}_{k}(x)}{P^{5}(x)}\big{(}\partial_{x}P(x)\big{)}^{2}$
$-\frac{1}{2P^{4}(x)}\partial_{x}P(x)\partial_{t}\partial_{x}\mathcal{R}_{k}\bigg{]}\ .$ (165)
They involve the inverse propagator in the presence of the cutoff
$P(x)=Kx+\mathcal{R}_{k}=K_{0}k^{2}\big{(}\tilde{z}y+r(y)\big{)}\ .$ (166)
Here we employ for the “normalized kinetial” the ratio
$\tilde{z}=\frac{K}{K_{0}}\ .$ (167)
Threshold functions
The momentum integrals (often called threshold functions) carry key information of the flow equation. Using
$\partial_{t}\mathcal{R}_{k}=K_{0}k^{2}\left(2-\eta-2\frac{\partial\ln r}{\partial\ln y}\right)r$ (168)
the factors $K_{0}^{p}$ cancel such that for a given choice of $r(y)$ these dimensionless integrals depend only on $\eta$ and $\tilde{z}$. (We omit the index on $\eta$ from now on.) The powers of $k$ drop out if we switch to dimensionless integration variables $y=x/k^{2}$,
$l_{1}^{4}(0,\eta,\tilde{z})=$ $\,\frac{1}{2}\int_{0}^{\infty}\text{d}y\,y(\tilde{z}y+r)^{-2}\big{[}(2-\eta)r-2y\partial_{y}r\big{]}\ ,$
$l_{2}^{6}(0,\eta,\tilde{z})=$ $\,\int_{0}^{\infty}\text{d}y\,y^{2}(\tilde{z}y+r)^{-3}\big{[}(2-\eta)r-2y\partial_{y}r\big{]}$
$=$ $\,-\frac{\partial}{\partial\tilde{z}}l_{1}^{4}(0,\eta,\tilde{z})\ ,$
$m_{4}^{8}(0,\eta,\tilde{z})=$  $\,2\int_{0}^{\infty}\text{d}y\,y^{4}(\tilde{z}y+r)^{-5}(\tilde{z}+\partial_{y}r)^{2}$
×$\qquad\times\big{[}(2-\eta)r-2y\partial_{y}r\big{]}$
$+2\int_{0}^{\infty}\text{d}$ $y\,y^{5}(\tilde{z}y+r)^{-4}(\tilde{z}+\partial_{y}r)\big{(}\partial_{y}^{2}r+\frac{\eta}{2y}\partial_{y}r\big{)}\ .$ (169)
As an example, we choose the Litim cutoff function Litim (2001)
$r(y)=(1-y)\theta(1-y)\ ,$ (170)
for which ($\tilde{z}\neq 0$)
$l_{1}^{4}(0,\eta,\tilde{z})=\frac{1}{2}\int_{0}^{1}\text{d}y\,y\big{(}1+(\tilde{z}-1)y\big{)}^{-2}\big{(}2-\eta(1-y)\big{)}$ (171)
The Litim cutoff needs to be defined as the limiting case of smooth functions in order to make $\partial_{y}^{2}r\partial_{y}r$ well defined. For our choice one finds
$m_{4}^{8}(0,\eta,\tilde{z})=$ $\,-\frac{1}{6}(\tilde{z}-1)^{2}\frac{\partial^{3}l_{1}^{4}(0,\eta,\tilde{z})}{\partial\tilde{z}^{3}}+\frac{2}{\tilde{z}^{3}}-\frac{1}{2\tilde{z}^{4}}$
$\,-\eta(\tilde{z}-1)\int_{0}^{1}\text{d}y\,y^{4}\big{[}1+(\tilde{z}-1)y\big{]}^{-4}\ .$ (172)
Evaluating the integral yields
$l_{1}^{4}(0,\eta,\tilde{z})=$ $\,\frac{\ln\tilde{z}}{(\tilde{z}-1)^{2}}-\frac{1}{\tilde{z}(\tilde{z}-1)}$
$-\frac{\eta}{(\tilde{z}-1)^{2}}\left(\frac{1}{2}\ln\tilde{z}-1+\frac{1}{\tilde{z}-1}\ln\tilde{z}\right)\ .$ (173)
For $\tilde{z}=1$ the integrals simplify
$\hat{l}_{1}^{4}(\eta)=$ $\,l_{1}^{4}(0,\eta,\tilde{z}=1)=\frac{1}{2}-\frac{\eta}{12}\ ,$
$\hat{l}_{2}^{6}(\eta)=$ $\,l_{2}^{6}(0,\eta,\tilde{z}=1)=\frac{2}{3}-\frac{\eta}{12}\ ,$
$\hat{m}_{4}^{8}(\eta)=$ $\,m_{4}^{8}(0,\eta,\tilde{z}=1)=\frac{3}{2}\ .$ (174)
For $\eta=0$ one obtains the explicit expressions
$\bar{l}_{1}^{4}(\tilde{z})=$ $\,l_{1}^{4}(0,0,\tilde{z})=\frac{\ln\tilde{z}}{(\tilde{z}-1)^{2}}-\frac{1}{\tilde{z}(\tilde{z}-1)}\ ,$ (175)
$\bar{l}_{2}^{6}(\tilde{z})=$ $\,l_{2}^{6}(0,0,\tilde{z})=\frac{2\ln\tilde{z}}{(\tilde{z}-1)^{3}}-\frac{3\tilde{z}-1}{\tilde{z}^{2}(\tilde{z}-1)^{2}}\ ,$
$\bar{m}_{4}^{8}(\tilde{z})=$ $\,m_{4}^{8}(0,0,\tilde{z})=\frac{4\ln\tilde{z}}{(\tilde{z}-1)^{3}}-\frac{38\tilde{z}^{3}-19\tilde{z}^{2}+8\tilde{z}-3}{6\tilde{z}^{4}(\tilde{z}-1)^{2}}\ .$
For gaining some intuition on the dependence on $\eta$ we also provide the integrals for $\eta=2$,
$l_{1}^{4}(0,2,\tilde{z})=$ $\,\frac{\tilde{z}+1}{\tilde{z}(\tilde{z}-1)^{2}}-\frac{2\ln\tilde{z}}{(\tilde{z}-1)^{3}}\ ,$
$l_{2}^{6}(0,2,\tilde{z})=$ $\,\tilde{z}^{-2}(\tilde{z}-1)^{-3}\left\{2\tilde{z}^{2}+5\tilde{z}-1-\frac{6\tilde{z}^{2}}{\tilde{z}-1}\ln\tilde{z}\right\}\ .$ (176)
There are no poles in these integrals for $\tilde{z}=1$, with a Taylor expansion around $\tilde{z}=1$ yielding the leading values (A).
For large $\tilde{z}$ the effect of the IR-cutoff is suppressed by powers of $\tilde{z}^{-1}$. In this range of $\tilde{z}$ we can scale out the dominant $\tilde{z}$-factors by changing the integration variable to $\tilde{y}=\tilde{z}y$,
$l_{1}^{4}(0,\eta,\tilde{z})=$ $\,\tilde{z}^{-2}\tilde{l}_{1}^{4}(\tilde{z},\eta)\ ,$
$l_{2}^{6}(0,\eta,\tilde{z})=$ $\,\tilde{z}^{-3}\tilde{l}_{2}^{6}(\tilde{z},\eta)\ ,$
$m_{4}^{8}(0,\eta,\tilde{z})=$ $\,\tilde{z}^{-3}\tilde{m}_{4}^{8}(\tilde{z},\eta)\ .$ (177)
Here $\tilde{l}_{1}^{4}(\tilde{z},\eta)$ obtains from $l_{1}^{4}(0,\eta,\tilde{z})$ by setting in eq. (A) $\tilde{z}\to 1$, $y\to\tilde{y}$, $r(y)\to\tilde{r}(\tilde{y})$, and similar for the other functions. The $\tilde{z}$-dependence of $\tilde{l}_{1}^{4}(\tilde{z},\eta)$ arises only through the argument $\tilde{r}(\tilde{y})=r\big{(}y(\tilde{y})\big{)}=r(\tilde{y}/\tilde{z})$. Comparison with the full result (171) shows a logarithmic dependence of $\tilde{l}_{1}^{4}$ on $\tilde{z}$. For $\tilde{l}_{2}^{6}$ one has the relation
$\tilde{l}_{2}^{6}(\tilde{z},\eta)=\left(2-\frac{\partial}{\partial\ln\tilde{z}}\right)\tilde{l}_{1}^{4}(\tilde{z},\eta)\ .$ (178)
The suppression of the threshold functions slows down the flow. This effect of “kinetic decoupling” arises since for $\tilde{z}$ increasing to large values the effect of the infrared cutoff becomes less and less important.
Field dependent cutoff
We next discuss a few characteristic solutions of the flow equation for the kinetial. Besides the dependence on the choice of the shape of the cutoff function $r(y)$ they also depend on the choice of $K_{0}$. We discuss here both a field-dependent and a field-independent choice of $K_{0}$.
Let us first consider the choice $K_{0}=K(\chi)$. This is often used effectively in the “background formalism”. Using a cutoff function depending on the macroscopic fields induces correction terms in the flow equation Wetterich (2018, 2021, 2024). These correction terms can be computed in the setting of the simplified flow equation Wetterich (2024). They have been found to be small and consistent with the relevant symmetries. We omit these correction terms here. The choice $K_{0}=K$ brings important technical simplifications since we can use $\tilde{z}=1$ in the threshold functions. We write eq. (A) in the form
$\partial_{t}K=$ $\,-\frac{k^{2}}{32\pi^{2}}\bigg{\{}2\frac{\partial^{2}\ln K}{\partial\chi^{2}}\hat{l}_{1}^{4}(\eta)$
$+\left(\frac{\partial\ln K}{\partial\chi}\right)^{2}\left[2\hat{l}_{1}^{4}(\eta)+\hat{m}_{4}^{8}(\eta)-\frac{9}{2}\hat{l}_{2}^{6}(\eta)\right]\bigg{\}}\ ,$ (179)
with $\hat{l}_{1}^{4}$ etc. given for the Litim cutoff by eq. (A). For $K_{0}=K$ and $K$ a power of $\chi$ the r.h.s. of eq. (A) or (A) scales $\sim\beta^{-2}$ for $\chi\to\beta\chi$.
For $\eta$ independent of $\chi$ we can make the ansatz of a powerlaw
$K(\chi)=c(k)\chi^{-\alpha}\ ,$ (180)
for which eq. (A) becomes
$\partial_{t}c\,\chi^{-\alpha}=-\frac{k^{2}}{32\pi^{2}}\chi^{-2}\left[2\alpha\hat{l}_{1}^{4}+\alpha^{2}\left(2\hat{l}_{1}^{4}+\hat{m}_{4}^{8}-\frac{9}{2}\hat{l}_{2}^{6}\right)\right]\ .$ (181)
For $\eta$ independent of $\chi$ the $\chi$-dependence drops out for $\alpha=0$ and $\alpha=2$. For $\alpha=0$ we simply have a free field theory with $K$ independent of $k$ or $\chi$. For $\alpha=2$ one finds
$\partial_{t}c=\frac{k^{2}}{32\pi^{2}}D(\eta)\ ,\quad\eta=-\partial_{t}\ln c\ ,$ (182)
with $D(\eta)$ given for the Litim cutoff by
$D(\eta)=-\big{(}12\hat{l}_{1}^{4}+4\hat{m}_{4}^{8}-18\hat{l}_{2}^{6}\big{)}=-\frac{\eta}{2}\ .$ (183)
Eq. (182) has two solutions. For the first, $\partial_{t}c=0$, $\eta=0$ we find the fixed point solution (2). This is in agreement with the vanishing of the loop correction $H=0$ in the saddlepoint approximation (33), (35). The second solution
$c=\frac{k^{2}}{64\pi^{2}}\ ,\quad\eta=-2\ ,$ (184)
results in a scaling form
$\Gamma_{k}=\frac{k^{2}}{128\pi^{2}}\int_{x}\frac{1}{\chi^{2}}\partial^{\mu}\chi\partial_{\mu}\chi\ .$ (185)
Both solutions are invariant under global dilatation or scale transformations $\chi\to\beta\chi$. For these solutions the kinetial diverges for $\chi\to 0$. The region of small $\chi$ is a candidate for the validity of the leading kinetial approximation. We will discuss this in appendix B.
Field independent cutoff
The solutions of the flow equation become more complex if we take $K_{0}=K_{0}(k)$ to be independent of $\chi$. In this case the infrared cutoff term does not involve the macroscopic field $\chi$ and there are no corrections to the flow equation. On the other hand, $\tilde{z}=K(\chi)/K_{0}$ becomes now a field- and $k$-dependent function. A cutoff with constant $K_{0}$ is no longer invariant under dilatations, such that quantum scale invariant solutions of the type (181) are at best an approximation. Of course, the free field solution $K=K_{0}$ continues to exist. For $\chi$-independent $K_{0}$ we translate the flow equation (A) into a flow equation for $\tilde{z}$,
$\partial_{t}\tilde{z}=$ $\,\eta\tilde{z}-\frac{k^{2}}{32\pi^{2}K_{0}}\bigg{\{}2\frac{\partial^{2}\tilde{z}}{\partial\chi^{2}}l_{1}^{4}(0,\eta,\tilde{z})$
$+\left(\frac{\partial\tilde{z}}{\partial\chi}\right)^{2}\left(m_{4}^{8}(0,\eta,\tilde{z})-\frac{9}{2}l_{2}^{6}(0,\eta,\tilde{z})\right)\bigg{\}}\ .$ (186)
We can set $K_{0}$ to an arbitrary constant by a multiplicative rescaling of $\chi$.
We first ask if there are “static solutions” with $k$-independent $K(\chi)$ and $K_{0}$. This requires the r.h.s. of eq. (A) to vanish, or
$2\frac{\partial^{2}\tilde{z}}{\partial\chi^{2}}\bar{l}_{1}^{4}(\tilde{z})+\left(\frac{\partial\tilde{z}}{\partial\chi}\right)^{2}\left(\bar{m}_{4}^{8}(\tilde{z})-\frac{9}{2}\bar{l}_{2}^{6}(\tilde{z})\right)=0\ .$ (187)
For the Litim cutoff we employ eq. (175) in order to obtain an explicit non-linear second order differential equation for $\tilde{z}(\chi)$,
$2\left(\frac{\ln\tilde{z}}{(\tilde{z}-1)^{2}}-\frac{1}{\tilde{z}(\tilde{z}-1)}\right)\frac{\partial^{2}\tilde{z}}{\partial\chi^{2}}$
$=$ $\,\left[\frac{5\ln\tilde{z}}{(\tilde{z}-1)^{3}}-\frac{13\tilde{z}^{2}-5\tilde{z}+3}{6\tilde{z}^{4}(\tilde{z}-1)}\right]\left(\frac{\partial\tilde{z}}{\partial\chi}\right)^{2}\ .$ (188)
In particular, for $\tilde{z}=1$ one finds
$\frac{\partial^{2}\tilde{z}}{\partial\chi^{2}}(\tilde{z}=1)=\frac{3}{2}\left(\frac{\partial\tilde{z}}{\partial\chi}(\tilde{z}=1)\right)^{2}\ ,$ (189)
with approximate solution near $\tilde{z}=1$ given for $\chi$ near $\chi_{1}$ by
$\tilde{z}(\chi)\approx 1-\frac{2}{3}\ln\left(\frac{\chi}{\chi_{1}}\right)\ ,\quad\tilde{z}(\chi_{1})=1\ ,$ (190)
or, alternatively, simply $\tilde{z}=1$.
For $\tilde{z}\to\infty$ the leading logarithmic term cancels for
$\frac{\partial^{2}\tilde{z}}{\partial\chi^{2}}=\frac{5}{2\tilde{z}}\left(\frac{\partial\tilde{z}}{\partial\chi}\right)^{2}\ ,$ (191)
with approximate solution
$\tilde{z}(\chi)\approx\left(\frac{\chi_{0}}{\chi}\right)^{\frac{2}{3}}\ .$ (192)
This limit should describe the limit of the static solution for $\chi\to 0$.
We conclude that the static solutions for a field-independent IR-cutoff differ substantially from the ones for a field-dependent cutoff with $K_{0}=K$, where $K\sim\chi^{-2}$ according to eq. (2). The origin of this difference arises from the fact that the field-independent cutoff is “inhomogeneous” in the sense that the relative weight as compared to $\Gamma_{k}^{(2)}$ depends strongly on the value of $\tilde{z}$. For large $\tilde{z}$ the cutoff becomes subdominant in the combined inverse propagator $P(x)$, rendering it relatively ineffective. In contrast, for $\tilde{z}\ll 1$ the IR-cutoff dominates $P(x)$ for momenta $q^{2}\approx k^{2}$. Furthermore, only the field-dependent cutoff respects the scale symmetry $\chi\to\beta\chi$ which is a crucial property of the fixed point (2). For an investigation of this possible fixed point the field-dependent cutoff seems clearly more appropriate.
Scaling solution for scale invariant fields
One is often interested in scaling solutions for which $K$ becomes a function of the “scaling field” or “scale invariant field”
$\tilde{\chi}=\frac{\chi}{k}\ ,$ (193)
without any further $k$-dependence. For a field-dependent cutoff $K_{0}=K$ the solution (185) is of this type. In this case the effective action can be written in terms of scale invariant fields if one also uses a scale invariant metric $\tilde{g}_{\mu\nu}=k^{2}g_{\mu\nu}$. In contrast, the effective action (2) only involves the scale invariant field $\tilde{\chi}$ if $\tilde{g}_{\mu\nu}=g_{\mu\nu}$.
We also may investigate scale invariant solutions for a field-independent cutoff with constant $K_{0}$. The flow equation for $\tilde{z}$ at fixed $\tilde{\chi}$ reads
$\partial_{t}\tilde{z}|_{\tilde{\chi}}=\partial_{t}\tilde{z}|_{\chi}-\frac{\partial\tilde{z}}{\partial\tilde{\chi}}\partial_{t}\tilde{\chi}|_{\chi}=\partial_{t}\tilde{z}|_{\chi}+\tilde{\chi}\frac{\partial\tilde{z}}{\partial\tilde{\chi}}\ .$ (194)
With primes denoting derivatives with respect to $\tilde{\chi}$ this yields
$\partial_{t}\tilde{z}=\eta\tilde{z}+\tilde{\chi}\tilde{z}^{\prime}-\frac{1}{32\pi^{2}K_{0}}\left\{2\tilde{z}^{\prime\prime}l_{1}^{4}+\big{(}\tilde{z}^{\prime}\big{)}^{2}\left(m_{4}^{8}-\frac{9}{2}l_{2}^{6}\right)\right\}\ .$ (195)
(By multiplicative rescaling of $\tilde{\chi}$ one may set $232\pi^{2}K_{0}=$ 2.) There is no longer any explicit dependence of the flow equation on the renormalization scale $k$. The effective action becomes
$\Gamma_{k}[\tilde{\chi}]=\frac{k^{2}K_{0}}{2}\int_{x}\tilde{z}(\tilde{\chi})\partial^{\mu}\tilde{\chi}\partial_{\mu}\tilde{\chi}\ .$ (196)
For a scaling solution $K_{0}$ is constant, $\eta=0$, $\partial_{t}\tilde{z}=0$, and
$\tilde{\chi}\tilde{z}^{\prime}=\frac{1}{32\pi^{2}K_{0}}\left\{2\tilde{z}^{\prime\prime}\bar{l}_{1}^{4}(\tilde{z})+\big{(}\tilde{z}^{\prime}\big{)}^{2}\left(\bar{m}_{4}^{8}(\tilde{z})-\frac{9}{2}\bar{l}_{2}^{6}(\tilde{z})\right)\right\}\ .$ (197)
With eq. (175) this reads
$32\pi^{2}K_{0}\tilde{\chi}\tilde{z}^{\prime}=$ $\,-\big{(}\tilde{z}^{\prime}\big{)}^{2}\bigg{\{}5\frac{\ln\tilde{z}}{(\tilde{z}-1)^{3}}-\frac{13\tilde{z}^{2}-5\tilde{z}+3}{6\tilde{z}^{4}(\tilde{z}-1)}\bigg{\}}$
$+\tilde{z}^{\prime\prime}\left\{\frac{2\ln\tilde{z}}{(\tilde{z}-1)^{2}}-\frac{2}{\tilde{z}(\tilde{z}-1)}\right\}\ .$ (198)
Neglecting the l.h.s. for $\tilde{\chi}\to 0$ one finds in this limit the constant solution $\tilde{z}=\text{const.}$., or the behavior
$\tilde{z}=a\tilde{\chi}^{-\frac{2}{3}}\ .$ (199)
Again, the degree of divergence of the scaling solution for $\tilde{\chi}\to 0$ differs from the behavior $\sim\tilde{\chi}^{-2}$ found in eq. (185) for the field dependent cutoff, reflecting the shortcoming of the inhomogeneity of the cutoff.
The region of small fields $\tilde{\chi}\to 0$ for solutions with diverging kinetial could be a candidate for the validity of the leading kinetial approximation. As for any truncation, one may investigate its validity by computing the flow of parts in the effective action that have been neglected, and assess their influence on the flow of the terms retained. In our case, this concerns in particular the neglected scalar potential which will be included in appendix B.
Appendix B Flow of kinetial and potential
In this appendix we investigate the functional flow for a single scalar field in leading order in the derivative expansion. This truncation includes two field-dependent functions, namely the potential and kinetial. This approximation has been studied in refs. Seide and Wetterich (1999); v. Gersdorff and Wetterich (2001) and employed frequently afterwards. What is new in this appendix is the focus on possible field regions for which the kinetial diverges. This divergence is not per se a shortcoming since it may be removed by a different choice of the field variable. In particular, we are interested in possible scaling solutions which could correspond to an ultraviolet fixed point defining an ultraviolet complete non-trivial theory.
In first order in a derivative expansion we need to determine the two dimensionless functions $u^{\prime}(\tilde{\chi})$ and $z(\tilde{\chi})$. Here $u(\tilde{\chi})=U(\tilde{\chi})/k^{4}$ is the dimensionless scalar potential and primes denote derivatives with respect to $\tilde{\chi}=\chi/k$. In this appendix we omit the tilde on $z\equiv\tilde{z}$.
We will find that the leading kinetial approximation discussed in appendix A is consistent for a field-dependent cutoff $K_{0}=K$. In contrast, the leading kinetial approximation fails for a field-independent cutoff. This could be taken as an additional point in favor of a field-dependent cutoff.
Field independent cutoff
The flow equation for the derivative of the potential at fixed $\tilde{\chi}$ reads for $K_{0}=1$, $\eta=0$
$\partial_{t}u^{\prime}=-3u^{\prime}+\tilde{\chi}u^{\prime\prime}-\frac{1}{16\pi^{2}}\left\{z^{\prime}l_{1}^{6}+u^{\prime\prime\prime}l_{1}^{4}\right\}\ .$ (200)
Similarly, the evolution of $z$ is given by
$\partial_{t}z=$ $\,\tilde{\chi}z^{\prime}-\frac{1}{32\pi^{2}}\bigg{\{}2z^{\prime\prime}l_{1}^{4}+(z^{\prime})^{2}\bigg{(}m_{4}^{8}-\frac{9}{2}l_{2}^{6}\bigg{)}$
$+(u^{\prime\prime\prime})^{2}m_{4}^{4}+u^{\prime\prime\prime}z^{\prime}\big{(}2m_{4}^{6}-4l_{2}^{4}\big{)}\bigg{\}}\ .$ (201)
The additional terms from the scalar potential involve the third derivative of $u$.
The threshold functions are modified in the presence of a potential. They are defined for $n\geq 1$ by
$l_{n}^{d}\equiv$ $\,l_{n}^{d}(u^{\prime\prime},z)$
$=$ $\,n\int_{0}^{\infty}\text{d}y\,y^{\frac{d}{2}-1}\big{(}zy+r+u^{\prime\prime}\big{)}^{-(n+1)}\big{(}r-y\partial_{y}r\big{)}\ ,$ (202)
and
$m_{n}^{d}\equiv$ $\,m_{n}^{d}(u^{\prime\prime},z)=n\int_{0}^{\infty}\text{d}y\,y^{\frac{d}{2}}\big{(}zy+r+u^{\prime\prime}\big{)}^{-(n+1)}$
×$\qquad\qquad\qquad\times\big{(}z+\partial_{y}r\big{)}^{2}\big{(}r-y\partial_{y}r\big{)}$
$+2\int_{0}^{\infty}\text{d}y\,y^{\frac{d}{2}+1}\big{(}zy+r+u^{\prime\prime}\big{)}^{-n}\big{(}z+\partial_{y}r\big{)}\partial_{y}^{2}r\ .$ (203)
The functions $l_{n}^{d}$ obey the recursive relations
$l_{n+1}^{d}(w,z)=$ $\,-(n+\delta_{n,0})^{-1}\frac{\partial}{\partial w}l_{n}^{d}(w,z)\ ,$
$l_{n+1}^{d+2}(w,z)=$ $\,-(n+\delta_{n,0})^{-1}\frac{\partial}{\partial z}l_{n}^{d}(w,z)\ ,$ (204)
with
$l_{0}^{d}(w,z)=\int\text{d}y\,y^{\frac{d}{2}-1}\big{(}zy+r+u^{\prime\prime}\big{)}^{-1}\big{(}r-y\partial_{y}r\big{)}\ .$ (205)
We observe the universal property
$l_{n}^{2n}(0,z)=z^{-n}\ .$ (206)
Similarly, one has for $n\geq 1$
$m_{n+1}^{d}(w,z)=-\frac{1}{n}\frac{\partial}{\partial w}m_{n}^{d}(w,z)\ .$ (207)
The threshold functions depend on the choice of the infrared cutoff encoded in $r(y)$.
For the Litim cutoff Litim (2001),
$r(y)=(1-y)\theta(1-y)\ ,$ (208)
one finds
$l_{0}^{d}(w,z)=\int_{0}^{1}\text{d}y\,y^{\frac{d}{2}-1}\big{[}(z-1)y+1+w\big{]}^{-1}\ ,$ (209)
or
$l_{0}^{4}(w,z)=\frac{1}{z-1}\bigg{\{}1-\frac{1+w}{z-1}\ln\bigg{(}1+\frac{z-1}{1+w}\bigg{)}\bigg{\}}\ .$ (210)
Expanding in small $(z-1)/(1+w)$ yields
$l_{0}^{4}(w,z)=\frac{1}{2(1+w)}-\frac{z-1}{3(1+w)^{2}}+\dots$ (211)
The threshold functions $l_{n}^{d}$ involved in eqs. (200), (B) follow from $l_{0}^{4}$ by use of eq. (B). For the Litim cutoff one finds
$m_{4}^{d}(w,z)=$  $\,4(z-1)^{2}\int_{0}^{1}\text{d}y\,y^{\frac{d}{2}}\big{[}(z-1)y+1+w\big{]}^{-5}$
$+\frac{2z-\frac{1}{2}}{(z+w)^{4}}$
$=$ $\,(z-1)^{2}l_{4}^{d+2}(w,z)+\frac{2z-\frac{1}{2}}{(z+w)^{4}}\ .$ (212)
A simple solution of the coupled system of differential equations (200), (B) is a free massless scalar field with constant $z$,
$u^{\prime}=0\ ,\quad z=z_{0}\ .$ (213)
Kinetic decoupling
For large $z\gg 1+w$ the threshold functions vanish with inverse powers of $z$. In the limit of this “kinetic decoupling” we evaluate
$l_{1}^{4}=$ $\,\frac{\ln z-1}{z^{2}}\ ,\quad l_{2}^{4}=l_{1}^{6}=\frac{1}{z^{2}}\ ,\quad l_{2}^{6}=\frac{2\ln z-3}{z^{3}}\ ,$
$m_{4}^{8}=$ $\,\frac{12\ln z-19}{3z^{3}}\ ,\quad m_{4}^{6}=\frac{1}{z^{2}}\ ,\quad m_{4}^{4}=\frac{1}{3z}\ .$ (214)
In this limit the flow equations become
$\partial_{t}u^{\prime}=$ $\,-3u^{\prime}+\tilde{\chi}u^{\prime\prime}-\frac{1}{16\pi^{2}z}\bigg{\{}\frac{z^{\prime}}{z}+\frac{(\ln z-1)u^{\prime\prime\prime}}{z}\bigg{\}}\ ,$
$\partial_{t}z=$ $\,\tilde{\chi}z^{\prime}-\frac{1}{32\pi^{2}z}\bigg{\{}2(\ln z-1)\frac{z^{\prime\prime}}{z}-\bigg{(}5\ln z-\frac{43}{6}\bigg{)}\bigg{(}\frac{z^{\prime}}{z}\bigg{)}^{2}$
$\qquad+\frac{1}{3}(u^{\prime\prime\prime})^{2}-2u^{\prime\prime\prime}\frac{z^{\prime}}{z}\bigg{\}}\ .$ (215)
In order to include a possible singular behavior for $z\to\infty$ we rewrite these equations in terms of the variable
$s=\frac{1}{z}\ ,$ (216)
resulting in
$\partial_{t}u^{\prime}=$ $\,-3u^{\prime}+\tilde{\chi}u^{\prime\prime}+\frac{1}{16\pi^{2}}\big{\{}s^{\prime}+(\ln s+1)s^{2}u^{\prime\prime\prime}\big{\}}\ ,$
$\partial_{t}s^{\prime}=$ $\,\tilde{\chi}s^{\prime}+\frac{1}{32\pi^{2}}\bigg{\{}2(\ln s+1)s^{2}s^{\prime\prime}+\bigg{(}\ln s+\frac{19}{6}\bigg{)}s(s^{\prime})^{2}$
$\qquad+\frac{s^{3}}{3}(u^{\prime\prime\prime})^{2}+2s^{2}s^{\prime}u^{\prime\prime\prime}\bigg{\}}\ .$ (217)
The region of interest concerns $s\ll 1$.
Scaling solutions
For scaling solutions the l.h.s. of the flow equation (B) vanishes, $\partial_{t}u^{\prime}=0$, $\partial_{t}s^{\prime}=0$. There are two simple solutions. The first corresponds to a free massless scalar field with constant $s$,
$u^{\prime}=0\ ,\quad s=s_{0}\ ,$ (218)
corresponding to eq. (213). The second corresponds to
$s=0\ ,\quad u=u_{0}+\frac{\lambda}{8}\tilde{\chi}^{4}.$ (219)
This corresponds to a massless scalar field with a quartic interaction. The fluctuation effects are switched off due to perfect kinetic decoupling for $s=0$. In view of the diverging kinetial the renormalized quartic coupling $\lambda_{R}\sim\lambda/z^{2}=\lambda s^{2}$ vanishes.
We are interested in possible scaling solutions which interpolate between the fixed points (218) and (219). Let us first consider a region of $\chi$ around some small $\bar{\chi}$ for which $s(\bar{\chi})$ is also small. We approximate $\ln s$ by a constant $\ln\bar{s}$, $\bar{s}=s(\bar{\chi})$. In this approximation we obtain the scaling solution
$s=a\tilde{\chi}\ ,\quad u=u_{0}+\frac{a\tilde{\chi}}{48\pi^{2}}+\frac{\lambda\tilde{\chi}^{4}}{8}\ ,$ (220)
where
$a^{2}=-\frac{32\pi^{2}}{\ln\bar{s}+19/6}\ .$ (221)
This approximation requires small enough $\bar{s}$ such that $\ln\bar{s}+19/6<0$, $a^{2}>0$. For positive $\tilde{\chi}$ we take $a>0$, for negative $\tilde{\chi}$ one chooses $a<0$. Then $s$ is always positive, while the approximate solution is not analytic at $\tilde{\chi}=0$. Besides the approximation for $\ln s$ we have neglected terms which lead to relative corrections $\sim a^{2}\lambda\tilde{\chi}^{3}$.
This approximate solution motivates for $\tilde{\chi}\to 0$ the ansatz
$s=4\pi\tilde{\chi}g\ ,$ (222)
for which the scaling solution has to obey the differential equations
$u^{\prime}=\frac{g+\tilde{\chi}g^{\prime}}{12\pi}+\frac{\tilde{\chi}}{3}\bigg{[}u^{\prime\prime}+\tilde{\chi}g^{2}\big{(}\ln(4\pi\tilde{\chi})+\ln g+1\big{)}u^{\prime\prime\prime}\bigg{]}\ ,$ (223)
and
$g+\tilde{\chi}g^{\prime}=$ $-\frac{1}{2}g(g+\tilde{\chi}g^{\prime})^{2}\bigg{(}\ln(4\pi\tilde{\chi})+\ln g+\frac{19}{6}\bigg{)}$
$-\tilde{\chi}g^{2}(2g^{\prime}+\tilde{\chi}g^{\prime\prime})(\ln(4\pi\tilde{\chi})+\ln g+1)$
$-\tilde{\chi}g^{2}(g+\tilde{\chi}g^{\prime})u^{\prime\prime\prime}-\frac{1}{6}\tilde{\chi}^{2}g^{3}(u^{\prime\prime\prime})^{2}\ .$ (224)
It is not clear at the present stage if a satisfactory solution with smoothly varying $g$ exists. We also do not know to what extent such a solution would depend on the precise choice of the cutoff function.
The general issue for solutions with $s$ non analytic for $\tilde{\chi}\to 0$ concerns the term $s^{\prime}/(16\pi^{2})$ in the flow equation (B) for $u^{\prime}$. Even a rather mild divergence of $s^{\prime}$ (e.g. logarithmic) typically induces a similar divergence in $u^{\prime}$. Such a divergence results in a much stronger divergence of $u^{\prime\prime\prime}$, which, in turn, can strongly affect the evolution of $s$. If this effect is too strong the leading kinetial approximation will fail.
For a possible scaling $z\sim\tilde{\chi}^{-\alpha}$, $s\sim\tilde{\chi}^{\alpha}$ as discussed in appendix A the leading term for a possible scaling solution obeys
$u^{\prime}-\frac{1}{3}\tilde{\chi}u^{\prime\prime}=\frac{s^{\prime}}{48\pi^{2}}\ .$ (225)
For $s=48\pi^{2}c\tilde{\chi}^{\alpha}$ one finds
$u^{\prime}=\frac{3c\alpha}{4-\alpha}\tilde{\chi}^{\alpha-1}\ ,$ (226)
and therefore
$u^{\prime\prime\prime}=\frac{3c\alpha(\alpha-1)(\alpha-2)}{4-\alpha}\tilde{\chi}^{\alpha-3}\ .$ (227)
This leads to contributions on the r.h.s. proportional to $\tilde{\chi}^{5\alpha-6}$ or $\tilde{\chi}^{4(\alpha-1)}$, unless $\alpha=1$ or $\alpha=2$. For $\alpha=2/3$ as discussed in appendix A for a field-independent cutoff these terms diverge and overwhelm the terms retained in the leading kinetial approximation. One concludes that the leading kinetial approximation is not valid for a field-independent cutoff. This issue would change for the particular case $z\sim\tilde{\chi}^{-2}$, $s\sim\tilde{\chi}^{2}$, $\alpha=2$ for which one obtains $u^{\prime}\sim\tilde{\chi}$, $u^{\prime\prime\prime}=0$.
Field dependent cutoff
The situation is different for a suitable field-dependent cutoff. If one employs a cutoff function proportional to the kinetial $K(\chi)$ the flow equation for the effective potential becomes
$\partial_{t}U=\frac{1}{2}\int_{q}\partial_{t}\big{(}K(\chi)k^{2}r\big{)}\big{(}\Gamma^{(2)}+K(\chi)k^{2}r\big{)}^{-1}\ .$ (228)
For a truncation in first order of a derivative expansion one has for constant $\chi$
$\Gamma^{(2)}=K(\chi)q^{2}+\partial_{\chi}^{2}U(\chi)\ .$ (229)
With $\eta=-\partial_{t}\ln K(\chi)$ this yields
$\partial_{t}U=\int_{q}\left(1-\frac{\eta}{2}+\frac{1}{2}\partial_{t}\ln r\right)\big{(}q^{2}+K^{-1}\partial_{\chi}^{2}U\big{)}^{-1}\ .$ (230)
For $\partial_{\chi}^{2}U=0$ and $\partial_{\chi}\eta=0$ the flow of $U$ vanishes up to a $\chi$-independent constant, $\partial_{t}\partial_{\chi}U=0$. A flat potential remains flat for all $K$ which are proportional to a power of $k$,
$K=f(\chi)k^{-\eta}\ .$ (231)
In this case the leading kinetial approximation may be a reasonable approximation.
The fixed point solution (2) obeys both the flow equation for $K$ and $U$. The flow equation (A) remains valid, admitting also the scaling solution (185) with $\eta=-2$. For both scaling solutions the effective action is invariant under the rescaling of $\chi$ by a constant factor $\beta$. They realize a version of quantum scale symmetry. This version of quantum scale symmetry, for which the metric (or the coordinates) are kept fixed, indeed requires a flat potential. It is respected by the field-dependent cutoff, but not by a field-independent cutoff. This could explain why no equivalent scaling solution is found for the field-independent cutoff.
The possible scaling solutions,
$\Gamma_{k}=\int_{x}\left\{\frac{\kappa^{2}}{2\chi^{2}}\partial^{\mu}\chi\partial_{\mu}\chi+u_{0}k^{4}\right\}\ ,$ (232)
with $\kappa=\kappa_{0}$ or $\kappa=ck^{2}$, describe a free massless composite scalar field. It may not be equivalent to the trivial scaling solution where $\kappa/\chi^{2}$ is replaced by one. It could define a new fixed point. For example, there could be “crossover scaling solutions” interpolating as a function of $\tilde{\chi}$ from eq. (232) for $\tilde{\chi}\to 0$ to the trivial scaling solution for $\tilde{\chi}\to\infty$.
The simplified flow equation Wetterich (2024) employs a field-dependent cutoff. In this approach one finds the scaling solution (2). The corrections due to the field-dependence of the cutoff can be computed and are found to remain consistent with the scaling solution (2).
Appendix C Free composite scalar fields
In this appendix we employ microscopic field transformations with a linear RG-kernel in order to discuss under which conditions a microscopic model with a non-quadratic classical action can describe a free composite scalar field. The issue concerns the question if one can find a microscopic realization of the scaling solution (2) or (232).
Simple microscopic field transformations
There are particular choices of microscopic variable transformations for which the RG-kernel remains simple and the flow equation keeps its closed form. Consider a transformation of the type
${\varphi}^{\prime}_{k}(x)=kf\left(\frac{\chi^{\prime}(x)}{k}\right)=kf\big{(}\tilde{\chi}^{\prime}(x)\big{)}\ ,$ (233)
which is adapted to scale invariant scalar fields $\tilde{\chi}^{\prime}=\chi^{\prime}/k$ in four dimensions. It implies
$\partial_{t}{\varphi}^{\prime}_{k}(x)=k\partial_{k}{\varphi}^{\prime}_{k}(x)=\left(1-\frac{\partial\ln f}{\partial\ln\tilde{\chi}}(x)\right){\varphi}^{\prime}_{k}(x)\ .$ (234)
For certain forms of $f$ one finds a simple behavior for $\partial_{t}{\varphi}^{\prime}_{k}(x)$, as
${\varphi}^{\prime}_{k}=ck\tilde{\chi}^{\prime\alpha}\ ,\quad\partial_{t}{\varphi}^{\prime}_{k}(x)=(1-\alpha){\varphi}^{\prime}_{k}(x)\ ,\gamma_{k}=(1-\alpha){\varphi}\ .$ (235)
This results in the flow equation
$\partial_{t}\Gamma_{{\varphi},k}=$ $\,\frac{1}{2}\mathrm{tr}\Big{\{}\big{(}\partial_{t}\mathcal{R}_{k}+(2-2\alpha)\mathcal{R}_{k}\big{)}\big{(}\Gamma_{{\varphi},k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}\Big{\}}$
$-(1-\alpha)\int_{x}\frac{\partial\Gamma_{{\varphi},k}}{\partial{\varphi}(x)}{\varphi}(x)\ .$ (236)
Another simple choice is
${\varphi}^{\prime}_{k}=ck\ln\tilde{\chi}^{\prime}\ ,\quad\partial_{t}{\varphi}^{\prime}_{k}(x)={\varphi}^{\prime}_{k}(x)-ck\ ,$ (237)
with flow equation
$\partial_{t}\Gamma_{{\varphi},k}=$ $\,\frac{1}{2}\mathrm{tr}\Big{\{}\big{(}\partial_{t}\mathcal{R}_{k}+2\mathcal{R}_{k}\big{)}\big{(}\Gamma_{{\varphi},k}^{(2)}+\mathcal{R}_{k}\big{)}^{-1}$
$-\int_{x}\frac{\partial\Gamma_{{\varphi},k}}{\partial{\varphi}(x)}\big{(}{\varphi}(x)-ck\big{)}\ .$ (238)
For these choices the flow equation keeps a simple closed form. This can be useful for approximations valid for certain ranges of the field. The flow equations (C) and (C) are particular cases of the more general setting of linear RG-kernels that lead to eq. (V).
Free composite scalar fields
A particular solution of the flow equations (V), (C), (C) is the effective action for a free massless scalar field,
$\Gamma_{{\varphi},k}=\frac{Z(k)}{2}\int_{x}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}\ .$ (239)
The first term in the flow equation (V) contributes only a ${\varphi}$-independent constant that may be neglected, and we conclude
$\partial_{t}\Gamma_{{\varphi},k}=-\int_{x}Z(k)a(k)\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}\ .$ (240)
Comparison of coefficients yields an anomalous dimension
$\partial_{t}Z=-2aZ\ ,\quad Z=Z_{0}\left(\frac{k}{k_{0}}\right)^{-2a}\ ,\quad\eta=-\partial_{t}\ln Z=2a\ ,$ (241)
with $a=1-\alpha$ and $\alpha=0$ for eq. (C). (This can be extended to free massive particles by adding to $\Gamma_{{\varphi},k}$ a $k$-dependent mass term, and possibly also a term linear in ${\varphi}$.)
If for large $k=\Lambda$ the microscopic effective average action $\Gamma_{{\varphi},\Lambda}[\chi^{\prime}]$ can be brought to the form (239) (with a possible additional mass term) by a suitable microscopic field transformation (129) (131), this form is preserved by the flow. This defines a family of models that correspond to free composite scalar fields. For a model formulated in terms of the canonical field $\chi^{\prime}$, as for our setting (3), the issue concerns the initial value $\Gamma_{{\varphi},\Lambda}[\chi]$. We found in sect. VI that for $\Lambda\to\infty$ one has $\Gamma_{\chi,\Lambda}[\chi]=S[\chi]$, while $\Gamma_{{\varphi},\Lambda}[\chi]$ will involve additional terms reflecting the Jacobian of the microscopic field transformation. For a non-linear kinetial (3) the classical action $S[\chi]$ needs additional terms, typically a scalar potential, in order to achieve the initial value (233) for $\Gamma_{{\varphi},\Lambda}$ despite the presence of a Jacobian. In contrast, a measure $\sim\text{d}\chi^{\prime}/\chi^{\prime}$, which is linear in ${\varphi}^{\prime}$ for the transformation (166), leads directly to an initial value of the form (239).
Conditions for linear RG-kernels
We next discuss a few properties of the condition (131) for linear RG-kernels. The “linearity condition” (131) imposes a simple constraint on the scale dependence of $\tilde{N}$,
$\partial_{t}\tilde{N}=$ $\,\partial_{t}\frac{\partial{\varphi}^{\prime}}{\partial\chi^{\prime}}=\left(1+\frac{\partial}{\partial\ln\chi^{\prime}}\right)\partial_{t}N$
$=$ $\,\left(1+\frac{\partial}{\partial\ln\chi^{\prime}}\right)\left(a(k)N+\frac{b(k)}{\chi^{\prime}}\right)=a(k)\tilde{N}\ .$ (242)
The general solution reads
$\tilde{N}=\left(\frac{Z(k)}{Z_{0}}\right)^{-\frac{1}{2}}\hat{N}(\chi^{\prime})\ ,\quad\eta(k)=-\partial_{t}\ln Z(k)=2a(k)\ ,$ (243)
with $\hat{N}(\chi^{\prime})$ independent of $k$. We conclude that the general form of the field transformation which leads to a linear RG-kernel reads
${\varphi}^{\prime}=\left(\frac{Z(k)}{Z_{0}}\right)^{-1/2}\int_{\chi_{0}}^{\chi^{\prime}}\text{d}\chi^{\prime\prime}\,\hat{N}(\chi^{\prime\prime})+f(k)\ .$ (244)
Choosing $Z(k)=Z_{0}$, $f(k)=0$, $\hat{N}(\chi^{\prime})=K^{1/2}(\chi^{\prime})$ one can bring the microscopic action (3) into the form for a free field ${\varphi}^{\prime}$, provided $K(\chi^{\prime})$ is positive and obeys some monotony properties. For a model with classical action (3) the microscopic effective average action $\Gamma_{{\varphi},\Lambda}[{\varphi}]$ differs, however, from the microscopic or classical action $S[\chi^{\prime}({\varphi})]$. Additional potential terms in $S[\chi^{\prime}]$ are necessary for a microscopic effective action of the form (239).
Eq. (131) admits logarithmic solutions,
${\varphi}^{\prime}=c(k)\ln\left(\frac{\chi^{\prime}}{k}\right)\ ,\quad a=\partial_{t}\ln c\ ,\quad b=-c\ .$ (245)
With
$\tilde{N}(\chi^{\prime})=\frac{\partial{\varphi}^{\prime}}{\partial\chi^{\prime}}=\frac{c(k)}{\chi^{\prime}}\ ,\quad K(\chi^{\prime})=Z(k)\tilde{N}^{2}(\chi^{\prime})\ ,$ (246)
this maps
$K(\chi^{\prime})\partial^{\mu}\chi^{\prime}\partial_{\mu}\chi^{\prime}=Z(k)\frac{c^{2}(k)}{\chi^{\prime 2}}\partial^{\mu}\chi^{\prime}\partial_{\mu}\chi^{\prime}=Z(k)\partial^{\mu}{\varphi}^{\prime}\partial_{\mu}{\varphi}^{\prime}\ .$ (247)
The flow of $Z$ according to eq. (241) becomes
$\partial_{t}Z=-2\partial_{t}\ln(c)Z\ .$ (248)
In particular, for a powerlaw behavior of $c(k)$ the anomalous dimension is constant
$c(k)=c_{0}\left(\frac{k}{k_{0}}\right)^{a}\ ,\quad\partial_{t}Z=-2aZ\ ,\quad\eta=2a\ .$ (249)
A particularly simple setting obtains by the limit $a\to 0$ of eq. (245) by choosing $c$ independent of $k$, such that $a=0$, $\eta=0$, and
$\Gamma_{{\varphi},k}=\frac{1}{2}\int_{x}\frac{Z_{0}c_{0}^{2}}{\chi^{2}}\partial^{\mu}\chi\partial_{\mu}\chi=\frac{1}{2}\int_{x}Z_{0}\partial^{\mu}{\varphi}\partial_{\mu}{\varphi}\ .$ (250)
For the particular field transformation,
${\varphi}^{\prime}=c_{0}\ln\left(\frac{\chi^{\prime}}{k}\right)\ ,$ (251)
the correction term in the flow equation vanishes (except for an irrelevant constant) despite the $k$-dependence of the field transformation.
Scale symmetry
The quantum effective action (250) realizes quantum scale symmetry Wetterich (2019). It is invariant under the scale- or dilatation transformations
$\chi(x)\to\beta\chi(x)\ ,\quad{\varphi}(x)\to{\varphi}(x)+\beta\ .$ (252)
In this simple case quantum scale symmetry can be related to classical scale symmetry. We can start with a scale invariant classical action
$\widehat{S}[\chi^{\prime}]=\frac{1}{2}\int_{x}\frac{Z_{0}c_{0}^{2}}{\chi^{\prime 2}}\partial^{\mu}\chi^{\prime}\partial_{\mu}\chi^{\prime}\ ,$ (253)
and a scale invariant measure
$\int\mathcal{D}\chi^{\prime}=\prod_{x}\int_{-\infty}^{\infty}\frac{\text{d}\chi^{\prime}(x)}{\chi^{\prime}(x)}\ .$ (254)
This is related to the formulation (4) with measure (6) by an action
$S[\chi^{\prime}]=\widehat{S}[\chi^{\prime}]+\ln J[\chi^{\prime}]\ ,$ (255)
with $J[\chi^{\prime}]$ the Jacobian (14).
With
$\frac{\text{d}\chi^{\prime}}{\chi^{\prime}}=\frac{\text{d}{\varphi}^{\prime}}{c_{0}}$ (256)
the Jacobian of the field transformation to ${\varphi}^{\prime}$ does not involve the scale $k$. It absorbs the term $\ln J[\chi^{\prime}]$ such that the microscopic or classical action for ${\varphi}^{\prime}$ is the one for a free massless field. Up to a multiplicative constant in the partition function or an additive constant in the effective action, this setting describes a free massless scalar field theory for the field ${\varphi}^{\prime}$, with measure $\prod_{x}\int\text{d}{\varphi}^{\prime}(x)$ and classical action
$\widehat{S}[{\varphi}^{\prime}]=\frac{Z_{0}}{2}\int_{x}\partial^{\mu}{\varphi}^{\prime}\partial_{\mu}{\varphi}^{\prime}\ .$ (257)
The difference to the action $\bar{S}[{\varphi}^{\prime}]$ in eq. (15) results from the different measure (254).
In this simple setting there are no field-dependent fluctuation effects. Up to an additive constant this results for $k=0$ directly in the effective action (250). We observe that the source term is invariant under the simultaneous shift ${\varphi}^{\prime}(x)\to{\varphi}^{\prime}(x)+\beta$, ${\varphi}(x)\to{\varphi}(x)+\beta$, which translates classical scale symmetry to quantum scale symmetry. An infrared cutoff term quadratic in ${\varphi}^{\prime}$ effectively maintains scale symmetry, if it can be written in the form
$\Delta_{k}S[{\varphi}^{\prime}]=\frac{1}{2}\int_{x}\partial^{\mu}{\varphi}^{\prime}\hat{r}_{k}\left(-\frac{\partial^{\mu}\partial_{\mu}}{k^{2}}\right)\partial_{\mu}{\varphi}^{\prime}\ .$ (258)
The issue is somewhat subtle if $\hat{r}_{k}(y)$ diverges for $y\to 0$. For example, $\hat{r}_{k}=y^{-1}$ results in a mass term $k^{2}$ in the classical action supplemented by the cutoff, with a similar term in $\Gamma_{{\varphi},k}^{\prime}[{\varphi}]$. This seems to indicate a violation of scale symmetry by the cutoff. The mass term is subtracted, however, by the definition of $\Gamma_{{\varphi},k}[{\varphi}]$, explaining the $k$-independence of $\Gamma_{{\varphi},k}$ also for $k>0$. For the measure (254), (256) we can write the functional integral defining $\Gamma_{{\varphi},k}$ in terms of $\hat{\varphi}(x)={\varphi}^{\prime}(x)-{\varphi}(x)$ which is invariant under the simultaneous shift in ${\varphi}^{\prime}$ and ${\varphi}$.
For a field-independent cutoff the definition of $\Gamma_{\chi,k}$ violates scale symmetry by the cutoff quadratic in $\chi^{\prime}$. In contrast, the formal definition of the quantum effective action $\Gamma_{\chi}[\chi]$ $(k=0)$ preserves scale symmetry if $S[\chi^{\prime}]$ is scale invariant. The source term is invariant under a simultaneous scaling of $\chi^{\prime}$ and $\chi$. The Jacobian for global rescalings of the linear measure (6) is a field independent factor which only adds an irrelevant additive constant to $\Gamma_{\chi}$. One may therefore ask if there exists an appropriate setting for the flow equation that maintains scale symmetry in case of the measure (6) for scale invariant $S$. For the measure (6) the transition to $\Gamma_{\varphi}$ violates scale symmetry by the Jacobian of the field transformation. Then $\Gamma_{\varphi}$ is not expected to be scale symmetric. For $\Gamma_{\chi}$ one may imagine a cutoff term involving the macroscopic field $\chi$ in the form $\mathcal{R}_{k}\sim\chi^{-2}$. In this setting the cutoff is invariant under simultaneous scale transformations of $\chi^{\prime}$ and $\chi$. We discuss this case of a field-dependent cutoff in appendices AB, where one finds indeed a scaling solution with scale symmetry.